### v1

# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

[…]
Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?
[…]
Can We Validate Counterfactual Estimations in the Presence of General Network Interference?
[…]
Network interference, message-passing model, experimental design, network data, high-dimensional data
1 Introduction
Randomized Control Trials (RCTs) are the gold standard for evaluating intervention effectiveness, such as assessing public campaigns to promote voting within communities (Imbens and Rubin, 2015). However, classic RCT methods often overlook the complex dynamics of belief adoption within social networks. In these networks, experimental units—individual voters, for instance—do not make decisions in isolation but are influenced by peers, family, and societal norms. A study of the 2010 US congressional elections by Bond et al., (2012) demonstrated that encouraging individuals to vote influenced not only their own behavior but also that of their friends and those beyond their immediate social circles.
This phenomenon, known as network interference, violates a fundamental RCT assumption—the Stable Unit Treatment Value Assumption (SUTVA)—which requires that the treatment status of one unit does not affect the outcomes of other units. This violation precludes clear separation of treatment and control groups, a challenge prevalent in modern social science contexts (Imbens, 2024). This raises a central question across scientific fields: “How can we efficiently estimate and validate the causal effect of an intervention within a network of interacting units?” (Eckles et al., 2016, Cai et al., 2015, Abaluck et al., 2022, Ogburn et al., 2024).
To articulate the challenges in the estimation and validation of causal effects, we first introduce a formal notation and visualization. Consider a randomized experiment conducted over time stamps $t=0,1,\ldots,T$, involving a population of $N$ units, indexed by $i=1,\ldots,N$. At each time $t$, unit $i$ receives a treatment (e.g., exposure to the voting campaign) according to a random variable $W^{i}_{t}$. In the voting example, these treatment variables are binary: 1 indicates assignment to the treatment group, and 0 indicates assignment to the control group. [^1] Our theoretical results apply to the more general case when the treatments can also be continuous-valued.
The treatment allocation across the entire population and time frame is represented by an $N\times(T+1)$ matrix, denoted by $\bm{W}{}$. Following the Rubin causal framework (Imbens and Rubin, 2015), we denote by $Y^{i}_{t}{}(\bm{W}{})$ the potential outcome of unit $i$ at time $t$. The counterfactual evolution, denoted by $\text{CFE}_{t}{(\bm{W}{})}$, represents the sequence of sample means of potential outcomes under the treatment assignment $\bm{W}{}$ over time [^2] Typically, outcomes are “non-anticipating,” meaning $\text{CFE}_{t}{(\bm{W}{})}$ and $Y^{i}_{t}{(\bm{W}{})}$ depend only on treatments up to time $t$. Our notation and theoretical results allow for more general scenarios where outcomes may be influenced by future treatments. However, for ease of reading, readers may assume the simpler case of non-anticipation where only the first $t+1$ columns of $\bm{W}{}$ impact $Y^{i}_{t}{(\bm{W}{})}$. Moreover, the first column of $\bm{W}{}$ is by convention equal to all zeros (or no-treatment state), and the first column of $\bm{Y}{}$ corresponds to outcomes in the all-control state.:
$\text{CFE}_{t}{(\bm{W}{})}:=\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{(\bm{W}{})},\quad\quad\quad t=0,1,\ldots,T.$ (1)
With $\bm{w}{}$ being a specific realization of $\bm{W}{}$, let $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ denote an $N\times(T+1)$ matrix that collectively represents the observed outcomes under the treatment allocation $\bm{w}{}$. For example, if per-unit treatment probability at all time periods is equal to $p$, $\text{CFE}_{t}{(\bm{W}{})}$ evolves as a stochastic function of time and $p$, where the randomness arises from the treatment allocation $\bm{W}{}$. Figure 1 visualizes this evolution as a two-dimensional surface parameterized by time and $p$, where each contour (at fixed $p$) represents one possible dynamic trajectory. [^3] Note that treatment probabilities can vary over time, making the space of possible counterfactual evolutions $\text{CFE}_{t}{(\bm{W}{}=\bm{w}{})}$ considerably more complex than the simplified visualization shown in Figure 1.
Figure 1: Evolution of outcomes sample mean (z-axis) with respect to time $t$ and treatment probability $p$. Red and blue contours highlight the counterfactual evolutions at treatment probabilities $p=0.25$ and $p=0.75$, respectively. The magenta line represents the equilibrium state, where the treatment effect has stabilized.
Our objective is to estimate counterfactual evolutions $\text{CFE}_{t}{(\bm{w}{}^{\prime})}$ under alternative treatment allocations $\bm{w}{}^{\prime}\neq\bm{w}{}$ and validate the accuracy of these estimations. This enables estimation of treatment effects over time through comparing counterfactual evolutions under different treatment conditions—for example, in the voting study, contrasting evolution under an intensive campaign against the status quo control. However, this objective faces a fundamental challenge: we can observe the system only under a single scenario, reflecting “the fundamental problem of causal inference” (Holland, 1986). Specifically, with treatment probability fixed at $p$, we observe only the corresponding contour in the surface shown in Figure 1. This limitation constrains both our ability to estimate counterfactual evolutions and validate such estimates, as we cannot directly observe outcomes under alternative scenarios.
To address the estimation challenge, we develop our approach by drawing an analogy between two systems: a network of units exchanging causal effects and a molecular system exchanging energy. In molecular systems, energy perturbations trigger redistribution, according to invariant physical laws governing molecular interactions, until a new equilibrium is reached. Similarly, in social networks, targeted interventions propagate through network connections following consistent social and behavioral principles until the system stabilizes. This analogy highlights a crucial insight: the temporal invariance of underlying mechanisms. Just as the equations governing molecular energy redistribution maintain their structural form, the mechanisms driving the propagation of intervention effects preserve their fundamental structure over time. This structural invariance presents a key opportunity: by observing the system’s evolution across different time periods, we can employ machine learning tools to learn the mathematical functions that characterize how intervention effects propagate through the network and aggregate into counterfactual evolutions [^4] This structural invariance differs from stationarity, as we allow the underlying mechanisms to exhibit temporal variations while maintaining their basic mathematical form..
To address the validation challenge, a natural candidate is cross-validation—a cornerstone of statistical learning (Hastie et al., 2009) whose theoretical foundations (Stone, 1974, Allen, 1974) build upon Fisher’s fundamental insight about the necessity of randomization for independent validation (Fisher, 1935). However, cross-validation’s application in causal inference has been notably limited by data scarcity. Existing analyses typically focus on observations at equilibrium reached at large $t$, depicted by the “equilibrium” curve in Figure 1 (Basse et al., 2019, Jackson et al., 2020, Li and Wager, 2022b ), providing insufficient samples for reliable validation. Even recent approaches that leverage temporal pre-equilibrium data (Shirani and Bayati, 2024, Bayati et al., 2024) provide only order $T$ summary statistics. This represents a significant constraint, given that extending experiment duration is often prohibitively costly (Holtz et al., 2020, Cooprider and Nassiri, 2023, Xiong et al., 2024a ). To overcome this limitation, we introduce a new distribution-preserving network bootstrap method (DPNB) that generates sufficient data to implement cross-validation, enabling rigorous validation of our estimation results.
By combining the temporal invariance principle with DPNB, we create a comprehensive framework that addresses both estimation and validation challenges in settings with unobserved network interference. Specifically, this work makes four main contributions. First, we extend the Causal Message-Passing framework (Shirani and Bayati, 2024) by incorporating unit- and network-level heterogeneities, deriving more general equations for outcome evolution. Second, we develop non-asymptotic versions of these dynamic equations using techniques from Li and Wei, (2022), leading to our DPNB method for generating and analyzing multiple subpopulations from the experimental population. This approach yields sufficient observations at each time period for accurate estimation through machine learning techniques.
Third, we combine temporal data with our DPNB method to develop a cross-validation framework for counterfactual estimation, implementing data-driven model selection procedures and preprocessing for complex time-trends to improve accuracy. Our DPNB method may have applications beyond causal inference, particularly in network sampling problems, where it offers an approach for generating representative subsamples from networked populations without requiring prior knowledge of network structure.
Our fourth contribution addresses a fundamental challenge in causal inference: the rigorous examination of estimation methods when ground-truth counterfactuals are typically unobservable. We develop a comprehensive benchmark toolbox with six diverse experimental environments that capture different forms of network interference and temporal dynamics. These environments encompass applications ranging from social network influence to data center load balancing and ride-sharing systems. A notable highlight is our social network environment, which simulates a social media platform through thousands of AI agents interacting with their local network connections via content feeds. Each environment is carefully constructed to maintain known ground truth values while incorporating realistic complexities such as time-varying trends and heterogeneous network effects. We use this toolbox to extensively evaluate our method against existing benchmarks. Importantly, we are making these environments publicly available to facilitate standardized testing of future methodologies. [^5] Code and environments available at https://github.com/CausalMP/CausalMP.git. This contribution advances the field by providing researchers with a much-needed framework for systematic evaluations of causal inference methods under general network interference (Basse and Airoldi, 2018, Sävje et al., 2021, Arkhangelsky and Imbens, 2023, Imbens, 2024).
The rest of the paper is organized as follows. Section 2 reviews related literature on causal inference under network interference, approximate message passing, and network sampling. Section 3 presents our generalization of the Causal Message-Passing framework (Shirani and Bayati, 2024). Section 4 introduces our distribution-preserving network bootstrap method for generating sufficient data to learn invariant aggregate dynamics. Section 5 presents our cross-validation framework for counterfactual estimation. Sections 6 and 7 describe our benchmark toolbox and empirical validation results. Section 8 concludes with discussion and future directions. Technical proofs and supplementary results appear in the appendices.
2 Related Literature
Recent literature on causal inference under network interference has evolved along several key dimensions. One common approach addresses settings with known network clusters (Sobel, 2006, Rosenbaum, 2007, Hudgens and Halloran, 2012, Tchetgen and VanderWeele, 2012, Auerbach and Tabord-Meehan, 2021), where researchers randomly assign entire clusters to different treatment intensities. This cluster-based design helps contain interference within cluster boundaries, though it requires prior knowledge of network clusters. Another significant development employs “exposure mappings” (Manski, 2013, Aronow and Samii, 2017, Leung, 2022, Harshaw et al., 2023, Sävje, 2024), which quantify how a unit’s treatment effect varies based on its neighbors’ treatment status. For scenarios with unknown interference patterns, researchers have proposed alternative strategies using historical data or staggered treatment rollouts (Yu et al., 2022, Cortez et al., 2022). A particularly promising direction has emerged through application-specific approaches that develop customized interference models for distinct contexts such as marketplace dynamics (Bajari et al., 2021, Holtz et al., 2020, Wager and Xu, 2021, Munro et al., 2021, Johari et al., 2022, Bright et al., 2022). These domain-specific solutions offer tailored frameworks for managing interference in their respective experimental settings. Our work contributes to this literature by developing methods that do not require knowledge of the interference structure or domain-specific context while still capturing complex network effects across diverse applications.
Our methodology is supported by rigorous theoretical results derived from Approximate Message-Passing (AMP), whose foundations trace back to Thouless et al., (1977), Kabashima, (2003) and Donoho et al., (2009) and were formally established by Bolthausen, (2014), Bayati and Montanari, (2011) and a large body of recent literature (Javanmard and Montanari, 2013, Bayati et al., 2015, Berthier et al., 2020, Chen and Lam, 2020, Zhong et al., 2021, Wang et al., 2022, Dudeja et al., 2023, Rush and Venkataramanan, 2018, Li and Wei, 2022). At its core, we employ the TAP equation and cavity method—a mathematical technique that characterizes system behavior by analyzing responses to carefully introduced perturbations (Mezard et al., 1986, Mezard and Montanari, 2009). AMP traditionally addresses high-dimensional estimation problems through iterative systems with known mixing matrices. Our approach is different and builds on the work of Shirani and Bayati, (2024); specifically, they assume the mixing matrix is unknown and develop techniques similar to those developed in the AMP literature to analyze observed outcomes and infer causal effects.
Our work on distribution-preserving network bootstrap relates to a rich body of research on network sampling and statistical inference in networked settings. The challenge of generating representative samples from network data has been central to network analysis, due to the inherent dependencies in network structures (Wormald, 1999, Newman, 2018). Network inference methods have evolved from basic configuration models (Bollobás, 1980) to sophisticated approaches handling non-independent observations (Snijders and Borgatti, 1999, Bickel and Chen, 2009) and robust bootstrapping techniques (Bhattacharyya and Bickel, 2015, Green and Shalizi, 2022). Recent advances encompass null models for complex network structures (Karrer and Newman, 2011, Bianconi, 2018) and machine learning approaches using graph neural networks (Kipf and Welling, 2017, Hamilton, 2020). Our approach differs fundamentally from this literature as it operates solely on node-level outcomes observed over time, without requiring access to the underlying network structure.
3 General Counterfactual Estimation
Estimating counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ based on the observed outcomes $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ enables addressing a broad range of causal questions, such as, “What if we had delivered the treatments according to $\bm{w}{}^{\prime}$ instead of $\bm{w}{}$?” For example, if outcomes are observed after treating 20% of potential voters, campaign organizers may want to explore how the population would have responded if 40% had received the campaign materials. In addition, by estimating the dynamics of the counterfactuals over time, we gain insights into how the treatment effect may strengthen or weaken as time progresses.
The total treatment effect (TTE) provides a formal way to compare any two counterfactual scenarios. For two treatment allocations $\bm{w}{}^{\prime}$ and $\bm{w}{}^{\prime\prime}$, the TTE measures the difference in population average outcomes:
$\text{TTE}_{t}{(\bm{w}{}^{\prime\prime},\bm{w}{}^{\prime})}=\text{CFE}_{t}{(\bm{w}{}^{\prime\prime})}-\text{CFE}_{t}{(\bm{w}{}^{\prime})},\quad\quad\quad t=0,1,\ldots,T\,.$ (2)
The literature typically focuses on a special case: when all entries of $\bm{w}{}^{\prime\prime}$ equal one and all entries of $\bm{w}{}^{\prime}$ equal zero, with all outcomes at equilibrium (Yu et al., 2022, Candogan et al., 2023, Ni et al., 2023, Ugander and Yin, 2023). However, this extreme scenario of treating the entire population versus treating no one is often impractical. As we demonstrate in subsequent sections, our framework enables the estimation of TTEs between any two treatment allocations, providing decision-makers with more realistic comparisons (Muralidharan and Niehaus, 2017, Egger et al., 2022).
3.1 Experimental Design Framework
We proceed by generalizing the experimental design. Specifically, we allow treatment assignments $W^{i}_{t}$ to vary over time following a Bernoulli distribution with mean $p_{t}$, and define the experimental design as the product distribution $\Pi=p_{0}\times\ldots\times p_{T}$. In this context, we use $\vec{W}_{t}{}:=(W^{1}_{t},\ldots,W^{N}_{t})^{\top}$ to represent the vector containing the treatment assignments for all experimental units at time $t$. This formulation encompasses a broad range of experimental designs, including staggered roll-out design (Xiong et al., 2024a ), micro-randomized trials (Li and Wager, 2022a ), and switchback experiments (Bojinov et al., 2023).
Remark 3.1
In Appendix 9, we further generalize this framework by allowing $W^{i}_{t}$ to follow more general probability distributions $\pi_{t}$ and defining the experimental design as $\Pi=\pi_{0}\times\ldots\times\pi_{T}$. The random variable $W^{i}_{t}$ can take values in either an integer set (e.g., different treatment types) or a continuous set (e.g., varying treatment doses). This generalization distinguishes our work from recent literature on network interference and longitudinal data, which primarily focuses on binary treatments (Arkhangelsky and Imbens, 2023).
For a fixed integer $n^{\text{x}}$, we can also consider covariates (might be observed or not) in the form of a $n^{\text{x}}$ by $N$ matrix $\bm{X}$, where each column (denoted by ${\vec{X}}^{i}:=(X^{1i},\ldots,X^{n^{\text{x}}i})^{\top}$) represents characteristics of unit $i$ (e.g., age and gender). The experimental data thus consists of treatment assignments $\bm{w}{}$, observed outcomes
$\bm{Y}{}(\bm{w}{}):=\bm{Y}{}(\bm{W}{}=\bm{w}{})\,,$
and covariates $\bm{X}$. Therefore, we observe outcomes only under one specific treatment allocation $\bm{w}{}$—a single realization among $2^{N(T+1)}$ distinct potential outcomes, where this number grows exponentially with population size and time horizon. Consequently, estimating causal effects under general interference is impossible due to non-identifiability (Karwa and Airoldi, 2018). To address this challenge, we propose a tractable outcome specification that aligns with and extends the causal message passing framework (Shirani and Bayati, 2024).
3.2 Potential Outcome Specification
Let $\bm{W}_{t}{}:=[\vec{W}_{0}{}|\ldots|\vec{W}_{t}{}]$ denote the treatment assignments up to time $t$; we represent by $\vec{Y}_{t}{}(\bm{W}_{t}{})=\big{(}Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})\big{)}^{\top}$ the potential outcome vector at time $t$. Consider unknown functions $g_{t}$ and $h_{t}{}$ which operate component-wise, and the expressions $g_{t}\big{(}\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big{)}$ and $h_{t}{}\big{(}\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\big{)}$ represent the corresponding column vectors. Given $\vec{Y}_{0}{}(\bm{W}_{0}{})$ as the initial outcome vector, we consider the following specification for potential outcomes:
$\vec{Y}_{t+1}{}(\bm{W}_{t+1}{})=\big{(}{\bf A}+{\bf B}_{t}\big{)}g_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+h_{t}{}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)+\vec{\epsilon}^{\;}_{t},\;\;t=0,1,\ldots,T-1,$ (3)
where ${\bf A}$ and ${\bf B}_{t}$ are $N\times N$ unknown matrices capturing the fixed and time-dependent interference effects, respectively. Additionally, $\vec{\epsilon}^{\;}_{t}=\big{(}\epsilon^{1}_{t},\ldots,\epsilon^{N}_{t}\big{)}^{\top}$ is the zero-mean noise term accounting for observation noise. Denoting by ${\text{A}}^{ij}$ and ${\text{B}}^{ij}_{t}$ the element in the $i^{th}$ row and $j^{th}$ column of ${\bf A}$ and ${\bf B}_{t}$, respectively, the value ${\text{A}}^{ij}+{\text{B}}^{ij}_{t}$ quantifies the impact of unit $j$ on unit $i$ at time $t$.
The specification in Eq. (3) captures several aspects of experimental data. It accommodates various types of interference, including treatment spillover effects, carryover effects, peer effects, and autocorrelation (Shirani and Bayati, 2024). Notably, this specification accounts for outcome dynamics by acknowledging the temporal interrelation of units’ outcomes. This contrasts with existing approaches to panel data analysis, which assume that time labels can be shuffled without affecting causal effects (Arkhangelsky and Imbens, 2023).
Remark 3.2
In Appendix 9, we further generalize the outcome specification in Eq. (3) to incorporate additional lag terms (e.g., $\vec{Y}_{t-1}{}(\bm{W}_{t-1}{})$) in the functions $g_{t}$ and $h_{t}{}$, the complete treatment matrix at any time point (allowing for anticipation effects), and time-dependent covariates.
Next, we analyze the state evolution of the experimental population, characterizing the asymptotic dynamics of unit outcomes. These theoretical foundations establish the basis for developing robust counterfactual estimators in subsequent sections.
3.3 Experimental State Evolution
In this section, we analyze the distribution of unit outcomes over time. This analysis requires the following two assumptions about the interference matrices. {assumption}[Gaussian interference structure] For all $i,j$, the element ${\text{A}}^{ij}$ in the $i^{th}$ row and $j^{th}$ column of ${\bf A}$ is an independent Gaussian random variable with mean $\mu^{ij}/N$ and variance $\sigma^{2}/N$. Similarly, ${\text{B}}^{ij}_{t}$, the element in the $i^{th}$ row and $j^{th}$ column of ${\bf B}_{t}$, is an independent Gaussian random variable with mean $\mu^{ij}_{t}/N$ and variance $\sigma^{2}_{t}/N$. {assumption}[Convergent interference pattern - informal statement] For all unit $i$ and any time $t$, the elements of vector $(\mu^{i1},\ldots,\mu^{iN})$ admit a weak limit [^6] This means that the empirical distribution of ${\mu^{i1},\ldots,\mu^{iN}}$ converges to a probability distribution as $N$ increases. and, separately, the elements of vector $(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})$ admit a weak limit, where both limits are invariant in $i$. According to Assumptions 3.3 and 3.3, the impact of unit $j$ on unit $i$ is captured by $\mu^{ij}$, $\mu^{ij}_{t}$, and two centered Gaussian random variables. This generalizes the model of Shirani and Bayati, (2024), which assumes i.i.d. interference matrix elements across all units. This extension accommodates more heterogeneous local interactions and varying levels of uncertainty about the interference structure. Specifically, a fully known interference network has exact values for $\mu^{ij}$ and $\mu^{ij}_{t}$ with $\sigma^{2}=0$ and $\sigma^{2}_{t}=0$, while a completely unknown interference means no knowledge of these quantities. Importantly, our estimation method is designed to handle the cases that we have no knowledge of these underlying quantities for implementation.
Example 3.3
Consider the case where $\mu^{ij}$ and $\mu^{ij}_{t}$ take values of 00 and $1$, generating time-dependent adjacency matrices of the graph representing the interference structure. If we have high confidence that interactions occur only through these adjacency matrices, we can imagine negligible values for $\sigma^{2}$ and $\sigma^{2}_{t}$. Conversely, significant uncertainty about potential interactions would be reflected in larger values of $\sigma^{2}$ and $\sigma^{2}_{t}$.
Remark 3.4
The formal version of Assumption 3.3 appears in Assumption 9.4 in Appendix 9.4, where we generalize the condition by allowing greater variation across units. This interference model accommodates complex interaction patterns and admits further generalizations, including extensions to non-Gaussian interference matrices. For detailed discussions, see the Appendix of Shirani and Bayati, (2024).
We then establish the following informal result for the large-sample regime that characterizes how outcome distributions evolve between consecutive time points, with its formal statement presented in Appendix 9.4.
Theorem 3.5 (State evolution - informal statement)
Let $N\to\infty$ and suppose Assumptions 3.3 and 3.3 hold. There exist mappings $f_{t},\;t=0,\ldots,T-1$, such that the distribution of outcomes at time $t+1$ is determined by:
$\mathcal{L}(Y_{t+1}{})=f_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{})\right),$ (4)
where $Y_{t}{}$ denotes the weak limit of unit outcomes $Y^{1}_{t}{}(\bm{W}_{t}{}),\ldots,Y^{N}_{t}{}(\bm{W}_{t}{})$ at time $t$, $W_{t+1}{}$ follows a Bernoulli distribution with mean $p_{t+1}$, and $\mathcal{L}(\cdot)$ refers to law or distribution of its argument.
The mapping $f_{t}$ in Eq. (4) characterizes how the distribution of outcomes at time t+1 emerges from the previous distribution at time $t$ through the functions $g_{t}$ and $h_{t}{}$, while accounting for both direct treatment effects and indirect effects from the interference matrices and unit covariates. Specifically, in the large-sample limit, even though each unit’s outcome depends on a complex network of interactions, the population-level distribution of outcomes follows a simpler evolution that depends only on the previous distribution and treatment assignment distribution. Building on (Shirani and Bayati, 2024), while extending their results to a broader setting, we refer to the relationship in (4) as the experimental state evolution (SE) equation.
3.4 A First Look at Estimation Through Theory and Practical Constraints
The state evolution equation (4) suggests a natural estimation strategy, motivated by the observation that in many settings, $f_{t}$ remains relatively stable over time. Specifically, we can often decompose $f_{t}$ into a substantial time-invariant component $f$ and a time-varying component.
This decomposition, combined with the fact that empirical distributions from experimental data with many units should approximate the distributions in equation (4), enables a supervised learning approach to estimate $f$. When strong time trends are present, one can first detrend the data to isolate the time-invariant component before applying these learning techniques. Once these mappings are estimated, we can generate any desired counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ through recursive application of Eq. (4) (see Algorithm 3 in Appendix 10). For the sake of building intuition, we will explain a simple version of this, adapted from Shirani and Bayati, (2024).
3.4.1 An Illustrative Example
Consider a special case of outcome specification (3) where functions $h_{t}{}$ are equal to zero, and functions $g_{t}$ satisfy,
$g_{t}\big{(}Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1},{\vec{X}}^{i}\big{)}=\delta+\theta Y^{i}_{t}{}(\bm{W}_{t}{})+\tau W^{i}_{t+1}+\lambda Y^{i}_{t}{}(\bm{W}_{t}{})W^{i}_{t+1}\,.$ (5)
In this case, if we denote the outcomes sample mean at time $t$ by $\nu_{t}{}$, the state evolution implies that as $N\rightarrow\infty$,
$\nu_{t+1}{}=f(\nu_{t}{},p_{t+1})\,,$ (6)
where $f$ has the form $f(a,b)=\delta+\theta a+\tau b+\lambda ab$, and plase the role of the aforementioned time-invariant function. Subsequently, given observations $\bm{Y}{}(\bm{w}{})$ and $\bm{w}{}$, Algorithm 1 enables consistent estimation of CFE for any target treatment allocation $\bm{w}{}^{\prime}$, provided that $\bm{w}{}$ and $\bm{w}{}^{\prime}$ share identical columns for treatment assignment at time $t=0$ and the design set $\{p_{0},\ldots,p_{T}\}$ contains at least two distinct values [^7] A detailed analysis of a more general setting appears in Appendix 10.1..
Algorithm 1 Causal message passing counterfactual estimator (simple case)
$\bm{Y}{}(\bm{w}{}),\bm{w}{}=[w^{i}_{t}{}]_{i,t}$, and $\bm{w}{}^{\prime}=[w^{{}^{\prime}i}_{t}{}]_{i,t}$
Step 1: Data processing
for $t=0,\ldots,T$ do
     $\hat{\nu}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{}(\bm{w}_{t}{})$,   $\hat{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{t}{}$,   $\check{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{{}^{\prime}i}_{t}{}$
end for
Step 2: Parameters estimation
$(\hat{\delta},\hat{\theta},\hat{\tau},\hat{\lambda})$leftarrow Estimation of $(\delta,\theta,\tau,\lambda)$ using OLS in $\hat{\nu}_{t+1}{}=\delta+\theta\hat{\nu}_{t}{}+\tau\hat{p}_{t+1}{}+\lambda\hat{\nu}_{t}{}\hat{p}_{t+1}{}$
Step 3: Counterfactual estimation
$\widehat{\text{CFE}}_{0}{(\bm{w}{}^{\prime})}\leftarrow\hat{\nu}_{0}{}$
for $t=1,\ldots,T$ do
     $\widehat{\text{CFE}}_{t}{(\bm{w}{}^{\prime})}\leftarrow\hat{\delta}+\hat{\theta}\widehat{\text{CFE}}_{t-1}{(\bm{w}{}^{\prime})}+\hat{\tau}\check{p}_{t}{}+\hat{\lambda}\widehat{\text{CFE}}_{t-1}{(\bm{w}{}^{\prime})}\check{p}_{t}{}$
end for
Estimated counterfactual evolution: $\{\widehat{\text{CFE}}_{t}{(\bm{w}{}^{\prime})}\}_{t=0}^{T}$.
3.4.2 General Estimation Theory
The above example demonstrates that our counterfactual estimation problem reduces to the consistent estimation of $f$. This relationship can be formalized for our general outcome specification (3). We provide an informal version of this result below, while the formal statement and complete proof appear in Theorem 10.1 in Appendix 10.
Theorem 3.6 (Consistency - informal statement)
Under certain regularity conditions, given consistent estimation of the mappings $f_{t}$ in the SE equation (4), any desired counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ can be consistently estimated.
3.4.3 Practical limitations
The effectiveness of approaches like Algorithm 1 in identifying treatment effects across various settings has been established by Shirani and Bayati, (2024). However, two significant constraints require consideration. First, the estimation procedure in the second step of Algorithm 1 faces sample size limitations due to the experiment horizon. Second, as demonstrated by Bayati et al., (2024), certain settings require accounting for more complex structures, necessitating a richer class of functions $f$. To capture these more complex settings, one approach is to use higher moments of the outcome and treatment distributions in state evolution (4).
For example, if we denote the variance of outcomes at time $t$ by $\rho_{t}{}$, and if functions $g$ and $h{}$ contain second-order terms, the state evolution becomes:
$\Big{(}\nu_{t+1}{},(\rho_{t+1}{})^{2}\Big{)}=f\Big{(}\nu_{t}{},(\rho_{t}{})^{2},p_{t+1},p_{t+1}^{2}\Big{)}\,.$ (7)
An alternative approach to capturing complex structure is to consider a richer family of functions $f$ during the estimation phase.
These considerations raise two key questions: First, how can we expand the sample size to improve estimation accuracy? Second, how can we determine the appropriate specification for the underlying experiment, similar to model selection in supervised learning? The following two sections address these questions systematically.
4 Distribution-preserving Network Bootstrap
To enable the use of more powerful supervised learning techniques, we introduce a theoretically supported resampling methodology that generates additional samples from the state evolution equation.
The methodology centers on batching experimental units into subpopulations, denoted by $\mathcal{S}$. Specifically, $\mathcal{S}$ represents a subset of units from $\{1,\ldots,N\}$ where membership is determined exclusively by the treatment allocation $\bm{W}{}$. Then, under the randomized treatment assumption—which specifies that $\bm{W}{}$ is random and independent of all other variables—each $\mathcal{S}$ constitutes a random sample from the experimental population. In this section, we demonstrate that each distinct selection of $\mathcal{S}$ yields a new observation of the state evolution equation, provided the empirical treatment distributions exhibit sufficient variation across subpopulations.
Our analysis proceeds in two stages. First, we extend the state evolution concept to derive subpopulation-specific state evolutions. We establish that our batching technique generates statistically efficient samples of the experimental state evolution while preserving its fundamental structure—hence the term distribution-preserving network bootstrap (DPNB). Second, we derive a decomposition rule that characterizes the finite-sample behavior of potential outcomes. This decomposition enables the identification of endogenous noise within the experimental data and elucidates how DPNB addresses this challenge.
Remark 4.1
The theoretical frameworks presented here extend, with minimal or no modification, to arbitrary subpopulations. While our analysis focuses specifically on treatment-allocation-based subpopulations in this section, the appendices provide complete theoretical statements applicable to general subpopulation structures.
4.1 Subpopulation-specific State Evolution
We begin by extending the state evolution analysis to subpopulations of experimental units.
Theorem 4.2 (informal statement)
Under the conditions of Theorem 3.5, consider a subpopulation $\mathcal{S}$ whose size $\left|{\mathcal{S}}\right|$ grows to infinity as $N\to\infty$. There exist mappings $\varphi_{t},\;t=0,\ldots,T-1$, such that:
$\mathcal{L}(Y^{\mathcal{S}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{\mathcal{S}}_{t}{}),\mathcal{L}(W^{\mathcal{S}}_{t+1}{})\right),$ (8)
where $Y^{\mathcal{S}}_{t}{}$ and $W^{\mathcal{S}}_{t+1}{}$ represent the weak limits of outcomes and treatments for units in subpopulation $\mathcal{S}$, while $Y_{t}{}$ and $W_{t+1}{}$ are their corresponding population-level analogs as defined in Theorem 3.5.
According to Equation (8), each subpopulation’s outcome distribution follows a distinct state evolution equation governed by mappings $\varphi_{t},\;t=0,\ldots,T-1$. The outcomes for units within $\mathcal{S}$ evolve through a dynamic interplay between population-level and subpopulation-specific outcomes and treatments. The formal statement of Theorem 4.2 appears in Theorem 9.8 of Appendix 9.4, where (27) explicitly defines $\varphi_{t}$ through mappings $g_{t}$ and $h_{t}{}$.
To demonstrate the efficacy of DPNB, consider two subpopulations $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ such that $W^{\mathcal{S}_{1}}_{t+1}{}\neq W^{\mathcal{S}_{2}}_{t+1}{}$ in (8). Then, two sequences $Y^{\mathcal{S}_{1}}_{0}{},\ldots,Y^{\mathcal{S}_{1}}_{T}{}$ and $Y^{\mathcal{S}_{2}}_{0}{},\ldots,Y^{\mathcal{S}_{2}}_{T}{}$ yield distinct observations of the state evolution described by (8). Furthermore, for each time step $t$, the mapping $\varphi_{t}$ exhibits fundamental similarities to $f_{t}$, with the experimental state evolution equation in (4) emerging as a special case of (8) when applied to the entire population. Accordingly, by strategically selecting subpopulations with distinct treatment allocations, we obtain multiple state evolution observations, enabling the estimation of $\varphi_{t}$ and, consequently, $f_{t}$.
Example 4.3
Consider two distinct subpopulations: treated units ($\mathcal{S}={\cal T}$) and control units ($\mathcal{S}={\cal C}$). From (8), we have:
$\mathcal{L}(Y^{{\cal T}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{{\cal T}}_{t}{}),1\right),$ (9)
$\mathcal{L}(Y^{{\cal C}}_{t+1}{})=\varphi_{t}\left(\mathcal{L}(Y_{t}{}),\mathcal{L}(W_{t+1}{}),\mathcal{L}(Y^{{\cal C}}_{t}{}),0\right).$
When the direct treatment effect is non-zero, the outcome distributions for the treated group $Y^{{\cal T}}_{t}{}$ and the control group $Y^{{\cal C}}_{t}{}$ are distinct, providing two different observations from the state evolution.
4.2 Outcomes Decomposition Rule
Building on the finite-sample analysis of AMP algorithms (Li and Wei, 2022), we obtain the following decomposition rule for the outcome specification in Eq. (3).
Theorem 4.4 (Unit-level decomposition rule)
Suppose Assumption 3.3 holds and let $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{T-1}$ denote i.i.d. random vectors in $\mathbb{R}^{N}$ following ${\cal N}(0,\frac{1}{N}{\bf I}_{N})$ distribution. For any unit $i$ and fixed $t\in\{0,\ldots,T-1\}$, we have
$Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=$ $\frac{1}{N}\sum_{j=1}^{N}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)+h_{t}{}\left(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i}\right)$ (10)
$+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|g_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)\right\|}R^{i}_{t}+\epsilon^{i}_{t},$
where $R^{i}_{t}$ represents a random variable such that $\vec{R}_{t}:=\left(R^{1}_{t},\ldots,R^{N}_{t}\right)^{\top}=\sum_{i=0}^{t}\beta^{i}_{t}\vec{Z}_{i}$, with $\vec{\beta}_{t}=(\beta^{0}_{t},\ldots,\beta^{t}_{t},0,\ldots,0)^{\top}\in\mathbb{R}^{N}$ denoting a random vector that is correlated with $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$ and satisfies ${\|\vec{\beta}_{t}\|}=1$. Furthermore,
${\cal W}_{1}\left(\mathcal{L}\left(\vec{R}_{t}\right),{\cal N}\left(0,\frac{1}{N}{\bf I}\right)\right)\leq c\sqrt{\frac{t\log N}{N}},$
where $c$ is a constant independent of $N$ and $t$, $\mathcal{L}(\vec{R}_{t})$ denotes the probability distribution of $\vec{R}_{t}$, and ${\cal W}_{1}$ is Wasserstein-1 distance.
Equation (10) provides a decomposition of unit outcomes: the first term captures the interference effect, the second term reflects the unit-specific effect, the third term arises from uncertainties in the network structure, and the fourth term accounts for observation noise. We can then adjust the first term in Eq. (10) to reflect the available knowledge about the interference network.
Example 4.5
When it is known that no interference exists, Eq. (10) simplifies to $Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=h_{t}{}\big{(}Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i}\big{)}+\epsilon^{i}_{t}$, representing a general specification for the potential outcomes.
Example 4.6
Consider a scenario where the interference network is partially known and exhibits a clustered structure with clusters $C^{1},\ldots,C^{K}$ (e.g., the experimental population consists of individuals from different cities). In this context, Eq. (10) can be rewritten as follows:
$Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=$ $\frac{1}{N}\sum_{k=1}^{K}\Bigg{[}\sum_{j\in C^{k},i\notin C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)$
$+\sum_{i,j\in C^{k}}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)\Bigg{]}$
$+h_{t}{}\left(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i}\right)+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|g_{t}\left(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X}\right)\right\|}R^{i}_{t}+\epsilon^{i}_{t},$
where we expect the magnitude of $\left(\mu^{ij}+\mu_{t}^{ij}\right)$ to be relatively larger whenever $i,j\in C^{k}$, indicating that units within the same cluster exhibit stronger ties.
The outcome decomposition rule in Eq. (10) leads to the following corollary, which characterizes the decomposition of the outcome sample mean for units belonging to subpopulation $\mathcal{S}$.
Corollary 4.7 (Subpopulation-level decomposition rule)
Under the assumptions of Theorem 4.4, we have
$\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t+1}{}(\bm{W}_{t+1}{})=$ $\frac{1}{N\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\sum_{j=1}^{N}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{}(\bm{W}_{t}{}),W^{j}_{t+1}{},{\vec{X}}^{j}\right)$ (11)
$+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}h_{t}{}\left(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i}\right)$
$+\sqrt{\frac{\sigma^{2}+\sigma_{t}^{2}}{{\left|{\mathcal{S}}\right|}}}{\left\|g_{t}(\vec{Y}_{t}{}(\bm{W}_{t}{}),\vec{W}_{t+1}{},\bm{X})\right\|}J_{t}+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\epsilon^{i}_{t}.$
where $\left|{\mathcal{S}}\right|$ denotes the size of the subpopulation $\mathcal{S}$ and $J_{t}$ is a random variable satisfying
${\cal W}_{1}\left(\mathcal{L}\left(J_{t}\right),{\cal N}\left(0,\frac{1}{N}\right)\right)\leq c\sqrt{\frac{t\log N}{N}}.$
While Theorem 4.2 establishes that distinct subpopulations yield different state evolution observations, Corollary 4.7 reveals two crucial implications for finite-sample analysis. First, the outcome distribution within each subpopulation incorporates unit-level heterogeneities, as demonstrated by the first two terms in Eq.(11). Second, the outcome uncertainty terms—represented by the final two expressions in Eq.(11)—decay at a rate of $\sqrt{\left|{\mathcal{S}}\right|}$. The subsequent sections examine these implications in detail, showing how DPNB addresses such observational complexities while generating additional samples.
4.3 Unit-level Heterogeneities
Experimental units exhibit distinct covariates, such as age and gender. Network interference also emerges through diverse local patterns, jointly inducing heterogeneous characteristics. These heterogeneities are captured in (11) through variations in $\mu^{ij}$, $\mu_{t}^{ij}$, and ${\vec{X}}^{i}$. This foundation motivates our strategy of selecting members for each subpopulation $\mathcal{S}$ solely based on treatment allocation. Specifically, this approach enables us to treat $\mathcal{S}$ as a random sample, supporting the assumption that each subpopulation represents the original experimental population. Indeed, Theorem 4.2 demonstrates the effectiveness of this batching strategy, evidenced by the invariance of $\varphi_{t}$ across subpopulations. Meanwhile, the result of Corollary 4.7 suggests maximizing subpopulation sizes to leverage stronger smoothing effects through averaging over larger groups of units.
4.4 Endogenous Observation Noise
The noise terms in the outcome decomposition rule exhibit complex correlation structures. According to Theorem 4.4, the random vectors $\vec{R}_{0},\ldots,\vec{R}_{T-1}$ in Eq. (10) asymptotically follow a centered Gaussian distribution as population size $N$ increases, thus functioning as noise terms. These vectors, however, demonstrate a complex endogenous dependency structure through several mechanisms.
Initially, for each $t$, the random vector $\vec{R}_{t}$ derives its randomness from $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$. Consequently, the random vectors $\vec{R}_{0},\ldots,\vec{R}_{T-1}$ exhibit correlation, generating temporal correlation in the observational noise.
Additionally, in finite populations where $N$ is bounded, the elements of $R^{i}_{t}$ demonstrate cross-unit correlation. The magnitude of this correlation is governed by the constant $c$ in Theorem 4.4, which varies across settings and remains unidentified.
Furthermore, the potential outcome of unit $i$ at time $t$ denoted by $Y^{i}_{t}{}(\bm{W}_{t}{})$ incorporates $R^{i}_{t-1}$, which correlates with elements of $\vec{R}_{t}$; this introduces correlation between noise terms and unit outcomes. The strength of this correlation depends on the magnitudes of $\sigma$ and $\sigma_{t}$, which remain unknown and can vary substantially across different settings.
These three factors collectively define a complex endogenous dependency structure, indicating how uncertainty in the interference structure can compromise outcome measurements and introduce substantial bias when overlooked. The following example examines an autoregressive model with unanticipated unit interactions, demonstrating how even minimal second-order interference induces endogeneity and generates biased estimates.
Example 4.8
Consider a simplified version of the outcome specification in Eq. (3) where $g_{t}(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i})=Y^{i}_{t}{}(\bm{W}_{t}{})$ and $h_{t}{}(Y^{i}_{t}{}(\bm{W}_{t}{}),W^{i}_{t+1}{},{\vec{X}}^{i})=\Theta^{i}Y^{i}_{t}{}(\bm{W}_{t}{})$, with $\mu^{ij}+\mu_{t}^{ij}=0$ for all $i$, $j$, and $t$. These conditions reflect the available partial information confirming negligible interference between units. Also, assume $\epsilon^{i}_{t}$ is independent from all other variables. We can write:
$Y^{i}_{t+1}{}(\bm{w}_{t+1}{})=\Theta^{i}Y^{i}_{t}{}(\bm{w}_{t}{})+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|\vec{Y}_{t}{}(\bm{w}_{t}{})\right\|}R^{i}_{t}+\epsilon^{i}_{t}.$
To estimate $\Theta^{i}$ for a specific unit $i$, we employ ordinary least squares regression, regressing outcomes $Y^{i}_{1}{}(\bm{w}_{1}{}),\ldots,Y^{i}_{T}{}(\bm{w}_{T}{})$ on their corresponding lagged values $Y^{i}_{0}{}(\bm{w}_{0}{}),\ldots,Y^{i}_{T-1}{}(\bm{w}_{T-1}{})$. Under the asymptotic regime where $N\rightarrow\infty$, we characterize the bias of estimator $\hat{\Theta}^{i}$ as follows:
$\mathbb{E}\left[\hat{\Theta}^{i}\Big{|}\bm{Y}{}(\bm{w}{})\right]-\Theta^{i}$ $=\frac{1}{\sum_{t=0}^{T-1}Y^{i}_{t}{}(\bm{w}_{t}{})^{2}}\sum_{t=0}^{T-1}\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|\vec{Y}_{t}{}(\bm{w}_{t}{})\right\|}Y^{i}_{t}{}(\bm{w}_{t}{})\mathbb{E}\left[R^{i}_{t}\big{|}\bm{Y}{}(\bm{w}{})\right].$
Based on the results of Corollary 4.7, the magnitude of this complex noise diminishes as the subpopulation size $\left|{\mathcal{S}}\right|$ increases, supporting the selection of larger subpopulations. Therefore, in the extreme case, when $\mathcal{S}$ encompasses the entire population ($\left|{\mathcal{S}}\right|=N$), the noise magnitude decays at a rate of $\sqrt{N}$.
4.5 How to use DPNB to generate efficient samples?
Both unit-level heterogeneities and endogenous observation noise suggest maximizing subpopulation sizes, which would ideally lead to selecting the entire experimental population. This approach aligns with the proposal in Section 3.4, where we observe a single instance of state evolution. Therefore, this presents a nuanced trade-off: larger subpopulations increase overlap between groups, reducing variation across subpopulations and thereby decreasing the number of effective samples. Given that the magnitude of endogenous noise varies across different settings, we aim to develop a data-driven methodology to address this trade-off optimally. This leads us to propose a cross-validation method, which we examine in detail in the subsequent section.
5 Counterfactual Cross Validation
The precision of estimated counterfactuals, as established in Theorem 3.6, depends on accurately estimating the state evolution mappings $f_{t}$. This estimation process faces two main challenges that require careful validation:
Function approximation and model selection
The true specification of potential outcomes is often unknown, requiring us to approximate the state evolution mappings $f_{t}$. This approximation involves two related aspects. First, we must choose appropriate summary statistics of the joint distribution of outcomes and treatments to serve as inputs to our model - an approach analogous to feature engineering in supervised learning. For example, Shirani and Bayati, (2024) used sample means of outcomes, treatments, and their products. While domain knowledge can guide this selection, we need a systematic way to validate these choices. Second, depending on the complexity of experimental settings, we may need to employ a range of estimation techniques, from simple linear regression to more sophisticated methods such as neural networks. The choice of technique and its specific implementation (e.g., architecture, hyperparameters) must be validated to prevent issues like model instability and estimation bias. These two aspects are linked, as both contribute to how well we can approximate the true mapping $f_{t}$. A data-driven validation methodology helps us jointly optimize these choices while reducing misspecification risks.
Optimal batch configuration
As established in §4.5, DPNB offers an effective strategy for addressing unit-level heterogeneities and controlling endogenous noise through batching. However, the choice of batch size presents a bias-variance trade-off. Larger batches better average out heterogeneities and noise, reducing estimation bias, but they also decrease the number of distinct batches available for learning, increasing estimation variance. Conversely, smaller batches provide more samples for learning, reducing variance, but may retain more heterogeneity and noise, potentially introducing bias. Beyond batch size, the number of batches creates a similar trade-off that interacts with the batch size selection. Therefore, data-driven validation helps find the optimal combination of batch size and number of batches that balances these competing effects in a given experimental context.
5.1 Counterfactual Cross Validation Algorithm
To address these challenges, we propose a counterfactual cross-validation algorithm that divides the time horizon into a list of time blocks $\mathfrak{L}_{\mathsf{tb}}$, which serve as natural cross-validation folds. For each time block, we use all other blocks as training data and the held-out block as validation data. The training process involves generating batches as discussed above, with their size and number serving as tuning parameters. For validation, we use a fixed set of pre-specified validation batches $\mathcal{S}_{1}^{v},\ldots,\mathcal{S}_{b_{v}}^{v}$ constructed as follows. We begin by computing each unit’s average treatment exposure across the experimental horizon (e.g., a unit that is treated in 60% of the time periods receives a treatment exposure of 0.6). We then rank the units in descending order based on their treatment exposure values and partition them into $b_{v}$ equal-sized groups, ensuring validation batches cover the full spectrum of treatment exposure. For each fold, the validation metric is averaged across these validation batches to provide a robust performance measure.
The parameters being selected through this cross-validation include both the choice of estimator from a candidate set $\mathfrak{L}_{\mathsf{E}}$ and the batching configuration from a candidate set $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$. Each estimator $\mathsf{E}\in\mathfrak{L}_{\mathsf{E}}$ represents a specific combination of feature generation (what summary statistics to use) and estimation method (e.g., linear regression or neural network). The batching parameters $(\mathsf{s_{b}},\mathsf{n_{b}})\in\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$ specify the size $\mathsf{s_{b}}$ and number $\mathsf{n_{b}}$ of training batches. For each held-out time block, we train models using all possible combinations of these parameters on the remaining blocks and evaluate their performance on the validation batches obtained from the held-out block.
Step 1: Reference counterfactual construction
After the experiment is completed, we obtain the following data. The treatment matrix $\bm{w}{}$ and observed outcomes matrix, $\bm{Y}{}(\bm{w}{})$. We first compute the sample mean of outcomes over time as ground truth (counterfactual) for evaluating estimations in subsequent steps. More precisely, for each time period $t\in\{0,1,\ldots,T\}$, and each validation batch $\mathcal{S}_{j}^{v}$ with $j\in\{1,\ldots,b_{v}\}$, we calculate average of outcomes at time $t$ across units in $\mathcal{S}_{j}^{v}$ and denote that by $\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}$.
Step 2: Leave-one-out and counterfactual estimation
For each time block $\mathsf{tb}\in\mathfrak{L}_{\mathsf{tb}}$, we construct training datasets $\bm{y}_{\text{train}}^{-\mathsf{tb}}$ and $\bm{w}_{\text{train}}^{-\mathsf{tb}}$ as submatrices of $\bm{Y}{}(\bm{w}{})$ and $\bm{w}{}$, respectively, excluding columns within $\mathsf{tb}$. We define $\bm{w}_{\text{test}}^{\mathsf{tb}}$ as the submatrix of $\bm{w}{}$ containing only columns in $\mathsf{tb}$. Each estimator $\mathsf{E}$ is then trained using $\bm{y}_{\text{train}}^{-\mathsf{tb}}$ and $\bm{w}_{\text{train}}^{-\mathsf{tb}}$, and used to estimate counterfactuals for treatment allocation $\bm{w}_{\text{test}}^{\mathsf{tb}}$ during the held-out time block $\mathsf{tb}$.
Step 3: Optimal Estimator Selection
Following the estimation across all time blocks, we identify the optimal estimator and batch parameters by comparing the results with observed counterfactuals using a predetermined loss function. For instance, using mean square error:
$\text{MSE}_{\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}}}=\frac{1}{b_{v}(T+1)}\sum_{j=1}^{b_{v}}\sum_{t=0}^{T}\Big{[}\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}-\widehat{\text{CFE}}_{t}^{\mathcal{S}_{j}^{v}}{(\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}})}\Big{]}^{2}.$
Algorithm 2 Counterfactual cross-validation
Treatment allocation $\bm{w}{}$, observed outcomes $\bm{Y}{}(\bm{w}{})$, validation batches $\{\mathcal{S}_{j}^{v}\}_{j=1}^{b_{v}}$, time blocks $\mathfrak{L}_{\mathsf{tb}}$, loss function $\ell:\mathbb{R}^{(T+1)b_{v}}\times\mathbb{R}^{(T+1)\times b_{v}}\to\mathbb{R}_{+}$, and candidate estimators $\mathfrak{L}_{\mathsf{E}}$ and batch parameters $\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$
Step 1: Reference Counterfactual Construction
$\big{\{}\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}\big{\}}_{t=0}^{T}\leftarrow$ mean of $\bm{Y}{}(\bm{w}{})$ for units belonging to $\mathcal{S}_{j}^{v},\;j=1,\ldots,b_{v}$
Step 2: Leave-one-out and Counterfactual Estimation
for $\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}}\in\mathfrak{L}_{\mathsf{E}}\times\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}$ do
     for $\mathsf{tb}\in\mathfrak{L}_{\mathsf{tb}}$ do
         $\bm{y}_{\text{train}}^{-\mathsf{tb}}\leftarrow$ columns of $\bm{Y}{}(\bm{w}{})$ outside of $\mathsf{tb}$
         $\bm{w}_{\text{train}}^{-\mathsf{tb}}\leftarrow$ columns of $\bm{w}{}$ outside of $\mathsf{tb}$
         $\bm{w}_{\text{test}}^{\mathsf{tb}}\leftarrow$ columns of $\bm{w}{}$ in $\mathsf{tb}$
         Train $\mathsf{E}$ with batching parameters $(\mathsf{s_{b}},\mathsf{n_{b}})$ on data $\bm{y}_{\text{train}}^{-\mathsf{tb}}$ and $\bm{w}_{\text{train}}^{-\mathsf{tb}}$
         for $j\in\{1,\ldots,b_{v}\}$ do
              $\big{\{}\widehat{\text{CFE}}_{t}^{\mathcal{S}_{j}^{v}}{(\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}})}\big{\}}_{t\in\mathsf{tb}}\leftarrow$ estimate CFE for $\bm{w}_{\text{test}}^{\mathsf{tb}}$ via the trained model, for units in $\mathcal{S}_{j}^{v}$
         end for
     end for
end for
Step 3: Optimal Estimator Selection
$\mathsf{E}^{*},\mathsf{s_{b}}^{*},\mathsf{n_{b}}^{*}\leftarrow\arg\min_{(\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}})\in\mathfrak{L}_{\mathsf{E}}\times\mathfrak{L}_{(\mathsf{s_{b}},\mathsf{n_{b}})}}\ell\left(\Big{\{}\big{\{}\text{CFE}_{t}^{\mathcal{S}_{j}^{v}}\big{\}}_{t=0}^{T}\Big{\}}_{j=1}^{b_{v}},\Big{\{}\big{\{}\widehat{\text{CFE}}_{t}^{\mathcal{S}_{j}^{v}}{(\mathsf{E},\mathsf{s_{b}},\mathsf{n_{b}})}\big{\}}_{t=0}^{T}\Big{\}}_{j=1}^{b_{v}}\right)$
Remark 5.1
Algorithms 5 and 7, presented in Appendices 10 and 11 respectively, provide examples of candidate estimators based on linear regression.
Remark 5.2
In experimental settings with strong temporal patterns such as seasonality, we first detrend the observed outcomes, e.g., see Algorithm 7 in Appendix 11. This procedure augments Step 2 of Algorithm 2 as follows. For each estimator $\mathsf{E}$, we utilize the complete observed dataset to estimate a baseline counterfactual—defined as the counterfactual outcome for all units under control—using an estimation model optimized for temporal pattern detection. We then generate filtered data by subtracting this baseline counterfactual from observed outcomes. This filtered data serves as input for the main estimator, which focuses specifically on treatment effect identification. The final estimates are obtained by combining the baseline counterfactual with the estimated treatment effects. The consistency of this estimation approach relies on a structural assumption of weak additive separability between baseline outcomes and treatment effects.
6 Benchmark Toolbox
We evaluate our framework using six semi-synthetic experiments that combine simulated environments with real-world data. This approach offers two key advantages: it maintains realistic data characteristics while allowing us to compute ground truth values for our estimands. Unlike real experimental settings where outcomes are observed under a single scenario, these settings provide ground truth values for any desired scenario. This enables rigorous evaluation of estimation methods under realistic conditions. In the following sections, we specify the treatment allocation, outcomes, and network structure for each experimental setting.
6.1 LLM-based Social Network
This environment simulates a social media platform like Facebook and Quora where user interactions occur through content feeds, designed to study the effects of feed ranking algorithms on user engagement. The environment employs Large Language Model (LLM) agents to represent users, with demographically realistic personas derived from US Census data (U.S. Census Bureau, 2023) following Chang et al., (2024)’s methodology. Each agent possesses two interests selected from a predefined set of keywords: Technology, Sports, Politics, Entertainment, Science, Health, and Fashion.
The treatment variable is the feed ranking algorithm, which determines content ordering for each user. In the control condition, users receive randomly ordered content, while the treatment condition presents content weighted by friend engagement. The outcomes of interest are user engagement metrics, measured as the sum of likes or replies generated by each user within a given time period, tracked in a comprehensive panel dataset.
The underlying directed follower-following relationships network is constructed based on a preferential attachment model (Barabási and Albert, 1999). The dynamics of the environment center on content generation and user interactions. Content originates from an LLM-generated bank focused on a specified topic, with cross-keyword variations (e.g., climate change intersecting with technology or politics). The system generates feeds by combining interest-based content (matching user interests), trending content (high engagement), and random content according to specified proportions. Users interact with their feeds through an LLM-driven decision process that considers content relevance, friend engagement, feed position, and demographic characteristics.
The simulation maintains comprehensive state information across multiple dimensions: engagement metrics, user interaction histories, network relationships, content visibility, and conversation threads. At each time step, agents have a 1% probability of generating new content, contributing to the platform’s organic content evolution. The interaction process follows a structured decision framework, where each agent evaluates content through a detailed prompt incorporating the agent’s persona attributes, content characteristics, social signals (including friend engagement), and feed positioning. This framework ensures that user behaviors and social influence patterns evolve naturally through the network while maintaining computational tractability.
6.2 Belief Adoption Model
This environment models the diffusion of competing opinions within interconnected communities, implementing the cascade model of Montanari and Saberi, (2010). The system examines how opinions spread through social networks when individuals make decisions through coordination games with their neighbors. Through this framework, we evaluate the effectiveness of promotional campaigns in influencing opinion adoption patterns.
The environment considers two competing stances: Opinion A (e.g., voting in an election) and Opinion B (e.g., declining to vote). The treatment represents a campaign aimed at increasing Opinion A adoption. The outcome for each unit in each period is binary: 1 if they adopt Opinion A, and 0 otherwise.
The opinion evolution follows a network-based coordination game where each individual $i$ assigns payoff values $\text{payoff}^{i}_{\text{A}}$ and $\text{payoff}^{i}_{\text{B}}$ to both opinions. The probability of Opinion A adoption in the next period depends on the neighbor configuration and relative payoffs:
$\mathbb{P}\big{(}\text{adopting Opinion~{}$\text{A}$}\big{|}n^{i}_{\text{A}},n^{i}_{\text{B}}\big{)}=\frac{e^{\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}}{e^{\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}+e^{-\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}}\,,$
where $h^{i}=\frac{\text{payoff}^{i}_{\text{A}}-\text{payoff}^{i}_{\text{B}}}{\text{payoff}^{i}_{\text{A}}+\text{payoff}^{i}_{\text{B}}}$ and $n^{i}_{\text{A}}$ represents the number of neighbors holding Opinion A out of $n^{i}$ total neighbors in the current period, and $\beta$ is a predetermined constant. The underlying network in our simulator is derived from the Pokec social network dataset (Takac and Zabovsky, 2012, Leskovec and Krevl, 2014), focusing on three regional networks: Krupina (3,366 users), Topolcany (18,246 users), and Zilina (42,971 users).
The environment also utilizes detailed profile data from the Pokec social network dataset to characterize each user. We extract three demographic variables from these profiles: age, profile activity, and gender. For base payoffs, we assign higher values for Opinion A to users aged 25-55 years and those maintaining active profiles. The treatment effectiveness follows a Gaussian distribution, reaching its maximum at age 35. Its impact scales directly with the user’s profile completion percentage, as measured in their Pokec data. These profile-driven calculations create distinct patterns of payoffs and treatment responses across the network, as visualized in Figures 2-4.
Figure 2: Distribution of base payoffs, treatment effects, and node degrees for individuals in Krupina.
Figure 3: Distribution of base payoffs, treatment effects, and node degrees for individuals in Topolcany.
Figure 4: Distribution of base payoffs, treatment effects, and node degrees for individuals in Zilina.
6.3 Ascending Auction Model
This environment simulates a competitive market where multiple bidders participate in an ascending auction for objects, following the model of Bertsekas, (1990). The auction mechanism creates a dynamic pricing system where bidder interactions generate complex patterns of market influence, even without direct object-to-object relationships.
The system operates with $N$ objects and $N$ bidders. Each object represents an experimental unit, with its final value in each round serving as the outcome variable. The treatment consists of promotional interventions that increase bidder valuations by $\tau$for randomly selected objects. This treatment affects all bidders interested in the selected objects.
The market evolution follows a structured bidding process. In each round, bidders evaluate objects based on their private valuations and current market prices. They submit bids for their preferred objects, with objects being assigned to the highest bidders. These assignments establish new price levels, which influence subsequent bidding behavior. As prices increase through competitive bidding, objects become progressively less attractive to competing bidders.
The environment demonstrates a unique form of interference: while objects do not directly influence each other, treatment effects propagate through the market via bidders’ strategic responses to price changes. This creates a network of indirect treatment effects, as promotional interventions for certain objects can influence market outcomes for others through shifts in bidder behavior and price dynamics.
6.4 New York City Taxi Routes
This environment models ride-sharing dynamics across New York City taxi zones using real-world TLC Trip Record Data (New York City Taxi and Limousine Commission, 2024). The framework adapts the established linear-in-mean outcome model (Eckles et al., 2016, Cai et al., 2015, Leung, 2022) to represent how passengers utilize ride-sharing services throughout the city. By incorporating actual travel data, passenger density metrics, and inter-zone relationships, the simulation effectively captures the complex network of interactions between taxi routes across the city.
In this setting, the experimental units are defined as routes (origin-destination pairs) between the city’s 263 taxi zones, with time segmented into 6-hour periods. The outcome variable measures the number of trips along each route during each time period. The treatment represents a program implemented on randomly selected routes and the goal is to evaluate travelers response patterns.
Given baseline outcomes $[y^{i}_{t}{}]_{i,t}$, the system’s evolution follows Equation (12), where outcomes depend on baseline patterns, network effects, and treatment status:
$Y^{i}_{t+1}{}=y^{i}_{t+1}{}+\theta\sum_{j=1}^{N}E^{ij}(Y^{j}_{t}{}-y^{i}_{t}{})+\tau_{\text{p}}\sum_{j=1}^{N}E^{ij}W^{j}_{t+1}+\tau^{i}_{\text{u}}W^{i}_{t+1},\quad t\geq 1,$ (12)
where we initiate the recursion by setting $Y^{i}_{0}{}=y^{i}_{0}{}$ and $W^{i}_{0}=0$ for all $i$. Here, ${\bf E}=[E^{ij}]_{i,j}$ represents the normalized route adjacency matrix, $\tau_{\text{u}}^{i}$ represents route-specific direct treatment effects, and parameters $(\theta,\tau_{\text{p}})=(0.4,0.2)$ control autocorrelation and spillover effects.
Figure 5: Mean and variance of the number of trips in each route during each period, revealing a strong daily and weekly seasonality pattern.
The environment incorporates real-world data through three components: First, it uses “High Volume For-Hire Vehicle Trip Records” (January-March 2024, 57,974,677 trips) to construct baseline outcomes (denoted by $[y^{i}_{t}{}]_{i,t}$), focusing on 18,768 active routes across 366 periods. As shown in Figure 5, the trips have a strong seasonality pattern, which is common in the ride-hailing application (Xiong et al., 2024b ). Second, it employs the “Taxi Zone Lookup Table” and Claude (Model 3.5 Sonnet, Anthropic, 2024) to generate passenger density scores that determine route-specific treatment effects (Figure 6 left). Third, it creates a route adjacency network based on geographic proximity, transit connections, shared roads, and functional relationships, yielding an average node degree of 8.32 (Figure 6 right). This data-driven approach ensures the simulation reproduces key real-world characteristics: temporal patterns, heterogeneous treatment effects, and localized network interactions.
Figure 6: The left panel displays the distribution of route-specific direct treatment effects, illustrating the heterogeneity in treatment responses. The right panel shows the histogram of node degrees in the route interference network.
6.5 Exercise Encouragement Program
This environment simulates an exercise intervention program that combines individual characteristics from the 1994 Census Bureau database (Kohavi and Becker, 1994) with social network effects. Drawing inspiration from mobile health intervention studies (Liao et al., 2016, Klasnja et al., 2015, 2019), the environment models how digital encouragement messages influence exercise decisions within a social network context.
The experimental units are individuals, with binary outcomes representing their exercise decisions in each period (1 for exercise, 0 for no exercise). The treatment consists of digital intervention messages designed to encourage physical activity. Inspired by Li and Wager, 2022a , we let the outcomes follow a Bernoulli distribution defined as follows:
$Y^{i}_{t+1}{}\sim\text{Bernoulli}\left(\frac{1}{1+\exp{-(\delta_{t}^{i}+\tau_{t}^{i}W^{i}_{t+1}+\theta Y^{i}_{t}{}Z^{i}_{t}+\lambda W^{i}_{t+1}Y^{i}_{t}{}Z^{i}_{t})}}\right),$ (13)
where $Z^{i}_{t}=\sum_{j=1}^{N}E^{ij}Y^{j}_{t}{}$ represents the count of neighboring individuals who exercised in the previous period. Here, $E^{ij}$’s are the elements of the adjacency matrix from Twitter social circles data (Leskovec and Mcauley, 2012), with an average of 21.74 connections per individual (Figure 9).
In (13), each component captures specific aspects of exercise behavior. The baseline probability ($\delta_{t}^{i}$) represents an individual’s inherent tendency to exercise, derived from Census Bureau demographic data. This probability incorporates age (with higher values for younger individuals), working hours (showing an inverse relationship with exercise likelihood), and occupation type (assigning higher probabilities to active or professional occupations). These baseline probabilities exhibit weekly patterns, showing peak values on weekends due to increased free time, elevated rates on Mondays from new week motivation, stable mid-week patterns, and slightly lower values on Fridays reflecting end-of-week fatigue (Figure 7).
The intervention effectiveness ($\tau_{t}^{i}$) quantifies how individuals respond to exercise encouragement messages. This response varies based on multiple demographic factors from the Census Bureau database. Younger individuals show higher responsiveness to digital interventions, while education level correlates positively with intervention effectiveness. Job-related factors, including occupation type and working hours, influence response rates by indicating flexibility and availability to act on interventions. The impact of messages follows weekly cycles, demonstrating maximum effectiveness during weekends and Mondays, with gradually decreasing impact through mid-week (Figure 8). The model also sets parameters $(\theta,\lambda)=(0.0.4,0.01)$ to capture peer influence and the interaction between treatment and peer effects.
Figure 7: Distribution of baseline exercise probabilities across the population.
Figure 8: Distribution of intervention message effects across the population
Figure 9: Distribution of node degrees in the Twitter social network.
6.6 Data Center Server Utilization
This environment simulates a server farm, and the goal is to evaluate interventions for improving server utilization. Given the increasing demand for cloud computing resources, optimizing data center utilization has become critical for addressing global sustainability concerns (Zhang et al., 2023, Saxena et al., 2023). Within this system, servers influence each other’s performance through the system’s physical characteristics, particularly via the join-the-shortest-queue routing policy (Gupta et al., 2007). This operational dynamic creates an implicit interference pattern in the absence of a pre-specified network structure.
The experimental units are individual servers within a parallel processing system of N servers. The outcome variable $Y^{i}_{t}{}$ represents server i’s utilization during the interval [t,t+1), measured as the proportion of time the server remains busy. The treatment consists of interventions that enhance the processing power of selected servers.
The system evolution follows a structured routing mechanism. When tasks arrive, the system identifies capable servers for each task type and selects a random sample among them. Following the join-the-shortest-queue policy, tasks are assigned to servers with minimal queue lengths within this sample, using random assignment to resolve ties. This routing approach naturally creates interference effects, as performance improvements in treated servers influence task distribution across the entire system.
The environment incorporates realistic workload patterns through a time-dependent Poisson arrival process. The demand model captures multiple temporal patterns: daily variations (night-time lows, morning increases, midday peaks, and evening declines), weekly cycles (heightened weekday activity), and stochastic elements (random fluctuations and event-driven spikes). Each server processes tasks with exponentially distributed service times, which can be modified by interventions. Through this design, the system replicates key characteristics of real-world data centers while enabling controlled experimentation.
Figure 10: Average demand for the data center over time shows a strong seasonality.
7 Results
Figures 11-18 present results from applying counterfactual cross-validation (Algorithm 2) across six benchmark scenarios detailed in §6. Below, we outline our implementation approach and key findings.
For the LLM-based social network model, we conducted 10 distinct runs, constrained by OpenAI API limitations; the current simulation, with $N=1000$ and $T=30$, required approximately 100,000 GPT-3.5 API calls to generate experimental and ground truth results. Each run employs a unique treatment allocation following a staggered rollout design across three stages with $\vec{p}=(0.2,0.5,0.8)$, each spanning 10 periods. This design implies that on average 20% of units received the intervention in the initial 10 periods, followed by an additional 30% in the subsequent 10 periods, and so forth.
For the remaining five experiments, we conducted 100 independent runs for each setting, utilizing a fresh treatment allocation for each run through a staggered rollout design. The design comprises four equal-length stages with treatment probabilities $\vec{p}=(0,0.2,0.4,0.6)$. In each figure’s leftmost panel, we display the temporal evolution of outcomes through their mean and standard deviation, along with the 95th percentile across runs.
Figure 11: LLM-based social network with $N=1,000$ agents.
Figure 12: Belief adoption model with Krupina network with $N=3,366$ users.
Figure 13: Belief adoption model with Topolcany network with $N=18,246$ users.
Figure 14: Belief adoption model with Zilina network with $N=42,971$ users.
Figure 15: Ascending auction model with $N=500$ objects.
The second panels of Figures 11-18 display the box plot of the average total treatment effect (TTE) across multiple time periods. The TTE contrasts the counterfactual of all units under treatment against all units under control:
$\text{TTE}:=\frac{1}{LN}\sum_{t=T-L+1}^{T}\sum_{i=1}^{N}\left[Y^{i}_{t}{(\mathbf{1})}-Y^{i}_{t}{(\mathbf{0})}\right].$ (14)
In each setting, we carefully select $L$ so that the TTE in (14) covers all periods with nonzero treatment probability, ensuring our benchmark estimators remain meaningful. The results compare ground truth (GT) values against estimates obtained from three methods: our proposed causal message-passing approach (CMP), the difference-in-means estimator (DM), and the Horvitz-Thompson estimator (HT) [^8] Difference-in-means (DM) and Horvitz-Thompson (HT) are expressed as: $\hat{\tau}_{\text{DM}}:=\frac{1}{L}\sum_{t=T-L+1}^{T}\Big{(}\frac{\sum_{i=1}^{N}Y^{i}_{t}{}W^{i}_{t}}{\sum_{i=1}^{N}W^{i}_{t}}-\frac{\sum_{i=1}^{N}Y^{i}_{t}{}(1-W^{i}_{t})}{\sum_{i=1}^{N}(1-W^{i}_{t})}\Big{)},\quad\quad\hat{\tau}_{\text{HT}}:=\frac{1}{LN}\sum_{t=T-L+1}^{T}\sum_{i=1}^{N}\left(\frac{Y^{i}_{t}{}W^{i}_{t}}{\mathbb{E}[W^{i}_{t}]}-\frac{Y^{i}_{t}{}(1-W^{i}_{t})}{\mathbb{E}[1-W^{i}_{t}]}\right).$ (Sävje et al., 2021). Finally, the rightmost two panels in each figure display the CFE under the ground truth and CMP estimates for all-control and all-treatment conditions, along with their respective 95th percentiles.
In implementing Algorithm 2, we employ five validation batches ($b_{v}=5$). To select candidate estimators, we begin with a base model where each outcome is expressed as a linear function of two components: the sample mean of outcomes from the previous round and the current treatment allocation means. We then systematically modify this model by incorporating additional first-order and higher-order terms. The configurations also span batch counts from 200 to 2000 and batch sizes ranging from 0.1 to 20 percent of the population size. To estimate parameters, we employ Ridge regression with penalty parameters logarithmically spaced from $10^{-4}$ to $10^{4}$. These parameters are comprehensively combined to generate a diverse set of potential estimators, with time blocks aligned to experimental stages. For example, when $T=40$ and the design $\vec{p}=(0,0.2,0.4,0.6)$ with equal length blocks is used, $\mathfrak{L}_{\mathsf{tb}}$ has four elements, one corresponding to each block with a fixed treatment probability. Then, the selection process incorporates both domain knowledge and observed data characteristics. For instance, the pronounced temporal patterns evident in the left panels of Figures 16-18, observed in the New York City taxi model, exercise encouragement program, and data center model, necessitate estimators with detrending steps (see Remark 5.2). Computational efficiency is maintained by constraining the estimator search space based on the experimental context.
Overall, our framework demonstrates robust performance across all six scenarios, successfully estimating counterfactual evolutions despite strong seasonality patterns and without requiring information about the underlying interference network. The proposed method achieves significantly better performance than both DM and HT estimators, even in settings with subtle treatment effects. As illustrated in Figures 11-16, CMP yields estimates with both smaller bias and variance in different scenarios. The effectiveness of our method is particularly evident in the challenging scenarios presented in Figures 12-15 and 18, where conventional estimators struggle to reliably determine even the direction of treatment effects. These comprehensive results establish our framework’s capability to deliver precise estimates of counterfactual evolutions and treatment effects across diverse experimental settings.
Figure 16: New York City Taxi model with $N=18,768$ Routes.
Figure 17: Exercise encouragement program with $N=30,162$ users.
Figure 18: Data Center model with $N=1,000$ servers.
Remark 7.1
Selecting a predetermined number of batches for a given batch size $n^{\mathcal{S}}$ presents a significant computational challenge, particularly in large-scale problems with time-varying treatment allocations across units. For staggered rollout designs, we implement a heuristic approach while deferring comprehensive analysis to future research. Our heuristic consists of three steps. First, we order units by their treatment duration, defined as the number of time periods under treatment. Second, we select two blocks of size $n^{\mathcal{S}}$—one that slides through the ordered list to cover all treatment durations, and another chosen randomly to ensure sufficient between-batch variation. Third, we select individual units from these merged blocks with equal probability to generate batches with average size $n^{\mathcal{S}}$. This procedure maintains computational efficiency while ensuring batches with diverse treatment allocations with high probability.
8 Concluding Remarks
This work presents several contributions to the study of causal inference under network interference: a distribution-preserving network bootstrap approach that enables resampling from networked populations with unobserved interaction patterns, a counterfactual cross-validation framework to validate estimation methods, and a benchmark toolbox containing six experimental environments. Our empirical results suggest the framework can be effective across diverse settings, showing promising performance compared to existing approaches when handling complex temporal patterns and unobserved network structures. Future work exploring advanced machine learning architectures and incorporating domain-specific knowledge may further enhance the framework’s practical utility while maintaining its theoretical properties.
Appendices Organization
These appendices provide detailed statements and proofs of theoretical results that support the main body of the paper.
The first appendix presents the technical results, beginning with essential notation and a reformulation of the outcome specification. We then establish the outcome decomposition rules through a sequence of proofs. This begins with Lemmas 9.1-9.3, culminating in the proof of Theorem 4.4 and Corollary 4.7. The discussion then progresses to a rigorous statement of batch-level state evolution in §9.4, which was informally introduced in Theorem 4.2. This is followed by a brief overview of the conditioning technique in §9.5, which is essential for proving the state evolution equation, as detailed in §9.6.
Appendix 10 demonstrates how consistent estimation of state evolution parameters enables consistent estimation of desired counterfactuals. We present the necessary assumptions, provide and prove the main theoretical results, and detail the corresponding algorithms. We then analyze the application of these results to Bernoulli randomized designs, concluding in presenting two families of estimators in §10.2.1 and §10.2.2.
Building on these results, §11 presents an extension to the causal message-passing methodology that addresses strong time-trends, enabling general counterfactual estimation even in the presence of seasonality or temporal patterns. The appendices conclude with §12, which presents auxiliary theorems necessary for our main proofs.
9 Technical Results
In this section, we delve into the analysis of the outcome specification in Eq. (3). Our analysis draws upon various results from the literature on approximate message-passing algorithms (Donoho et al., 2009, Bayati and Montanari, 2011, Rush and Venkataramanan, 2018, Li and Wei, 2022).
In the following, we first introduce several notations necessary for the rigorous presentation of our theoretical results. We then rewrite the potential outcome specification, facilitating the ensuing discussions. Next, we provide a rigorous proof for the outcome decomposition rule. This is rooted in the non-asymptotic results for the analysis of AMP algorithms (Li and Wei, 2022) and can be seen as the finite sample analysis of the causal message-passing framework (Shirani and Bayati, 2024), within a broader class of outcome specifications.
9.1 Notations
For any vector $\vec{v}\in\mathbb{R}^{n}$, we denote its Euclidean norm as ${\left\|\vec{v}\right\|}$. For a fixed $k\geq 1$, we define ${{\cal C}{\cal P}}(k)$ as the class of functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that are continuous and exhibit polynomial growth of order $k$. In other words, there exists a constant $c$ such that $|f(\vec{v})|\leq c(1+{\left\|\vec{v}\right\|}^{k})$. Moreover, we consider a probability space $(\Omega,\mathbb{F},\mathbb{P})$, where $\Omega$ represents the sample space, $\mathbb{F}$ is the sigma-algebra of events, and $\mathbb{P}$ is the probability measure. We denote the expectation with respect to $\mathbb{P}$ as $\mathbb{E}$. Additionally, for any other probability measure $p$, we use $\mathbb{E}_{p}$ to denote the expectation with respect to $p$.
For any set $S$, the indicator function $\mathds{1}_{S}(\omega)$ evaluates to $1$ if $\omega$ belongs to $S$, and 00 otherwise. We define $\mathbb{R}^{n\times m}$ as the set of matrices with $n$ rows and $m$ columns. Given a matrix $\bm{M}$, we denote its transpose as $\bm{M}^{\top}$. Additionally, we represent a matrix of ones with dimensions $n\times m$ as $\bm{1}_{n\times m}\in\mathbb{R}^{n\times m}$. The symbol =$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}$ is used to denote equality in distribution, while =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;$ is used for equalities that hold almost surely with respect to the reference probability measure $\mathbb{P}$.
9.2 Preliminaries
We initiate by rewriting the outcome model in Eq. (3). Specifically, we consider the following more general specification:
$\vec{Y}_{t+1}{}(\bm{W}{})=\vec{\Upsilon}_{t+1}{}+{\bf M}_{t}g_{t}\left(\vec{Y}_{0}{}(\bm{W}{}),\ldots,\vec{Y}_{t}{}(\bm{W}{}),\bm{W}{},\bm{X}\right)+h_{t}{}\left(\vec{Y}_{0}{}(\bm{W}{}),\ldots,\vec{Y}_{t}{}(\bm{W}{}),\bm{W}{},\bm{X}\right)+\vec{\epsilon}^{\;}_{t},$ (15)
where
$\vec{\Upsilon}_{t+1}{}=\big{(}{\bf A}+{\bf B}_{t}\big{)}g_{t}\left(\vec{Y}_{0}{}(\bm{W}{}),\ldots,\vec{Y}_{t}{}(\bm{W}{}),\bm{W}{},\bm{X}\right).$ (16)
Above, ${\bf M}_{t}$ is the matrix with the element $(\mu^{ij}+\mu_{t}^{ij})/N$ in the $i^{th}$ row and $j^{th}$ column. With a slight abuse of notation, we consider the entries of matrices ${\bf A}$ and ${\bf B}_{t}$ to have zero mean. Formally, we state the following assumption regarding the interference matrices. {assumption} Entries of ${\bf A}$ are i.i.d. Gaussian variables with zero mean and variance $\sigma^{2}/N$. Similarly, for any $t$, entries of ${\bf B}_{t}$ are i.i.d. Gaussian variables with zero mean and variance $\sigma^{2}_{t}/N$, independent of other components of the model.
Compared to (3), the specification in (15) and (16) is more comprehensive, incorporating both the complete outcome history and the full treatment allocation matrix. This generalization enables us to capture more complex temporal dynamics, including additional lag terms and potential anticipation effects of treatments. The functions $g_{t}$ and $h_{t}{}$ can also be specified to include only a finite number $l\in\mathbb{N}$ of historical lag terms. While all subsequent results hold for the general model in (15) and (16), for notational simplicity, we present the proofs using only one lag term $\vec{Y}_{t}{}(\bm{W}{})$. Furthermore, when the context is clear, we omit the treatment matrix notation $\bm{W}{}$ and write simply $\vec{Y}_{t}{}$ as the vector of outcomes at time $t$.
9.3 Outcome Decomposition Rule
Fixing $N$, we proceed by setting more notations. Given $\vec{Y}_{0}{}$ as the vector of initial outcomes, for $1\leq t<N$, define:
$\vec{V}_{0}{}:=\frac{g_{0}\left(\vec{Y}_{0}{},\bm{W}{}{},\bm{X}\right)}{{\left\|g_{0}\left(\vec{Y}_{0}{},\bm{W}{}{},\bm{X}\right)\right\|}},\quad\quad\quad\quad\vec{V}_{t}{}:=\frac{\left({\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}\right)g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)}{{\left\|\left({\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}\right)g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)\right\|}},$ (17)
where, ${\bf I}_{N}$ is $N\times N$ identity matrix, and
$\bm{V}_{t-1}{}=\left[\vec{V}_{0}{}\Big{|}\ldots\Big{|}\vec{V}_{t-1}{}\right].$ (18)
Note that ${\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}$ functions as a projection onto the subspace that is orthogonal to the column space of $\bm{V}_{t-1}{}$. As a result, the vectors $\{\vec{V}_{0}{},\ldots,\vec{V}_{N-1}{}\}$ constitute an orthonormal basis by definition. Therefore, we can represent the vector $g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)$ with respect to this basis as $\vec{\alpha}_{t}:=(\alpha^{0}_{t},\ldots,\alpha^{t}_{t},0,\ldots,0)^{\top}\in\mathbb{R}^{N}$; that is:
$g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)=\sum_{j=0}^{t}\alpha^{j}_{t}\vec{V}_{j}{},\quad\quad\quad\quad\alpha^{j}_{t}=\left<g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right),\vec{V}_{j}{}\right>.$ (19)
Then, it is immediate to get ${\left\|\vec{\alpha}_{t}\right\|}={\|g_{t}(\vec{Y}_{t}{},\bm{W}{}{},\bm{X})\|}$. In addition, we let $\bm{V}^{\perp}_{t-1}{}\in\mathbb{R}^{N\times(N-t)}$ denote the orthogonal complement of $\bm{V}_{t-1}{}$ such that $(\bm{V}^{\perp}_{t-1}{})^{\top}\bm{V}^{\perp}_{t-1}{}={\bf I}_{N-t}$.
We also define the following sequence of matrices based on the fixed interference matrix ${\bf A}$:
${\bf A}_{0}:={\bf A},\quad\quad\quad\quad{\bf A}_{t}:={\bf A}_{t-1}\left({\bf I}_{N}-\vec{V}_{t-1}{}\vec{V}^{\top}_{t-1}{}\right).$ (20)
Then, Eq. (20) enables us to write:
${\bf A}_{0}={\bf A}_{t}+\sum_{j=0}^{t-1}\left({\bf A}_{j}-{\bf A}_{j+1}\right)={\bf A}_{t}+\sum_{j=0}^{t-1}{\bf A}_{j}\vec{V}_{j}{}\vec{V}^{\top}_{j}{}.$ (21)
Further, for $t=1,\ldots,N$, we define the following sequence of matrices:
$\widetilde{\bf A}_{t}:={\bf A}_{t}\bm{V}^{\perp}_{t-1}{}={\bf A}_{t-1}\left({\bf I}_{N}-\vec{V}_{t-1}{}\vec{V}^{\top}_{t-1}{}\right)\bm{V}^{\perp}_{t-1}{}={\bf A}_{t-1}\bm{V}^{\perp}_{t-1}{}=\ldots={\bf A}\bm{V}^{\perp}_{t-1}{}\in\mathbb{R}^{N\times(N-t)},$ (22)
where we used the fact that $\vec{V}^{\top}_{s-1}{}\bm{V}^{\perp}_{t-1}{}=\vec{0}\in\mathbb{R}^{N-t}$, for any $s\leq t$. Also, by (20), we can write
${\bf A}_{t}={\bf A}_{t-1}\left({\bf I}_{N}-\vec{V}_{t-1}{}\vec{V}^{\top}_{t-1}{}\right)=\ldots={\bf A}\left({\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}\right)={\bf A}\bm{V}^{\perp}_{t-1}{}(\bm{V}^{\perp}_{t-1}{})^{\top}=\widetilde{\bf A}_{t}(\bm{V}^{\perp}_{t-1}{})^{\top}.$ (23)
We prove the outcome decomposition rule in multiple steps, beginning with the following lemma.
Lemma 9.1
For any $t=0,\ldots,N-1$, we have
$\vec{Y}_{t+1}{}={\bf M}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\sum_{j=0}^{t}\alpha^{j}_{t}\left({\bf A}_{j}+{\bf B}_{t}\right)\vec{V}_{j}{}+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\vec{\epsilon}^{\;}_{t}.$ (24)
Proof. By outcome model given in (15) and (21), we can write:
$\vec{Y}_{t+1}{}$ $={\bf A}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)+\sum_{j=0}^{t-1}{\bf A}_{j}\vec{V}_{j}{}\vec{V}^{\top}_{j}{}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)+\big{(}{\bf M}_{t}+{\bf B}_{t}\big{)}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\vec{\epsilon}^{\;}_{t}$
$={\bf A}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)+\sum_{j=0}^{t-1}\alpha^{j}_{t}{\bf A}_{j}\vec{V}_{j}{}+\sum_{j=0}^{t}\alpha^{j}_{t}{\bf B}_{t}\vec{V}_{j}{}+{\bf M}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\vec{\epsilon}^{\;}_{t}$
$={\bf M}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\sum_{j=0}^{t}\alpha^{j}_{t}({\bf A}_{j}+{\bf B}_{t})\vec{V}_{j}{}+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\vec{\epsilon}^{\;}_{t},$
where in the third line, we used (19) and the fact that the vectors $\{\vec{V}_{0}{},\ldots,\vec{V}_{t}{}\}$ constitute an orthonormal set. Also, in the last line, we utilized the following fact
${\bf A}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)={\bf A}_{t}\left({\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}\right)g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)=\alpha^{t}_{t}{\bf A}_{t}\vec{V}_{t}{},$
that holds true because ${\bf A}_{t}\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}=0$ by (20); additionally, the term $\left({\bf I}_{N}-\bm{V}_{t-1}{}\bm{V}^{\top}_{t-1}{}\right)g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)$ is equal to the projection of $g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right)$ on the subspace perpendicular to the column space of $\bm{V}_{t-1}{}$, which is $\alpha^{t}_{t}\vec{V}_{t}{}$. It completes the proof. $\Box$
The next lemma characterizes the distribution of $\widetilde{\bf A}_{t}$, defined in (22), based on the rotational invariance property of Gaussian matrices.
Lemma 9.2
Fix $1\leq t<N$. Conditional on $\vec{Y}_{0}{}$, $\vec{Y}_{1}{},\ldots,\vec{Y}_{t-1}{}$, $\bm{W}{}{}$, and $\bm{X}$, entries of the matrix $\widetilde{\bf A}_{t}\in\mathbb{R}^{N\times(N-t)}$ are i.i.d. with distribution ${\cal N}(0,{\sigma^{2}}/{N})$, and the matrix $\widetilde{\bf A}_{t}$ (and so ${\bf A}_{t}$) is independent of $\vec{Y}_{t}{},\vec{V}_{t}{},$ as well as ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{t-1}\vec{V}_{t-1}{}$.
Proof. First, note that given $\bm{W}{}{}$ and $\bm{X}$, by (17), conditioning on $\vec{Y}_{0}{}$, $\vec{Y}_{1}{},\ldots,\vec{Y}_{t-1}{}$, is equivalent to conditioning on $\vec{V}_{0}{}$, $\vec{V}_{1}{},\ldots,\vec{V}_{t-1}{}$. Now, we use an induction on $t$ to prove the result.
Step 1. Let $t=1$. By (22) and the rotational invariance property of Gaussian matrices, we have
$\widetilde{\bf A}_{1}={\bf A}\bm{V}^{\perp}_{0}{}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}{\bf A}\vec{\text{e}}^{\perp}_{1},$ (25)
where $\vec{\text{e}}^{\perp}_{1}$ denotes the orthogonal complement of $\vec{\text{e}}_{1}$, which is the first standard basis vector. Letting $\vec{\text{e}}^{\perp}_{1}=[\vec{\text{e}}_{2}|\ldots|\vec{\text{e}}_{N}]$, by Assumption 9.2, we get that the entries of $\widetilde{\bf A}_{1}\in\mathbb{R}^{N\times(N-1)}$ are i.i.d. with distribution ${\cal N}(0,{\sigma^{2}}/{N})$. Furthermore, considering that $\bm{V}_{0}{}=\vec{V}_{0}{}$, the matrix $\widetilde{\bf A}_{1}$ is independent from ${\bf A}_{0}\vec{V}_{0}{}$. Then, conditional on $\vec{Y}_{0}{}$, the outcome model in (15) implies that $\widetilde{\bf A}_{1}$ is independent of $\vec{Y}_{1}{}$ and $\vec{V}_{1}{}$ (note that the randomness of $\vec{Y}_{1}{}$ comes from ${\bf A}_{0}\vec{V}_{0}{}$, ${\bf B}_{0}$, and $\vec{\epsilon}^{\;}_{0}$). Finally, considering (23), the same results about the independency hold true for the matrix ${\bf A}_{1}$.
Step 2. Suppose that the result is true for $t=1,\ldots,s-1$. Given $\vec{Y}_{0}{},\ldots,\vec{Y}_{s-1}{}$, we show the result also holds for $t=s$. Note that by the induction hypothesis, conditional on $\vec{Y}_{0}{},\ldots,\vec{Y}_{s-2}{}$, the matrix ${\bf A}_{s-1}$ is independent of $\vec{Y}_{s-1}{}$ and ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{s-2}\vec{V}_{s-2}{}$. Thus, ${\bf A}_{s-1}$ and $\vec{Y}_{s-1}{}$ are conditionally independent. This implies that conditional on $\vec{Y}_{0}{},\ldots,\vec{Y}_{s-1}{}$ (we added $\vec{Y}_{s-1}{}$), the matrix ${\bf A}_{s-1}$ (and so ${\bf A}_{s}:={\bf A}_{s-1}({\bf I}_{N}-\vec{V}_{s-1}{}\vec{V}^{\top}_{s-1}{})$, see (20), as well as $\widetilde{\bf A}_{s}$) is still independent of ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{s-2}\vec{V}_{s-2}{}$.
Next, we show that ${\bf A}_{s}$ is also independent from ${\bf A}_{s-1}\vec{V}_{s-1}{}$. By (22) and the rotational invariance property of Gaussian matrices, we can write
$\widetilde{\bf A}_{s}={\bf A}\bm{V}^{\perp}_{s-1}{}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}{\bf A}[\vec{\text{e}}_{1}|\ldots|\vec{\text{e}}_{s}]^{\perp}.$
Here, $[\vec{\text{e}}_{1}|\ldots|\vec{\text{e}}_{s}]^{\perp}$ represents the orthogonal complement of the first $s$ standard basis vectors. Then, a similar argument to the one in Step 1 implies that the matrix $\widetilde{\bf A}_{s}\in\mathbb{R}^{N\times(N-s)}$ has i.i.d. entries with a distribution of ${\cal N}(0,{\sigma^{2}}/{N})$. Furthermore, it yields that $\widetilde{\bf A}_{s}$ (and consequently ${\bf A}_{s}$, see Eq. (23)) is independent of the vector ${\bf A}_{s-1}\vec{V}_{s-1}{}={\bf A}({\bf I}_{N}-\bm{V}_{s-2}{}\bm{V}^{\top}_{s-2}{})\vec{V}_{s-1}{}={\bf A}\vec{V}_{s-1}{}$; this holds true because of Eq. (23) and the fact that $\vec{V}^{\top}_{j}{}\vec{V}_{s-1}{}=0$, for $j=0,\ldots,s-2$.
Now, by Lemma 9.1, we have
$\vec{Y}_{s}{}={\bf M}_{s-1}g_{s-1}\left(\vec{Y}_{s-1}{},\bm{W}{}{},\bm{X}{}\right)\sum_{j=0}^{s-1}\alpha^{j}_{s-1}\left({\bf A}_{j}+{\bf B}_{s-1}\right)\vec{V}_{j}{}+h_{s-1}{}\left(\vec{Y}_{s-1}{},\bm{W}{}{},\bm{X}{}\right)+\vec{\epsilon}^{\;}_{s-1}.$
As a result, ${\bf A}_{s}$ and $\widetilde{\bf A}_{s}$ are also independent of $\vec{Y}_{s}{}$ and $\vec{V}_{s}{}$. This concludes the proof. $\Box$
By combining the results of Lemmas 9.1 and 9.2, we arrive at the conclusion of Lemma 9.3.
Lemma 9.3
For any $t=0,\ldots,N-1$, we have
$\vec{Y}_{t+1}{}={\bf M}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|g_{t}(\vec{Y}_{t}{},\bm{W}{}{},\bm{X})\right\|}\sum_{j=0}^{t}\beta^{j}_{t}\vec{Z}_{j}+\vec{\epsilon}^{\;}_{t},$ (26)
where $\vec{Z}_{0},\vec{Z}_{1},\ldots,\vec{Z}_{t}$ are i.i.d. random vectors in $\mathbb{R}^{N}$ following ${\cal N}(0,\frac{1}{N}{\bf I}_{N})$ distribution. Additionally, $\beta^{j}_{t}:=\alpha^{j}_{t}/{\|g_{t}(\vec{Y}_{t}{},\bm{W}{},\bm{X})\|}$, making $\vec{\beta}_{t}=(\beta^{0}_{t},\ldots,\beta^{t}_{t},0,\ldots,0)^{\top}\in\mathbb{R}^{N}$ a unit random vector (i.e., ${\|\vec{\beta}_{t}\|}=1$). Note that $\beta^{j}_{t}$ and $\vec{Z}_{j}$ are not independent.
Proof. Fixing $t$, we prove the result in two steps. First, we demonstrate that ${\bf A}_{j}\vec{V}_{j}{}$, for $j=0,\ldots,t$, follows a Gaussian distribution with specified mean and variance. Then, we show that ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{t}\vec{V}_{t}{}$ are independent.
Conditional on $\vec{Y}_{0}{}$, $\vec{Y}_{1}{},\ldots,\vec{Y}_{j-1}{}$, $\bm{W}{}{}$, and $\bm{X}$, Lemma 9.2 implies that the matrix ${\bf A}_{j}$ and the vector $\vec{V}_{j}{}$ are independent. Also, by (23), we can write
${\bf A}_{j}\vec{V}_{j}{}={\bf A}\bm{V}^{\perp}_{j-1}{}(\bm{V}^{\perp}_{j-1}{})^{\top}\vec{V}_{j}{}={\bf A}\vec{V}_{j}{},$
where we used the fact that the vector $\vec{V}_{j}{}$ is perpendicular to the column space of the matrix $\bm{V}_{j-1}{}$. Thus, conditional on the value of $\vec{V}_{j}{}$, as well as $\vec{Y}_{0}{}$, $\vec{Y}_{1}{},\ldots,\vec{Y}_{j-1}{}$, $\bm{W}{}{}$, and $\bm{X}$, by the rotational invariance property of Gaussian matrices, the elements of ${\bf A}_{j}\vec{V}_{j}{}$ are i.i.d. random variables with distribution ${\cal N}(0,\frac{\sigma^{2}}{N})$. Furthermore, note that this conditional distribution of ${\bf A}_{j}\vec{V}_{j}{}$ remains the same regardless of the value of $\vec{V}_{j}{}$. As a result, we can conclude that the elements of ${\bf A}_{j}\vec{V}_{j}{}$ are i.i.d. Gaussian, even without conditioning on $\vec{V}_{j}{}$ as well as $\vec{Y}_{0}{}$, $\vec{Y}_{1}{},\ldots,\vec{Y}_{j-1}{}$, $\bm{W}{}{}$, and $\bm{X}$. Precisely, for a deterministic vector $\vec{v}$, we can write:
$\mathbb{P}\left({\bf A}_{j}\vec{V}_{j}{}\leq\vec{v}\right)=\mathbb{E}\left[\mathbb{E}\left[\mathbb{P}\left({\bf A}_{j}\vec{V}_{j}{}\leq\vec{v}\right)\Big{|}\vec{V}_{j}{}\right]\bigg{|}\vec{Y}_{0}{},\ldots,\vec{Y}_{j-1}{},\bm{W}{}{},\bm{X}\right]=\mathbb{E}\left[\mathbb{E}\left[\Phi(\vec{v})\Big{|}\vec{V}_{j}{}\right]\bigg{|}\vec{Y}_{0}{},\ldots,\vec{Y}_{j-1}{},\bm{W}{}{},\bm{X}\right]=\Phi(\vec{v}),$
where $\Phi$ denotes the CDF of a vector whose entries follow a normal distribution ${\cal N}(0,\frac{\sigma^{2}}{N})$.
We proceed by establishing the independence property. Note that by Lemma 9.2, conditional on the values of $\vec{Y}_{0}{},\ldots,\vec{Y}_{t}{},\bm{W}{}{},\bm{X}$ (and so on the values of $\vec{V}_{0}{},\ldots,\vec{V}_{t-1}{},\vec{V}_{t}{}$), it follows that ${\bf A}_{t}$ (and so ${\bf A}_{t}\vec{V}_{t}{}$) is independent of ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{t-1}\vec{V}_{t-1}{}$. Importantly, our previous demonstration confirmed that the distribution of ${\bf A}_{t}\vec{V}_{t}{}$ remains unchanged across different values of $\vec{Y}_{0}{},\ldots,\vec{Y}_{t}{},\bm{W}{}{},\bm{X}$. Consequently, we can assert that the random vector ${\bf A}_{t}\vec{V}_{t}{}$ is independent of ${\bf A}_{0}\vec{V}_{0}{},\ldots,{\bf A}_{t-1}\vec{V}_{t-1}{}$. More precisely, we can repeat this argument multiple times and, for deterministic vectors $\vec{v}_{0},\ldots,\vec{v_{t}}$, show that
$\mathbb{P}\left({\bf A}_{0}\vec{V}_{0}{}\leq\vec{v}_{0},\ldots,{\bf A}_{t}\vec{V}_{t}{}\leq\vec{v}_{t}\right)$ $=\mathbb{E}\left[\mathbb{P}\left({\bf A}_{0}\vec{V}_{0}{}\leq\vec{v}_{0},\ldots,{\bf A}_{t}\vec{V}_{t}{}\leq\vec{v}_{t}\right)\Big{|}\vec{Y}_{0}{},\ldots,\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right]$
$=\mathbb{E}\left[\mathbb{P}\left({\bf A}_{0}\vec{V}_{0}{}\leq\vec{v}_{0},\ldots,{\bf A}_{t-1}\vec{V}_{t-1}{}\leq\vec{v}_{t-1}\right)\Phi(\vec{v}_{t})\Big{|}\vec{Y}_{0}{},\ldots,\vec{Y}_{t}{},\bm{W}{}{},\bm{X}\right]$
$=\Phi(\vec{v}_{t})\mathbb{E}\left[\mathbb{P}\left({\bf A}_{0}\vec{V}_{0}{}\leq\vec{v}_{0},\ldots,{\bf A}_{t-1}\vec{V}_{t-1}{}\leq\vec{v}_{t-1}\right)\Big{|}\vec{Y}_{0}{},\ldots,\vec{Y}_{t-1}{},\bm{W}{}{},\bm{X}\right]$
=$\ldots$= …
$=\Phi(\vec{v}_{0})\Phi(\vec{v}_{1})\ldots\Phi(\vec{v}_{t}).$
Based on the fact that the time-dependent interference matrix ${\bf B}_{t}$ and the noise vector $\vec{\epsilon}^{\;}_{t}$ are independent of everything else in the model, the proof is complete and we obtain the desired result in (26). $\Box$
The dependence between $\beta^{j}_{t}$ and $\vec{Z}_{j}$ in Lemma 9.3 implies that the Gaussianity of the elements in $\vec{Y}_{t+1}{}$ cannot be inferred directly from the result of this lemma alone.
Proof of Theorem 4.4. To obtain the desired result, we apply Lemma 9.3 and Lemma 12.3 together. To be more specific, in view of Lemma 9.3, we know that
$\vec{Y}_{t+1}{}={\bf M}_{t}g_{t}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+h_{t}{}\left(\vec{Y}_{t}{},\bm{W}{}{},\bm{X}{}\right)+\sqrt{\sigma^{2}+\sigma_{t}^{2}}{\left\|g_{t}(\vec{Y}_{t}{},\bm{W}{}{},\bm{X})\right\|}\vec{R}_{t}+\vec{\epsilon}^{\;}_{t},$
where $\vec{R}_{t}=\sum_{i=0}^{t}\beta^{i}_{t}\vec{Z}_{i}$. But, by Lemma 12.3, we have
$W_{1}\left(\mathcal{L}\left(\vec{R}_{t}\right),{\cal N}\left(0,\frac{1}{N}{\bf I}_{N}\right)\right)\leq c\sqrt{\frac{t\log N}{N}},$
which concludes the proof. $\Box$
Proof of Corollary 4.7. By the result of Theorem 4.4, we can write
$\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t+1}{}=\;$ $\frac{1}{N\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\sum_{j=1}^{N}\left(\mu^{ij}+\mu_{t}^{ij}\right)g_{t}\left(Y^{j}_{t}{},\vec{W}^{j}{},{\vec{X}}^{j}\right)+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}h_{t}{}\left(Y^{i}_{t}{},\vec{W}^{i}{},{\vec{X}}^{i}\right)$
$+\sqrt{\frac{\sigma^{2}+\sigma_{t}^{2}}{\left|{\mathcal{S}}\right|}}{\left\|g_{t}(\vec{Y}_{t}{},\bm{W}{}{},\bm{X})\right\|}\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{i\in\mathcal{S}}R^{i}_{t}+\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\epsilon^{i}_{t}.$
Letting $J_{t}:=\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{i\in\mathcal{S}}R^{i}_{t}$ and applying Lemma 12.4 for $\vec{\Phi}=\vec{R}_{t}$, we get the result. $\Box$
9.4 Batch-level State Evolution
Next, we analyze the large-sample behavior of the outcomes for a subpopulation of units. Specifically, let $\mathcal{S}\subset[N]$ represent an arbitrary subpopulation, with its size $\left|{\mathcal{S}}\right|$ increasing indefinitely as the population size $N$ grows large to infinity. We investigate the asymptotic behavior of the elements in $\bm{Y}{}$ as $N$ approaches infinity, This provides valuable insights into the evolution of outcomes within the subpopulation. {assumption} Fixing $T\in\mathbb{N}$ and $k\geq 2$, we assume that
  1. (i)
    For all $t\in[T]$, the function $g_{t}:\mathbb{R}^{1+T+M}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(\frac{k}{2})$ function.
  2. (ii)
    For all $t\in[T]$, the function $h_{t}{}:\mathbb{R}^{1+T+M}\rightarrow\mathbb{R}$ is a ${{\cal C}{\cal P}}(1)$ function.
  3. (iii)
    The sequence of initial outcome vectors $\vec{Y}_{0}{}$, the treatment allocation matrices $\bm{W}{}$, the covariates $\bm{X}$, and the function $g_{0}$ are such that for a deterministic value $\tilde{\rho}_{1}{}>0$, we have
    $(\tilde{\rho}_{1}{})^{2}$ $=\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{0}^{2}}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}^{2}<\infty.$
Assumption 9.4 comprises a collection of regularity conditions on the model attributes. The first two parts ensure that the functions $g_{t}$ and $h_{t}{}$ do not demonstrate fast explosive behavior, thereby ensuring the well-posedness of the large system asymptotic. The final part pertains to the initial observation of the network and corroborates that the function $g_{0}$ is non-degenerate. This ensures that initial observations provide meaningful information and contribute to the evolution of outcomes. {assumption} Fix $T\in\mathbb{N}$, $k\geq 2$, and a subpopulation $\mathcal{S}\subset[N]$, where the size $\left|{\mathcal{S}}\right|$ grows to infinity as $N\rightarrow\infty$. We assume the following statements hold:
  1. (i)
    $p_{y}^{\mathcal{S}}(N)$ denotes the empirical distribution of the initial outcomes $Y^{i}_{0}{(N)}$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{y}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{y}^{\mathcal{S}}}\big{[}{\|Y_{0}{}\|}^{k}\big{]}<\infty$,
  2. (ii)
    $p_{x}^{\mathcal{S}}(N)$ denotes the empirical distribution of the covariate vectors ${\vec{X}}^{i}(N)$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{x}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{x}^{\mathcal{S}}}\big{[}{\|{\vec{X}}\|}^{k}\big{]}<\infty$,
  3. (iii)
    $p_{w}^{\mathcal{S}}(N)$ denotes the empirical distribution of the treatment vectors $\vec{W}^{i}_{N}$ with $i\in\mathcal{S}$, and converges weakly to a probability measure $p_{w}^{\mathcal{S}}$ such that $\mathbb{E}_{p_{w}^{\mathcal{S}}}\big{[}{\|\vec{W}\|}^{k}\big{]}<\infty$.
  4. (iv)
    For all $i$ and any $t\in[T]_{0}$, let $p_{\mu^{i}}(N)$ and $p_{\mu_{t}^{i}}(N)$ be, respectively, the empirical distribution of the elements of the vectors $\vec{\mu}^{\;i\cdot}:=(\mu^{i1},\ldots,\mu^{iN})^{\top}$ and $\vec{\mu}^{\;i\cdot}_{t}:=(\mu^{i1}_{t},\ldots,\mu^{iN}_{t})^{\top}$. Then, $p_{\mu^{i}}(N)$ converges weakly to $p_{\mu^{i}}$ and $p_{\mu_{t}^{i}}(N)$ converges weakly to $p_{\mu_{t}^{i}}$. Also, $\mathbb{E}_{p_{\mu^{i}}}\big{[}{\|{M}^{i}\|}^{k}\big{]}<\infty$ and $\mathbb{E}_{p_{\mu^{i}_{t}}}\big{[}{\|{M}_{t}^{i}\|}^{k}\big{]}<\infty$.
    In addition, the limit distributions ($p_{\mu^{i}}$ and $p_{\mu^{i}_{t}}$) are independent of other randomnesses in the model and if $\bar{\mu}^{i}$ denotes the mean of a random variable under probability measure $p_{\mu^{i}}$ (i.e., $\bar{\mu}^{i}=\mathbb{E}_{p_{\mu^{i}}}[{M}^{i}]$), the empirical distributions of $\bar{\mu}^{i},\;i\in\mathcal{S}$, denoted by $p_{\mu}^{\mathcal{S}}(N)$, converges weakly to a probability measure $p_{\mu}^{\mathcal{S}}$. Likewise, we let $p_{\mu_{t}}^{\mathcal{S}}$ denote the weak limit of the empirical distribution of the means under probability measures $p_{\mu^{i}_{t}}$. Finally, $\mathbb{E}_{p_{\mu}^{\mathcal{S}}}\big{[}{\|{M}^{\mathcal{S}}\|}^{k}\big{]}<\infty$ as well as $\mathbb{E}_{p_{\mu_{t}}^{\mathcal{S}}}\big{[}{\|{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}<\infty$.
  5. (v)
    For all $i$ and $t\in[T]_{0}$, as $N\rightarrow\infty$, we have
    $\;\mathbb{E}_{p_{y}^{\mathcal{S}}(N)\times p_{x}^{\mathcal{S}}(N)\times p_{w}^{\mathcal{S}}(N)\times p_{\mu^{i}}(N)\times p_{\mu_{t}^{i}}(N)\times p_{\mu}^{\mathcal{S}}(N)\times p_{\mu_{t}}^{\mathcal{S}}(N)}\big{[}{\|Y_{0}{},{\vec{X}},\vec{W},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}$
    $\rightarrow$ $\;\mathbb{E}_{p_{y}^{\mathcal{S}}\times p_{x}^{\mathcal{S}}\times p_{w}^{\mathcal{S}}\times p_{\mu^{i}}\times p_{\mu^{i}_{t}}\times p_{\mu}^{\mathcal{S}}\times p_{\mu_{t}}^{\mathcal{S}}}\big{[}{\|Y_{0}{},{\vec{X}},\vec{W},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}.$
Remark 9.4
We can replace the second part of Assumption 9.4-(iv) by assuming that $p_{\mu^{i}}$ is the same across all units, and similarly, $p_{\mu^{i}_{t}}$ is identical for all units.
Remark 9.5
We can drop Assumptions 9.4 and 9.4-(v) by confining the functions $g_{t}$ and $h_{t}{}$ to be bounded and continuous.
Remark 9.6 (Notation convention)
In Assumption 9.4, when $\mathcal{S}$ represents the entire experimental population, we omit the superscript $\mathcal{S}$ from all notations.
Remark 9.7 (Time-dependent covariates)
For each time $t=0,1,\ldots,T$, let ${\cal X}_{t}\in\mathbb{R}^{n^{\text{x}}_{t}\times N}$ denote the time-dependent covariate matrix, where the $i^{th}$ column corresponds to the covariates for unit $i$ at time $t$. We extend the covariate matrix $\bm{X}$ by incorporating these time-dependent covariates, resulting in a new $(n^{\text{x}}+n^{\text{x}}_{0}+\ldots+n^{\text{x}}_{T})\times N$ covariate matrix in our model. The functions $g_{t}$ and $h_{t}{}$ are then modified accordingly to reference the appropriate portion of this extended covariate matrix at each time period. Along this, we exclude the noise vectors (i.e., $\vec{\epsilon}^{\;}_{t}$) from subsequent discussions on the potential outcome specification Eq. (15).
Assumption 9.4 is a standard assumption in statistical theory, ensuring that the empirical distributions of system attributes remain stable and do not diverge as the sample size increases. This assumption holds, for example, when units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i}$ follow an i.i.d. distribution with finite moments of order $k$. Moreover, a wide range of treatment assignments satisfies the conditions of Assumption 9.4, including cases where the support of $\pi$ is bounded, such as the Bernoulli design. By imposing such conditions, we ensure the reliability of estimation by keeping the experimental design moments finite and manageable.
To state the main theoretical results, we need to define the Batch State Evolution (BSE) equations as follows:
$\tilde{\nu}_{1}^{\mathcal{S}}{}$ $=\mathbb{E}\left[({M}^{\mathcal{S}}+{M}_{0}^{\mathcal{S}})g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}\right],\quad\tilde{\rho}_{1}{}=(\sigma+\sigma_{0})\mathbb{E}\left[g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}^{2}\right],\quad H^{\mathcal{S}}_{0}{}=h_{0}{}\big{(}Y^{\mathcal{S}}_{0}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)},$ (27)
$H^{\mathcal{S}}_{t}{}$ $=h_{t}{}\big{(}\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)},$
$\tilde{\nu}_{t+1}^{\mathcal{S}}{}$ $=\mathbb{E}\left[({M}^{\mathcal{S}}+{M}_{t}^{\mathcal{S}})g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right],$
$(\tilde{\rho}_{t+1}{})^{2}$ $=(\sigma^{2}+\sigma_{t}^{2})\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\vec{W},{\vec{X}}\big{)}^{2}\right],$
$\nu_{t+1}^{\mathcal{S}}{}$ $=\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\mathbb{E}\left[H^{\mathcal{S}}_{t}{}\right],$
$(\rho_{t+1}{})^{2}$ $=(\tilde{\rho}_{t+1}{})^{2}+\mathrm{Var}\left[H_{t}{}\right].$
where
  • $Y_{0}{}\sim p_{y}$ and $Y^{\mathcal{S}}_{0}{}\sim p_{y}^{\mathcal{S}}$ represent the weak limits of the population and subpopulation initial outcomes;
  • $\vec{W}\sim p_{w}$ and $\vec{W}^{\mathcal{S}}\sim p_{w}^{\mathcal{S}}$ are the weak limits of the population and subpopulation treatment assignments;
  • ${\vec{X}}\sim p_{x}$ and ${\vec{X}}^{\mathcal{S}}\sim p_{x}^{\mathcal{S}}$ represent the weak limits of the population and subpopulation covariates;
  • ${M}\sim p_{{M}}$ and ${M}_{t}\sim p_{{M}_{t}}$ represent the weak limits of interference elements at the population level, as specified in Assumption 9.4. Similarly, the corresponding subpopulation quantities are denoted by ${M}^{\mathcal{S}}\sim p_{{M}}^{\mathcal{S}}$ and ${M}_{t}^{\mathcal{S}}\sim p_{{M}_{t}}^{\mathcal{S}}$;
  • and independent from all of them, $Z_{t}$ follows a standard Gaussian distribution.
The following theorem characterizes the distribution of units’ outcomes $Y^{i}_{1}{},\ldots,Y^{i}_{t+1}{}$ within the large sample asymptotic, based on the BSE equations outlined in Eq. (27).
Theorem 9.8
Fixing $k\geq 2$, consider the sequence of units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i,t}$ and suppose that Assumptions 3.3-9.4 hold. Then, in view of the BSE equations given in Eq. (27), for any function $\psi\in{{\cal C}{\cal P}}(k)$, we have,
$\;\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{T}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$ (28)
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{T}^{\mathcal{S}}{}+\tilde{\rho}_{T}{}Z_{T}+H^{\mathcal{S}}_{T-1}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]},$
where $Z_{t}\sim{\cal N}(0,1),\;t=1,\ldots,T,$ independent of $\vec{W}^{\mathcal{S}}\sim p_{w}^{\mathcal{S}}$ and ${\vec{X}}^{\mathcal{S}}\sim p_{x}^{\mathcal{S}}$. In addition, for any $t=1,\ldots,T$, the random variables $Z_{t}$ and $H_{t-1}{}$ are independent.
We prove the result of Theorem 9.8 by extending the theoretical results of Shirani and Bayati, (2024), which in turn build on the AMP framework developed by Bayati and Montanari, (2011). The proof mainly relies on a conditioning technique introduced by Bolthausen, (2014). Below, we first present a version of the conditioning technique adapted to our specific setting, followed by a detailed proof of the theorem.
Remark 9.9
To derive the result in the second part of Theorem 4.2 from Theorem 9.8, we note that when $\mathcal{S}$ depends solely on the treatment allocation, which is independently distributed from all other variables, $\mathcal{S}$ effectively functions as a random sample from the experimental population. As a result, the distributions of $Y^{\mathcal{S}}_{0}{}$, ${\vec{X}}^{\mathcal{S}}$, ${M}^{\mathcal{S}}$, and ${M}_{t}^{\mathcal{S}}$ are equivalent to those of $Y_{0}{}$, ${\vec{X}}$, ${M}$, and ${M}_{t}$, respectively. This equivalence follows from the fact that the empirical distribution of quantities in a random sample converges to the empirical distribution of the entire population.
9.5 Conditioning Technique
Recalling (16) and letting $\vec{U}_{t}{}=g_{t}\big{(}\vec{Y}_{t}{},\bm{W}{},\bm{X}\big{)}$ (and $U^{i}_{t}{}=g_{t}\big{(}Y^{i}_{t}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$), we denote
$\bm{Q}_{t}:=\left[\vec{U}_{0}{}\Big{|}\vec{U}_{1}{}\Big{|}\ldots\Big{|}\vec{U}_{t-1}{}\right],\quad\quad\bm{R}_{t}:=\left[\vec{\Upsilon}_{1}{}-{\bf B}_{0}\vec{U}_{0}{}\Big{|}\ldots\Big{|}\vec{\Upsilon}_{t}{}-{\bf B}_{t-1}\vec{U}_{t-1}{}\right].$ (29)
According to Eq. (29), $\bm{Q}_{t}$ and $\bm{R}_{t}$ are matrices with columns of $\vec{U}_{s-1}{}$ and $\vec{\Upsilon}_{s}{}-{\bf B}_{s-1}\vec{U}_{s-1}{}$, when $s=1,\ldots,t$, respectively. Then, we denote by $\vec{U}^{\parallel}_{t}{}$ the projection of $\vec{U}_{t}{}$ onto the space generated by the columns of $\bm{Q}_{t}$ and define $\vec{U}^{\perp}_{t}{}=\vec{U}_{t}{}-\vec{U}^{\parallel}_{t}{}$. We also define $\vec{\gamma}_{t}=(\gamma^{0}_{t},\gamma^{1}_{t},\ldots,\gamma^{t-1}_{t})^{\top}$ such that
$\vec{U}^{\parallel}_{t}{}=\sum_{s=0}^{t-1}\gamma^{s}_{t}\vec{U}_{s}{}=\sum_{s=0}^{t-1}\gamma^{s}_{t}g_{s}\big{(}\vec{Y}_{s}{},\bm{W}{},\bm{X}\big{)},$ (30)
where
$\vec{\gamma}_{t}=\left(\bm{Q}_{t}^{\top}\bm{Q}_{t}\right)^{-1}\bm{Q}_{t}^{\top}\vec{U}_{t}{}.$ (31)
Now, note that the available observation at any time $t$ implicitly reveals information about the fixed interference matrix ${\bf A}$. To manage this intricate randomness, we define ${\cal G}_{t}$ as the $\sigma$-algebra generated by $\vec{Y}_{0}{},\vec{Y}_{1}{},\ldots,\vec{Y}_{t}{}$, $\vec{\Upsilon}_{1}{},\ldots,\vec{\Upsilon}_{t}{}$, ${\bf M}_{t}$, ${\bf B}_{0},\ldots,{\bf B}_{t-1}$, $\bm{W}{}$, and $\bm{X}$. We then compute the conditional distribution of ${\bf A}$ given ${\cal G}_{t}$. In this framework, conditioning on ${\cal G}_{t}$ is equivalent to conditioning on the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$. When conditioned on ${\cal G}_{t}$, the entries of both $\bm{Q}_{t}$ and $\bm{R}_{t}$ become deterministic values, leading to the following lemma.
Lemma 9.10
Fix $t$ and assume that $\bm{Q}_{t}$ is a full-row rank matrix. Then, for the conditional distribution of the fixed interference matrix ${\bf A}$ given ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$, we have
${\bf A}|_{{\bf A}\bm{Q}_{t}=\bm{R}_{t}}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}\bm{R}_{t}\left(\bm{Q}_{t}^{\top}\bm{Q}_{t}\right)^{-1}\bm{Q}_{t}^{\top}+\widetilde{{\bf A}}P^{\perp}.$ (32)
where $\widetilde{{\bf A}}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}{\bf A}$ independent of ${\bf A}$ and $P^{\perp}=({\bf I}-P)$ that P denotes the orthogonal projector onto the column space of $\bm{Q}_{t}$.
The proof of Lemma 9.10 relies on the rotational invariance of the Gaussian distribution and utilizes Lemma 11 from Bayati and Montanari, (2011). The proofs of this lemma and the subsequent lemma— which describes the distribution of $\vec{\Upsilon}_{t+1}{}$ given the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$— follow a similar approach to the proofs of Lemmas 2 and 3 in Shirani and Bayati, (2024), and we refer readers to that work for detailed derivations.
Lemma 9.11
Fix $t$ and assume that $\bm{Q}_{t}$ is a full-row rank matrix. The following holds for the conditional distribution of the vector $\vec{\Upsilon}_{t+1}{}$:
$\vec{\Upsilon}_{t+1}{}\big{|}_{{\cal G}_{t}}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}$ $\;\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{},$ (33)
where the matrix $\widetilde{{\bf A}}$ is independent of ${\bf A}$ and has the same distribution.
9.6 Detailed Proof of Theorem 9.8
Here, we first state Lemma 9.12 which is an expanded version of Theorem 9.8. To this end, we need some new notations. Specifically, for vectors $\vec{u},\vec{v}\in\mathbb{R}^{m}$, we define the scalar product $\vec{u},\vec{v}\right>:=\frac{1}{m}\sum_{i=1}^{m}u_{i}v_{i}$. Also, considering (27), for $t\geq 1$, we define
$(\bar{\rho}_{1}{})^{2}$ $=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}^{2}<\infty$ (34)
$(\bar{\rho}_{t+1}{})^{2}$ $=\sigma^{2}\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\vec{W},{\vec{X}}\big{)}^{2}\right].$
Lemma 9.12
For a fixed $k\geq 2$ and a specified subpopulation of experimental units $\mathcal{S}$, consider the following conditions. Given that Assumption 9.4 holds, and both the subpopulation $\mathcal{S}$ and the complete experimental population satisfy Assumption 9.4, the following statements are valid for all time steps $t$ under the BSE equations defined in (27):
  1. (a)
    For any function $\psi\in{{\cal C}{\cal P}}(k)$, we have
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$ (35)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]},$
    where $Z_{1},\ldots,Z_{t}$ are standard Gaussian random variables.
  2. (b)
    For all $0\leq r\neq s\leq t$, the following equations hold and all limits exist, are bounded, and have degenerate distribution (i.e. they are constant random variables)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (36a)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (36b)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (36c)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (36d)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (36e)
  3. (c)
    Letting $\bm{\Upsilon}_{t}=[\vec{\Upsilon}_{1}{}|\ldots|\vec{\Upsilon}_{t}{}]$, the following matrices are positive definite almost surely:
    $\lim_{N\rightarrow\infty}\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\succ 0,\quad\quad\quad\lim_{N\rightarrow\infty}\frac{\bm{\Upsilon}_{t}^{\top}\bm{\Upsilon}_{t}}{N}\succ 0.$ (37)
In the following section, we will provide a comprehensive explanation of the conditioning technique, which will be employed to establish the results presented in Lemma 9.12.
Proof. For all $t$, we assume, without loss of generality, that the mapping $y\mapsto g_{t}(y,\vec{W},{\vec{X}})$ is a non-constant function with positive probability with respect to the randomness of $\vec{W}$ and ${\vec{X}}$; otherwise, the result is trivial and does not require further analysis. We prove the result by induction on $t$.
Step 1. Let $t=0$. By definition, the matrices $\bm{Q}_{0}$ and $\bm{R}_{0}$ are empty, and the $\sigma$-algebra ${\cal G}_{0}$ is generated by $\vec{Y}_{0}{}$, $\bm{W}{}$, and $\bm{X}$. As the induction base case, we establish Parts (a) and (b) for $t=0$ and Part (c) for $t=1$.
  1. (a)
    Conditioning on the values of $\vec{Y}_{0}{}$, $\bm{W}{}$, $\bm{X}$, and so on the value of $\vec{U}_{0}{}=g_{0}\big{(}\vec{Y}_{0}{},\bm{W}{},\bm{X}\big{)}$, the elements of $\vec{\Upsilon}_{1}{}$ are i.i.d. Gaussian random variables with zero mean and variance $(\tilde{\rho}_{1N}{})^{2}$:
    $(\tilde{\rho}_{1N}{})^{2}$ $:=\mathrm{Var}\left[\Upsilon^{i}_{1}{}\Big{|}\vec{U}_{0}{}\right]=\frac{\sigma^{2}+\sigma^{2}_{0}}{N}\sum_{i=1}^{N}g_{0}\left(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\right)^{2}.$ (38)
    Then, Assumption 9.4-(iii) implies that the value of $(\tilde{\rho}_{1N}{})^{2}$ is bounded independent from $N$, and
    $(\tilde{\rho}_{1}{})^{2}:=\lim_{N\rightarrow\infty}(\tilde{\rho}_{1N}{})^{2}.$ (39)
    Now, let $Z$ denote a standard Gaussian random variable. Fixing $l\geq 1$, it is straightforward to show that
    $\mathbb{E}\left[\big{|}\Upsilon^{i}_{1}{}\big{|}^{l}\big{|}\vec{U}_{0}{}\right]$ $=\mathbb{E}\left[\big{|}\tilde{\rho}_{1N}{}Z\big{|}^{l}\big{|}\vec{U}_{0}{}\right]\leq c,$ (40)
    where $c$ is a constant independent of $N$ and might alter in different lines.
    We next focus on the second term on the right-hand side of Eq. (15) and define
    $\tilde{\nu}_{1N}^{i}{}$ $:=\frac{1}{N}\sum_{j=1}^{N}(\mu^{ij}+\mu_{0}^{ij})g_{0}\big{(}Y^{j}_{0}{},\vec{W}^{j},{\vec{X}}^{j}\big{)},\quad\quad\quad i\in[N].$ (41)
    In view of Assumption 9.4, we can apply Theorem 12.2. Consequently, for all $i$, we can write
    $\lim_{N\rightarrow\infty}\tilde{\nu}_{1N}^{i}{}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[({M}^{i}+{M}_{0}^{i})g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}\right]=(\bar{\mu}^{i}+\bar{\mu}^{i}_{0})\mathbb{E}\left[g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}\right]=\tilde{\nu}_{1}^{i}{}<\infty.$ (42)
    Above, we let $\bar{\mu}^{i}:=\mathbb{E}[{M}^{i}]$ and $\bar{\mu}^{i}_{0}:=\mathbb{E}[{M}_{0}^{i}]$, where ${M}^{i}\sim p_{\mu^{i}}$ and ${M}_{0}^{i}\sim p_{\mu^{i}_{0}}$. Additionally, $Y_{0}{}$, $\vec{W}$, and ${\vec{X}}$ represent the weak limits of the initial outcomes, treatment allocations, and covariates for the entire population as outlined in Assumption 9.4, respectively. In this context, $\bar{\mu}^{i}+\bar{\mu}^{i}_{0}$ determines the average interaction level of unit $i$ at time 00.
    Also, note that (42) yields the boundedness of $\tilde{\nu}_{1N}^{i}{}$ for all $i$ independent of $N$. Now, using Assmption 9.4 and the fact that $\psi\in{{\cal C}{\cal P}}(k)$, for $\kappa>0$, we have the following,
    $\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}\left[\Big{|}\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}-\mathbb{E}_{{\bf A},{\bf B}_{0}}\left[\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}\right]\Big{|}^{2+\kappa}\right]\leq c\left|{\mathcal{S}}\right|^{\kappa/2},$
    where $\mathbb{E}_{{\bf A},{\bf B}_{0}}$ is the expectation with respect to the randomness of the interference matrices ${\bf A},{\bf B}_{0}$ and $c$ is a constant independent of $N$. Then, applying the Strong Law of Large Numbers (SLLN) for triangular arrays in Theorem 12.1, we obtain the following result:
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\Big{(}\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}-\mathbb{E}_{{\bf A},{\bf B}_{0}}\left[\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}\right]\Big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;0.$ (43)
    On the other hand, employing the dominated convergence theorem, e.g., Theorem 16.4 in Billingsley, (2008), allows us to interchange the limit and the expectation in view of the fact that $\psi\in{{\cal C}{\cal P}}(k)$. We also utilize the continuous mapping theorem, e.g., Theorem 2.3 in Van der Vaart, (2000), to pass the limit through the function. As a result, considering $\Upsilon^{i}_{1}{}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}\tilde{\rho}_{1N}{}Z$, we get
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{{\bf A},{\bf B}_{0}}\left[\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}\right]\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{Z}\left[\psi\big{(}Y^{i}_{0}{},\tilde{\rho}_{1}{}Z,\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{}\big{)}\right].$ (44)
    Then, applying Theorem 12.2 for the function $f(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{})=\mathbb{E}_{Z}\left[\psi\big{(}Y^{i}_{0}{},\tilde{\rho}_{1}{}Z,\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{}\big{)}\right]$, we can write
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\rho}_{1}{}Z,\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},\tilde{\nu}_{1}^{\mathcal{S}}{}\big{)}\Big{]},$ (45)
    where $Y^{\mathcal{S}}_{0}{},\;\vec{W}^{\mathcal{S}},\;{\vec{X}}^{\mathcal{S}},\;\tilde{\nu}_{1}^{\mathcal{S}}{}$ represent the weak limits of $Y^{i}_{0}{},\;\vec{W}^{i},\;{\vec{X}}^{i},\;\tilde{\nu}_{1}^{i}{}$ over the subpopulation units, as specified in Assumption 9.4 and
    $\tilde{\nu}_{1}^{\mathcal{S}}{}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[({M}^{\mathcal{S}}+{M}_{0}^{\mathcal{S}})g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}\right],$ (46)
    where ${M}^{\mathcal{S}}+{M}_{0}^{\mathcal{S}}$ captures the weak limit of the average interference level for units in the subpopulation.
    In Eq. (45), the function $f$ is within ${{\cal C}{\cal P}}(k)$, since $\psi\in{{\cal C}{\cal P}}(k)$ and expectation is a linear operator. It is important to note that above, $Z$ is independent of $Y^{\mathcal{S}}_{0}{},\;\vec{W}^{\mathcal{S}},\;{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}}$, and ${M}_{0}^{\mathcal{S}}$. This is true because the randomness of $Z$ arises from the interference matrices which are assumed to be independent of everything in the model, see Assumption 9.2.
    Now, we use Eq. (45) to derive the main result. Fix an arbitrary function $\psi\in{{\cal C}{\cal P}}(k)$ and based on Eqs. (15) and (16), define the function $\widetilde{\psi}$ such that
    $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}=\psi\left(Y^{i}_{0}{},\Upsilon^{i}_{1}{}+\tilde{\nu}_{1N}^{i}{}+h_{0}{}\big{(}Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\big{)},\vec{W}^{i},{\vec{X}}^{i}\right)=\widetilde{\psi}\left(Y^{i}_{0}{},\Upsilon^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1N}^{i}{}\right).$ (47)
    The function $\widetilde{\psi}$ is within ${{\cal C}{\cal P}}(k)$ by Assumption 9.4. Then, applying (45) for the function $\widetilde{\psi}$, we obtain,
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\psi\left(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z+H^{\mathcal{S}}_{0}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\right)\right]$ (48)
    where $H^{\mathcal{S}}_{0}{}=h_{0}{}\big{(}Y^{\mathcal{S}}_{0}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}$ is a random variable independent of $Z$.
    In the second step of the induction, we also require the following results. Note that the result in (48) represents a specific instance of the more comprehensive result (50). These results can be derived by following the same procedure outlined above.
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}$ $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\Upsilon^{i}_{1}{},\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{},\vec{W}^{i},{\vec{X}}^{i},\mu^{ji},\mu^{ji}_{r}\big{)}$ (49)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z+H_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big{)}\Big{]},$
    as well as
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}$ $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\Upsilon^{i}_{1}{},\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{},\vec{W}^{i},{\vec{X}}^{i},\bar{\mu}^{i},\bar{\mu}^{i}_{r}\big{)}$ (50)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z+H^{\mathcal{S}}_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big{)}\Big{]},$
    for any $j$ and all $r\in[T]_{0}$; here, $Z^{\prime}$ is a standard Gaussian random variable and $\vec{\text{B}}^{i\cdot}_{0}$ detnoes the $i^{th}$ row of the time-dependent interference matrix ${\bf B}_{0}$.
  2. (b)
    By (39) and (49), we get
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{0}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (51)
    Similarly, by (34) and (49), we can write
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (52)
    Finally, given $\vec{U}_{0}{}$, we know that $\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}$ and $\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}$ are statistically independent. Then, applying (49) together with (52), we obtain the following result:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{1}{}\big{(}\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}+\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}\big{(}\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{n=1}^{N}\left(U^{n}_{0}{}\right)^{2}.$
    where $\vec{\text{A}}^{i\cdot}$ and $\vec{\text{B}}^{i\cdot}_{0}$ denote the $i^{th}$ row of the fixed interference matrix ${\bf A}$ and time-dependent interference matrix ${\bf B}$.
  3. (c)
    For $t=1$, the matrix $\bm{Q}_{1}$ is equal to the vector $\vec{U}_{0}{}$ and $\bm{V}_{1}$ is equal to the vector $\vec{Y}_{1}{}$. By Assumption 9.4-(iii) and (51), we have
    $\lim_{N\rightarrow\infty}\frac{\bm{Q}_{1}^{\top}\bm{Q}_{1}}{N}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}>0,\quad\lim_{N\rightarrow\infty}\frac{\bm{\Upsilon}_{1}^{\top}\bm{\Upsilon}_{1}}{N}=\lim_{N\rightarrow\infty}\left<\vec{\Upsilon}_{1}{},\vec{\Upsilon}_{1}{}\right>>0.$
Induction Hypothesis (IH). Now, we assume that the following results hold true:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}$ $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{1}{},\ldots,\Upsilon^{i}_{t}{},\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{},\ldots,\Upsilon^{i}_{t}{}-\vec{\text{B}}^{i\cdot}_{t-1}\vec{U}_{t}{},\vec{W}^{i},{\vec{X}}^{i},\mu^{ji},\mu^{ji}_{r}\big{)}$ (IH-1)
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z_{1}+H_{0}{},\ldots,\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big{)}\Big{]},$
and
$\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}$ $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{1}{},\ldots,\Upsilon^{i}_{t}{},\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{},\ldots,\Upsilon^{i}_{t}{}-\vec{\text{B}}^{i\cdot}_{t-1}\vec{U}_{t}{},\vec{W}^{i},{\vec{X}}^{i},\bar{\mu}^{i},\bar{\mu}^{i}_{r}\big{)}$ (IH-2)
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big{)}\Big{]}.$
Also, for $10\leq s\neq r\leq t-$ 1, we have
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (IH-3)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-4)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (IH-5)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-6)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (IH-7)
Finally, the following condition holds almost surely:
$\lim_{N\rightarrow\infty}\frac{\bm{\Upsilon}_{t-1}^{\top}\bm{\Upsilon}_{t-1}}{N}\succ 0.$ (IH-8)
Step 2. To establish the second step of the induction, we prove the assertions in reverse order, starting with Part (c), followed by Part (b), and concluding with Part (a).
  1. (c)
    We begin the second step by applying (IH-1) to the function $g_{s}\big{(}Y^{i}_{s}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}g_{r}\big{(}Y^{i}_{r}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$, for $1\leq r,s\leq t$. Precisely, by Assumption 9.4-(iii) as well as (IH-1), we get
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(U^{i}_{0}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})^{2}\big{)}\right]=\frac{(\tilde{\rho}_{1}{})^{2}}{\sigma^{2}+\sigma_{0}^{2}}>0,$ (53)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{0}{}U^{i}_{s}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}\right],$
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}g_{r}\big{(}\tilde{\nu}_{r}{}+\tilde{\rho}_{r}{}Z_{r}+H_{r-1}{},\vec{W},{\vec{X}}\big{)}\Big{]}.$
    Now, let $\vec{u}=(u_{1},\ldots,u_{t})^{\top}\in\mathbb{R}^{t}$ be a non-zero vector. By Assumption 9.4-(iii) and (53), we have
    $\vec{u}^{\top}\left(\lim_{N\rightarrow\infty}\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\right)\vec{u}$ $=\lim_{N\rightarrow\infty}\vec{u}^{\top}\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\vec{u}$ (54)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\left(u_{1}g_{0}(Y_{0}{},\vec{W},{\vec{X}})+\sum_{s=1}^{t-1}u_{s+1}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}\right)^{2}\right]\geq 0.$
    We show that the inequality in (54) is strict. To this end, note that $\vec{u}$ is a non-zero vector, and there exists some $1\leq i\leq t$ such that $u_{i}\neq 0$. Whenever $u_{1}\neq 0=u_{2}=\ldots=u_{t}$, the result is immediate. Otherwise, recall that $y\mapsto g_{s}(y,\vec{W},{\vec{X}})$ is a non-constant function with a positive probability with respect to $(\vec{W},{\vec{X}})$; consequently, the mapping $(y_{0},\ldots,y_{t-1})\mapsto\sum_{s=0}^{t-1}u_{s}g_{s}\big{(}y_{s},\vec{W},{\vec{X}}\big{)}$ is a non-constant function as well. Considering $H_{s}{}=h_{s}{}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}$ and $H_{0}{}=h_{0}{}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}$, the randomness of $u_{1}g_{0}(Y_{0}{},\vec{W},{\vec{X}})+\sum_{s=1}^{t-1}u_{s+1}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}$ comes solely from $Y_{0}{},Z_{1},\ldots,Z_{t-1},\vec{W},{\vec{X}}$; as a result, there exists a measurable continuous function $\phi$ (which depends on $\vec{u},g_{0},\ldots,g_{t-1},h_{0}{},\ldots,h_{t-1}{}$) such that we can rewrite the right-hand side of (54) as follows,
    $\;\mathbb{E}\left[\left(u_{1}g_{0}(Y_{0}{},\vec{W},{\vec{X}})+\sum_{s=1}^{t-1}u_{s+1}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}\right)^{2}\right]$
    $=$ $\;\mathbb{E}\left[\phi\left(Y_{0}{},\tilde{\nu}_{1}{},\ldots,\tilde{\nu}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t-1}{}Z_{t-1},\vec{W},{\vec{X}}\right)^{2}\right]$
    $=$ $\;\mathbb{E}\left[\mathbb{E}\left[\phi\left(y_{0},\tilde{\nu}_{1}{},\ldots,\tilde{\nu}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t-1}{}Z_{t-1},\vec{W},{\vec{X}}\right)^{2}\Bigg{|}Y_{0}{}=y_{0}\right]\right],$
    where in the last equality we used the tower property of conditional expectations. Then, it suffices to show that the random variable $\phi\left(y_{0},\tilde{\nu}_{1}{},\ldots,\tilde{\nu}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t-1}{}Z_{t-1},\vec{W},{\vec{X}}\right)$ has a non-degenerate distribution. To obtain that, by (IH-1), it is straightforward to obtain the following,
    $\mathrm{Cov}\left[\left(\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t-1}{}Z_{t-1}\right)\right]\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\bm{\Upsilon}_{t-1}^{\top}\bm{\Upsilon}_{t-1}}{N},$
    which is positive definite by (IH-8), and the proof of the first claim is complete.
    To proceed to the proof of the second claim, for $1\leq r,s\leq t$, let us denote
    $v_{r,s}:=\left[\frac{\bm{\Upsilon}_{t}^{\top}\bm{\Upsilon}_{t}}{N}\right]^{r,s}=\frac{\vec{\Upsilon}^{\top}_{r}{}\vec{\Upsilon}_{s}{}}{N}.$
    By (IH-4), whenever $r\neq s$, we can write
    $\lim_{N\rightarrow\infty}v_{r,s}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r-1}{}U^{i}_{s-1}{},$
    and if $r=s$, by (IH-3), we have
    $\lim_{N\rightarrow\infty}v_{r,r}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{r}^{2}}{N}\sum_{i=1}^{N}(U^{i}_{r-1}{})^{2}.$
    Then, the result follows directly, as we have just established the almost sure positive definiteness of $\bm{Q}_{t}$.
    Corollary 9.13
    The vector $\vec{\gamma}_{t}$ defined in (31) has a finite limit as $N\rightarrow\infty$.
    Proof. By (31), we can write
    $\lim_{N\rightarrow\infty}\vec{\gamma}_{t}=\lim_{N\rightarrow\infty}\left(\bm{Q}_{t}^{\top}\bm{Q}_{t}\right)^{-1}\bm{Q}_{t}^{\top}\vec{U}_{t}{}=\lim_{N\rightarrow\infty}\left(\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\right)^{-1}\lim_{N\rightarrow\infty}\frac{\bm{Q}_{t}^{\top}\vec{U}_{t}{}}{N}.$ (55)
    Using the result of part (c), for large values of $N$, the matrix $\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}$ is positive definite (this is true because the eigenvalues of a matrix vary continuously with respect to its entries). Then, note that the mapping $\bm{G}\mapsto\bm{G}^{-1}$ is continuous for any invertible matrix $\bm{G}$. As a result, we get
    $\lim_{N\rightarrow\infty}\left(\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\right)^{-1}=\left(\lim_{N\rightarrow\infty}\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}\right)^{-1}.$
    Since the matrix $\lim_{N\rightarrow\infty}\frac{\bm{Q}_{t}^{\top}\bm{Q}_{t}}{N}$ is positive definite, the left term in (55) is well-defined and finite. The finiteness of the right term is the consequence of (53). $\Box$
  2. (b)
    We first derive several intermediate results and then utilize them to demonstrate that (36) holds true for $0\leq r,s\leq t$. In this process, we apply the Strong Law of Large Numbers (SLLN) from Theorem 12.1 multiple times, without explicitly verifying the conditions, as they are straightforward.
    By Lemma 9.11, conditioning on ${\cal G}_{t}$, the terms $\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}$ for $i\in[N]$ are i.i.d. Gaussian random variables. Similarly, the terms $\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}$ for $i\in[N]$ are also i.i.d. Gaussian random variables:
    $\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\sim{\cal N}\left(0,\frac{\sigma^{2}}{N}\sum_{j=1}^{N}(U^{\perp,j}_{t}{})^{2}\right),\quad\quad\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\sim{\cal N}\left(0,\frac{\sigma_{t}^{2}}{N}\sum_{j=1}^{N}(U^{j}_{t}{})^{2}\right).$ (56)
    Now, applying Theorem 12.1, we get the following results:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\right)^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2},$ (57)
    as well as
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2}.$ (58)
    Also, considering (29) and applying (IH-6), we obtain
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)^{2}$ (59)
    $=$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{s=0}^{t-1}\gamma^{t}_{s}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\right)^{2}$
    $=$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\sum_{0\leq s,r<t}\gamma^{t}_{s}\gamma^{t}_{r}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}$
    $=$ $\sum_{0\leq s,r<t}\gamma^{t}_{s}\gamma^{t}_{r}\left(\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{0\leq s,r<t}\gamma^{t}_{s}\gamma^{t}_{r}U^{i}_{s}{}U^{i}_{r}{}$
    $=$ $\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\parallel,i}_{t}{}\right)^{2},$
    where in the last line we used (30).
    Now, we first obtain (36a) for $r=t$. By (33), (57), (58), and (59), we can write
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ (60)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\parallel,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}$
    $\quad+\lim_{N\rightarrow\infty}\frac{2}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)+\lim_{N\rightarrow\infty}\frac{2}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
    $\quad+\lim_{N\rightarrow\infty}\frac{2}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right).$
    Note that the only random elements in the right-hand side of (60) are $\vec{\widetilde{\text{A}}}^{i\cdot}$ and $\vec{\text{B}}^{i\cdot}_{t}$. Thus, by (56) and applying Theorem 12.1, we can demonstrate that the last three terms vanish, resulting in the following:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{t+1}{})^{2},$ (61)
    where the last equality is immediate by (IH-1).
    Next, we derive (36b) for $r=t$ and $10\leq s\leq t-$ 1. Considering (33), we can write
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\Upsilon^{i}_{t+1}{}$ (62)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\Upsilon^{i}_{s+1}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\Upsilon^{i}_{s+1}{}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\Upsilon^{i}_{s+1}{}\right).$
    Note that by conditioning on ${\cal G}_{t}$, the value of $\Upsilon^{n}_{s+1}{}$ is deterministic. Then, applying Theorem 12.1 and considering (56), (29), and by (IH-7), we get the desired result:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\Upsilon^{i}_{t+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\Upsilon^{i}_{s+1}{}$ (63)
    $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{r=0}^{t-1}\gamma^{r}_{t}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\Upsilon^{i}_{s+1}{}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\sum_{r=0}^{t-1}\gamma^{r}_{t}\left(\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\left(\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{r=0}^{t-1}\gamma^{r}_{t}U^{i}_{r}{}U^{i}_{s}{}\right)$
    $=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\parallel,i}_{t}{}U^{i}_{s}{}\right)$
    $=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}U^{i}_{s}{}\right),$
[…]
    The derivations for (36c) and (36d) follow a similar procedure, which we omit here for brevity. We then apply a similar approach to obtain (36e). Specifically, fixing $10\leq r\leq t-$ 1 and setting $s=t$, we can write the following using (33) and (63):
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}-\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\big{)}\Upsilon^{i}_{r+1}{}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\Upsilon^{i}_{r+1}{}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}U^{i}_{r}{}\right).$
    Likewise, we can show the result for the case that $r=t$ and $10\leq s\leq t-$ 1:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\Upsilon^{i}_{t+1}{}$ (64)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
    $=$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\right)\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$
    $+\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)},$
    $+\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}.$
    Then, by Theorem 12.1, the first and second terms on the right-hand side are zero. Additionally, (29), (30), and (IH-6) imply that
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\sum_{r=0}^{t-1}\gamma^{r}_{t}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ (65)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{},$
[…]
  3. (a)
    We use induction hypotheses to establish the following result:
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$ (66)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]}.$
    More general results related to the extension of (49) and (50) follow a similar procedure and are omitted for brevity.
    We proceed by introducing a new notation. Fixing $i$ as an arbitrary unit, we define
    $\Psi^{i}(N)$ $:=\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big{)},$
    where
    $\tilde{\nu}_{(t+1)N}^{i}{}$ $=\frac{1}{N}\sum_{k=1}^{N}(\mu^{ik}+\mu_{t}^{ik})g_{t}\big{(}Y^{k}_{t}{},\vec{W}^{k},{\vec{X}}^{k}\big{)}.$
    Using (IH-1), this implies that
    $\lim_{N\rightarrow\infty}\tilde{\nu}_{(t+1)N}^{i}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[({M}^{i}+{M}_{t}^{i})g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right]$ (67)
    $=(\bar{\mu}^{i}+\bar{\mu}^{i}_{t})\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right]=\tilde{\nu}_{t+1}^{i}{}<\infty.$
    Now, by (33), we can write
    $\Psi^{i}(N)\Big{|}_{{\cal G}_{t}}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize d}}}}{{=}}\psi\left(Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\right),$
    where $\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i}$ represent the $i^{th}$ element in the vector $\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}$. We also let
    $\widetilde{\Psi}^{i}(N)=\Psi^{i}(N)-\mathbb{E}_{{\bf A},{\bf B}_{t}}[\Psi^{i}(N)].$
    where $\mathbb{E}_{{\bf A},{\bf B}_{t}}$ denotes the expectation with respect to the randomness of the interference matrices ${\bf A}$ and ${\bf B}_{t}$. We follow the same approach as Step 1-(a). Note that given ${\cal G}_{t}$, the elements of $\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}$ are i.i.d. Gaussian random variables with a zero mean and variance $(\hat{\rho}_{tN})^{2}$:
    $(\hat{\rho}_{tN})^{2}$ $:=\mathrm{Var}\left[[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}]^{i}\Big{|}\vec{U}_{t}{}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}+\frac{\sigma_{t}^{2}}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2},$ (68)
    where $U^{j}_{t}{}=g_{t}\big{(}Y^{j}_{t}{},\vec{W}^{j},{\vec{X}}^{j}\big{)}$ is the $j^{th}$ element of the column vector $\vec{U}_{t}{}$, and similarly, $U^{\perp,j}_{t}{}$ is the $j^{th}$ element of the column vector $\vec{U}^{\perp}_{t}{}$. Letting
    $(\hat{\rho}_{t})^{2}=\lim_{N\rightarrow\infty}(\hat{\rho}_{tN})^{2},$ (69)
    we have the following lemma regarding the finiteness of $\hat{\rho}_{t}$.
    Lemma 9.14
    $\hat{\rho}_{t}$ is almost surely finite.
    Proof of Lemma 9.14. We show that the following relations hold with a probability of 1:
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}<\infty,\;\;\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}(U^{j}_{t}{})^{2}<\infty.$ (70)
    By definition, we can write
    $\frac{1}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}=\left<\vec{U}^{\perp}_{t}{},\vec{U}^{\perp}_{t}{}\right>$ =$\vec{U}_{t}{},\vec{U}_{t}{}\right>-\left<\vec{U}^{\parallel}_{t}{},\vec{U}^{\parallel}_{t}{}\right>=\frac{1}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2}-\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}.$ (71)
    Then, by (IH-1) for the function $g_{t}\big{(}Y^{j}_{t}{},\vec{W}^{j},{\vec{X}}^{j}\big{)}^{2}$, we get
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}g_{t}\big{(}Y^{j}_{t}{},\vec{W}^{j},{\vec{X}}^{j}\big{)}^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big{)}^{2}\right]<\infty,$ (72)
    where $Z\sim{\cal N}(0,1)$. Further, by (30), we have
    $\frac{1}{N}\sum_{j=1}^{N}U^{\parallel,j}_{t}{}$ $=\frac{1}{N}\sum_{j=1}^{N}\sum_{s=0}^{t-1}\gamma^{s}_{t}U^{j}_{s}{}=\sum_{s=0}^{t-1}\frac{\gamma^{s}_{t}}{N}\sum_{j=1}^{N}U^{j}_{s}{}$
    $\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2},$ $=\frac{1}{N}\sum_{j=1}^{N}\left(\sum_{s=0}^{t-1}\gamma^{s}_{t}U^{j}_{s}{}\right)^{2}=\sum_{r,s=0}^{t-1}\gamma^{r}_{t}\gamma^{s}_{t}\left<\vec{U}_{r}{},\vec{U}_{s}{}\right>.$
    Considering Corollary 9.13, the vector $\vec{\gamma}_{t}$ has a finite limit as $N\rightarrow\infty$. Similar to (72), the induction hypothesis for the function $\psi=g_{r}\big{(}Y^{j}_{r}{},\vec{W}^{j},{\vec{X}}^{j}\big{)}g_{s}\big{(}Y^{j}_{s}{},\vec{W}^{j},{\vec{X}}^{j}\big{)}$ implies that almost surely
    $\quad\quad\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}=\lim_{N\rightarrow\infty}\sum_{r,s=0}^{t-1}\gamma^{r}_{t}\gamma^{s}_{t}\left<\vec{U}_{r}{},\vec{U}_{s}{}\right><\infty.$ (73)
    Consequently, by (71)-(73), we get the result in (70) and the proof is complete. $\Box$
    An immediate corollary of the result of Lemma 9.14 is that $\hat{\rho}_{tN}$, in (68), is almost surely bounded independent of $N$. Then, for $l\geq 1$, it is straightforward to show that,
    $\mathbb{E}\left[\left|\big{[}\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}\big{]}^{i}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right|^{l}\right]\leq 2^{l-1}\mathbb{E}\left[\left|\big{[}\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}\big{]}^{i}\right|^{l}+\left|\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right|^{l}\right]\leq c,$ (74)
    where $c$ is a constant independent of $N$ and we used the inequality $(v_{1}+v_{2})^{l}\leq 2^{l-1}(v_{1}^{l}+v_{2}^{l}),\;v_{1},v_{2}\geq 0$. Note that in (74), given ${\cal G}_{t}$, the term $\bm{R}_{t}\vec{\gamma}_{t}$ is deterministic and bounded by Corollary 9.13. Then, fixing $10<\kappa<$ 1 and using the fact that $\psi\in{{\cal C}{\cal P}}(k)$ and so $|\psi(\vec{\omega})|\leq c(1+{\left\|\vec{\omega}\right\|}^{k})$, by Assumption 9.4, we get,
    $\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}\left[\left|\widetilde{\Psi}^{i}(N)\right|^{2+\kappa}\right]\leq c\left|{\mathcal{S}}\right|^{\kappa/2}.$ (75)
    Therefore, we can apply the SLLN for triangular arrays in Theorem 12.1 to obtain the following result:
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}$ $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big{)}$ (76)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\;\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{{\bf A},{\bf B}_{t}}\Big{[}\psi$ $\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\hat{\nu}_{(t+1)N}^{i}{}\Big{)}\Big{]}.$
    Now, considering (67) and (69), we can write
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{{\bf A},{\bf B}_{t}}\Big{[}\psi\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\Big{)}\Big{]}$ (77)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{Z}\Big{[}\psi\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big{)}\Big{]},$
    where similar to (44) in the first step, we utilized the dominated convergence theorem and the continuous mapping theorem to pass the limit through the expectation and the function, respectively.
    From Eq. (67), recall that $\tilde{\nu}_{t+1}^{i}{}=(\bar{\mu}^{i}+\bar{\mu}^{i}_{t})\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right]$; accordingly, we define
    $\widehat{\psi}$ $\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{},\ldots,\Upsilon^{i}_{t}{}-\vec{\text{B}}^{i\cdot}_{t-1}\vec{U}_{t-1}{},\vec{W}^{i},{\vec{X}}^{i},\bar{\mu}^{i},\bar{\mu}^{i}_{t}\big{)}$
    $:=\mathbb{E}_{Z}\Big{[}\psi$ $\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big{)}\Big{]}.$
    Considering (IH-2) as well as (76)-(77), for the function $\widehat{\psi}$, we have
    $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big{)}$ (78)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg{[}\widehat{\psi}\bigg{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\bigg{)}\bigg{]}$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg{[}\psi\bigg{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},\tilde{\nu}_{t+1}^{\mathcal{S}}{}\bigg{)}\bigg{]},$
    where $Z$ is a standard Gaussian random variable, independent of all other variables, as the inherent randomness arises from $\widetilde{{\bf A}}$ and ${\bf B}_{t}$. Additionally, $\widehat{\psi}$ belongs to ${{\cal C}{\cal P}}(k)$, since $\tilde{\nu}_{t+1}^{i}{}$ in the calculations of (78) can be viewed as a linear function depending solely on $\bar{\mu}^{i}$ and $\bar{\mu}^{i}_{t}$.
    Now, we need to show that
    $\mathrm{Var}\left[\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\big{(}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1}\big{)}\right]$ $=(\tilde{\rho}_{t+1}{})^{2}.$ (79)
    Considering that $(Z^{\prime}_{1},\ldots,Z^{\prime}_{t})$ follows a joint Normal distribution independent of $Z$, the random variable $\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1}$ is Gaussian. To obtain (79), we let $\psi=(\Upsilon^{i}_{t+1}{})^{2}$ in (78). This yields
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}\big{)}^{2}$ $=\mathbb{E}\left[\left(\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1}\right)^{2}\right].$ (80)
    Meanwhile, by (61), we have
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{t+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{t+1}{})^{2}.$ (81)
    Combining (80) and (81), we derive the desired result as stated in (79).
    Finally, similar to (47) and based on outcome representation in (15), we define the function $\widetilde{\psi}$ such that
    $\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},Y^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$ $=\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{}+\tilde{\nu}_{(t+1)N}^{i}{}+h_{t}{}\big{(}Y^{i}_{t}{},\vec{W}^{i},{\vec{X}}^{i}\big{)},\vec{W}^{i},{\vec{X}}^{i}\big{)}$
    $=\widetilde{\psi}\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big{)}.$
    Whenever $\psi\in{{\cal C}{\cal P}}(k)$, by Assumption 9.4, we get that $\widetilde{\psi}\in{{\cal C}{\cal P}}(k)$, and applying the result in Eq. (78) for the function $\widetilde{\psi}$ yields the desired result. $\Box$
10 Estimation of Counterfactual Evolutions
In this section, we present a general framework for counterfactual estimation based on the outcome specification in Eq. (3). Specifically, given a desired treatment allocation matrix $\bm{w}_{u}{}$, and observing $\bm{Y}{}(\bm{W}{}=\bm{w}_{o}{})$, $\bm{w}_{o}{}$, and $\bm{X}$, we aim to estimate the counterfactual evolution denoted by $\text{CFE}_{t}{(\bm{w}_{u}{})}$ as defined in (1).
To state the main theoretical result of this section, we parameterize the unknown functions in the state evolution equations. Specifically, let $g_{t}\big{(}\cdot;{\vec{\Theta}}_{g_{t}}\big{)}$ and $h_{t}{}\big{(}\cdot;{\vec{\Theta}}_{h_{t}{}}\big{)}$ represent the parameterized forms of $g_{t}(\cdot)$ and $h_{t}{}(\cdot)$ for $t=0,\ldots,T-1$, respectively. Here, $\widehat{\vec{\Theta}}_{g_{t}}$ and $\widehat{\vec{\Theta}}_{h_{t}{}}$ are vectors of appropriate dimensions, denoting the parameters of the respective functions. {assumption} Considering (27), for all $t$, we assume there exists a modification of the function $g_{t}$, which, with a slight abuse of notation, is also denoted by $g_{t}$ such that:
$\tilde{\nu}_{t+1}{}=\mathbb{E}\left[g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right].$ (82)
A simple example where Assumption 10 holds is when the random variable ${M}+{M}_{t}$ is independent of all other sources of randomness. In this case, we can normalize the mean to 1 by adjusting the function $g_{t}$, scaling it by an appropriate constant factor.
For further simplicity in notation, we also define $\sigma_{t}:=\sigma+\sigma_{t}$. Then, we can collect the unknown parameters in the state evolution equations in (27) as follows:
${\cal U}:=\left(\{\sigma_{t}\},\{{\vec{\Theta}}_{g_{t}}\},\{{\vec{\Theta}}_{h_{t}{}}\}\right).$ (83)
We denote an estimation of ${\cal U}$ as $\widehat{{\cal U}}:=\left(\left\{\hat{\sigma}_{t}\right\},\{\widehat{\vec{\Theta}}_{g_{t}}\},\{\widehat{\vec{\Theta}}_{h_{t}{}}\}\right)$. We can then employ Algorithm 3 to compute the desired counterfactual.
Algorithm 3 General counterfactual estimation
$\bm{Y}{}(\bm{W}{}=\bm{w}_{o}{}),\bm{w}_{o}{},\bm{X}$, $\bm{w}_{u}{}$, and $Z_{i}\sim{\cal N}(0,1)$ for $i=1,\ldots,N$
Step 1: Parameters Estimation
Estimate the set of unknown parameters ${\cal U}$ by $\widehat{{\cal U}}$.
Step 2: Counterfactual Estimation
$\widehat{\tilde{\nu}}_{1}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{0}}\big{)}$
$\widehat{\tilde{\rho}}_{1}{}\leftarrow\frac{\hat{\sigma}_{0}}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{0}}\big{)}^{2}$
$\widehat{H}^{i}_{0}{}\leftarrow h_{0}{}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{0}{}}\big{)},\;i=1,\ldots,N$
$\widehat{\text{CFE}}_{1}{(\bm{w}_{u}{})}\leftarrow\widehat{\tilde{\nu}}_{1}{}+\frac{1}{N}\sum_{i=1}^{N}\widehat{H}^{i}_{0}{}$
for $t=1,\ldots,T-1$ do
     $\widehat{H}^{i}_{t}{}\leftarrow h_{t}{}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}{}}\big{)},\;i=1,\ldots,N$
     $\widehat{\tilde{\nu}}_{t+1}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}\big{)}$
     $\widehat{\tilde{\rho}}_{t+1}{}\leftarrow\frac{\hat{\sigma}_{t}}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}\big{)}^{2}$
     $\widehat{\text{CFE}}_{t+1}{(\bm{w}_{u}{})}\leftarrow\widehat{\tilde{\nu}}_{t+1}{}+\frac{1}{N}\sum_{i=1}^{N}\widehat{H}^{i}_{t}{}$
end for
$\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$, $t=1,\ldots,T$.
{assumption}
[…]
In the following, we demonstrate the consistency of the results from Algorithm 3 under above assumptions.
Theorem 10.1 (Consistency)
Let the conditions of Lemma 9.12 and Assumptions 10-3 hold. In particular, Assumption 9.4 holds for both observed and desired treatment allocations $\bm{w}_{o}{}$ and $\bm{w}_{u}{}$. Then, for any $t$, $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ provides a consistent estimator for $\nu_{t}{(\bm{w}_{u}{})}$.
Remark 10.2
If the consistency in Assumption 3 is strong, i.e., $\widehat{{\cal U}}\xrightarrow{a.s.}{\cal U}$ as $N\rightarrow\infty$, then the consistency in Theorem 10.1 also holds strongly.
Remark 10.3
We can generalize Algorithm 3 in several ways. First, when considering outcome specifications with $l\geq 1$ lag terms, we can modify the algorithm accordingly, and extend Assumption 3 to require $l$ historical observations: $W^{i}_{t}=0$ for all $t\leq l$ and $i\in[N]$. We can also relax Assumption 3 by beginning from any arbitrary state, provided we enforce the initial conditions to the desired counterfactual scenario— specifically, requiring that $\bm{w}_{o}{}$ and $\bm{w}_{u}{}$ match for the first $l$ time periods.
Proof. We use an induction argument on $t\geq 1$ to prove the following more general statement. As $N\rightarrow\infty$, we show that
[…]
Then, it is straightforward to see that $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ also converges to $\nu_{t}{(\bm{w}_{u}{})}=\tilde{\nu}_{t}{}+\mathbb{E}\left[H_{t-1}{}\right]$ in probability, whenever $N\rightarrow\infty$.
Step 1. Let $t=1$. We begin by proving the first result in (84). To this end, we add the notation $(N)$ to the quantities associated with the system containing $N$ experimental units. We have
$\widehat{\tilde{\nu}}_{1}{(N)}$ $=\frac{1}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{0}}(N)\big{)}$ (85)
Note that, by Assumption 3, the right-hand side of (85) can be seen as a continuous function of $\widehat{\vec{\Theta}}_{g_{0}}(N)$. Therefore, applying the continuous mapping theorem, e.g., Theorem 2.3 in Van der Vaart, (2000), implies that
$\lim_{N\rightarrow\infty}\widehat{\tilde{\nu}}_{1}{(N)}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{0}}(N)\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};{\vec{\Theta}}_{g_{0}}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\tilde{\nu}_{1}{}.$ (86)
In the last equality, we used the result of Theorem 9.8. A similar argument, combined with Assumption 9.4-(iii), yields the second result in (84) for $t=1$. The third result also follows immediately.
Fixing an arbitrary function $\psi\in{{\cal C}{\cal P}}(k)$, we also need the following intermediary result:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{0}{},\vec{W}_{u},{\vec{X}}\big{)}\right],$ (87)
where $\vec{W}_{u}$ represents the weak limit of $\vec{w}^{\;i}_{u}$’s and $Z\sim{\cal N}(0,1)$. To obtain this result, let $\widetilde{\psi}$ be the function such that
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},h_{0}{}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{0}{}}(N)\big{)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$
$=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},\widehat{\vec{\Theta}}_{h_{0}{}}(N)\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},{\vec{\Theta}}_{h_{0}{}}\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\widetilde{\psi}\big{(}Z,Y_{0}{},\vec{W}_{u},{\vec{X}},{\vec{\Theta}}_{h_{0}{}}\big{)}\right].$
Above, we used the continuous mapping theorem and Theorem 12.2 in view of Assumption 9.4. Note that $\widetilde{\psi}\in{{\cal C}{\cal P}}(k)$ because of Assumption 9.4-(ii).
Induction Hypothesis (IH). Suppose that the limits in (84) hold true for $t$ and also
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{t-1}{},\vec{W}_{u},{\vec{X}}\big{)}\right].$ (88)
[…]
$\lim_{N\rightarrow\infty}\widehat{\tilde{\nu}}_{t+1}{(N)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{\tilde{\rho}}_{t}{(N)}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}(N)\big{)}$ (89)
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};{\vec{\Theta}}_{g_{t}}\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\tilde{\nu}_{t+1}{}.$
[…]
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{t}{},\vec{W}_{u},{\vec{X}}\big{)}\right].$ (90)
For this purpose, considering $\widehat{H}^{i}_{t}{(N)}=h_{t}{}\big{(}\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{\tilde{\rho}}_{t}{(N)}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}{}}(N)\big{)}$, for a proper choice of the functions $\widetilde{\psi}$ and $\widetilde{\psi}^{\prime}$, we can write:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},\widehat{\tilde{\nu}}_{t}{(N)},\widehat{\tilde{\rho}}_{t}{(N)},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i},\widehat{\vec{\Theta}}_{h_{t}{}}(N)\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},\tilde{\nu}_{t}{},\tilde{\rho}_{t}{},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i},{\vec{\Theta}}_{h_{t}{}}\big{)}$
$=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}^{\prime}\big{(}Z^{i},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\widetilde{\psi}^{\prime}\big{(}Z,H_{t-1}{},\vec{W}_{u},{\vec{X}}\big{)}\right],$
where in the last line we used the induction hypothesis. Considering the definition of functions $\widetilde{\psi}$ and $\widetilde{\psi}^{\prime}$, the proof is complete. $\Box$
10.1 Application to Bernoulli Randomized Design
We showcase the applicability of our framework by considering a Bernoulli randomized design, where each unit $i$ at time $t$ receives treatment with a probability denoted by $p_{t}$. Therefore, $W^{i}_{t}\sim Bernoulli(p_{t})$, and the $W^{i}_{t}$’s are independent across experimental units.
We consider a first-order yet non-linear approximation of functions $g_{t}$ and $h_{t}{}$ in the outcome specification in (3). Specifically, we let ${\vec{X}}^{i}=(\Delta_{g}^{i},\Theta_{gl}^{i},\ldots,\Theta_{g1}^{i},\mathcal{T}_{g}^{i},\Lambda_{g}^{i},\Delta_{h}^{i},\Theta_{hl}^{i},\ldots,\Theta_{h1}^{i},\mathcal{T}_{h}^{i},\Lambda_{h}^{i})^{\top}$ and define
$g_{t}\left(Y^{i}_{t-l+1}{},\ldots,Y^{i}_{t}{},\vec{W}^{i},{\vec{X}}^{i}\right)$ $=\Delta_{g}^{i}+\Theta_{gl}^{i}Y^{i}_{t-l+1}{}+\ldots+\Theta_{g1}^{i}Y^{i}_{t}{}+\mathcal{T}_{g}^{i}W^{i}_{t+1}+\Lambda_{g}^{i}Y^{i}_{t}{}W^{i}_{t+1}$ (91)
$h_{t}{}\left(Y^{i}_{t-l+1}{},\ldots,Y^{i}_{t}{},\vec{W}^{i},{\vec{X}}^{i}\right)$ $=\Delta_{h}^{i}+\Theta_{hl}^{i}Y^{i}_{t-l+1}{}+\ldots+\Theta_{h1}^{i}Y^{i}_{t}{}+\mathcal{T}_{h}^{i}W^{i}_{t+1}+\Lambda_{h}^{i}Y^{i}_{t}{}W^{i}_{t+1}.$
Remark 10.4
In (91), we allow the parameters of functions $g_{t}$ and $h_{t}{}$ to vary across units. These parameters can be viewed as unobserved unit-specific covariates that characterize how each unit responds to interventions. Additional unit-specific characteristics can be incorporated as observed covariates in the vectors ${\vec{X}}^{i},\;i\in[N]$. We omit these details for brevity.
We continue by considering a subpopulation of experimental units denoted by $\mathcal{S}$. We assume that the sampling rule determining $\mathcal{S}$ depends only on the treatment allocations. Because the treatment allocation is independent of other variables, we can assume in the state evolution equations outlined in Eq. (27) that ${M}_{\cdot}^{\mathcal{S}}$ and ${\vec{X}}^{\mathcal{S}}$ are equal to their global counterparts ${M}_{\cdot}$ and ${\vec{X}}$, respectively. This reflects the idea that sampling based on treatment allocation is equivalent to random sampling from the experimental population. Thus, by the state evolution equations, for $t=0,1,\ldots,l-1$, we can write,
$\nu_{t}{}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{},\quad\nu_{t}^{\mathcal{S}}{}=\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t}{},$ (92)
and for $t\geq l-1$,
$\tilde{\nu}_{t+1}{}$ $=\delta_{g}+\theta_{gl}\nu_{t-l+1}{}+\ldots+\theta_{g1}\nu_{t}{}+\tau_{g}p_{t+1}+\lambda_{g}p_{t+1}\nu_{t}{}$ (93)
$\nu_{t+1}^{\mathcal{S}}{}$ $=\tilde{\nu}_{t+1}{}+\delta_{h}+\theta_{hl}\nu_{t-l+1}^{\mathcal{S}}{}+\ldots+\theta_{h1}\nu_{t}^{\mathcal{S}}{}+\tau_{h}p_{t+1}^{\mathcal{S}}+\lambda_{h}p_{t+1}^{\mathcal{S}}\nu_{t}^{\mathcal{S}}{},$
where
$(\delta_{g},\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\delta_{h},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top}$ (94)
::=:= $\mathbb{E}\left[{\vec{X}}=(\Delta_{g},\Theta_{gl},\ldots,\Theta_{g1},\mathcal{T}_{g},\Lambda_{g},\Delta_{h},\Theta_{hl},\ldots,\Theta_{h1},\mathcal{T}_{h},\Lambda_{h})^{\top}\right].$
Here, ${\vec{X}}$ represents the weak limit of ${\vec{X}}^{1},\ldots,{\vec{X}}^{N}$ when $N\rightarrow\infty$, as specified by Assumption 9.4. Then, Equation (93) follows from Assumption 10 and the additional assumption about the elements of ${\vec{X}}$. For example, (93) holds when the elements of ${\vec{X}}$ are random variables independent of all other sources of randomness in the model.
To proceed with counterfactual estimation, we consider $b$ distinct subpopulations, denoted by $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$, each determined solely by treatment allocations. With convention $\delta:=\delta_{g}+\delta_{h}$ in (93), we can write the following linear regression model:
$\begin{bmatrix}\nu_{l}^{\mathcal{S}_{1}}{}\\ \nu_{l+1}^{\mathcal{S}_{1}}{}\\ \vdots\\ \nu_{T}^{\mathcal{S}_{1}}{}\\ \\ \vdots\\ \\ \nu_{l}^{\mathcal{S}_{b}}{}\\ \nu_{l+1}^{\mathcal{S}_{b}}{}\\ \vdots\\ \nu_{T}^{\mathcal{S}_{b}}{}\end{bmatrix}=\delta\begin{bmatrix}1\\ 1\\ \vdots\\ 1\\ \\ \vdots\\ \\ 1\\ 1\\ \vdots\\ 1\end{bmatrix}$ $+\theta_{gl}\begin{bmatrix}\nu_{0}{}\\ \nu_{1}{}\\ \vdots\\ \nu_{T-l}{}\\ \\ \vdots\\ \\ \nu_{0}{}\\ \nu_{1}{}\\ \vdots\\ \nu_{T-l}{}\end{bmatrix}+\ldots+\theta_{g1}\begin{bmatrix}\nu_{l-1}{}\\ \nu_{l}{}\\ \vdots\\ \nu_{T-1}{}\\ \\ \vdots\\ \\ \nu_{l-1}{}\\ \nu_{l}{}\\ \vdots\\ \nu_{T-1}{}\end{bmatrix}+\tau_{g}\begin{bmatrix}p_{l}\\ p_{l+1}\\ \vdots\\ p_{T}\\ \\ \vdots\\ \\ p_{l}\\ p_{l+1}\\ \vdots\\ p_{T}\end{bmatrix}+\lambda_{g}\begin{bmatrix}p_{l}\nu_{l-1}{}\\ p_{l+1}\nu_{l}{}\\ \vdots\\ p_{T}\nu_{T-1}{}\\ \\ \vdots\\ \\ p_{l}\nu_{l-1}{}\\ p_{l+1}\nu_{l}{}\\ \vdots\\ p_{T}\nu_{T-1}{}\end{bmatrix}$ (95)
$+\theta_{hl}\begin{bmatrix}\nu_{0}^{\mathcal{S}_{1}}{}\\ \nu_{1}^{\mathcal{S}_{1}}{}\\ \vdots\\ \nu_{T-l}^{\mathcal{S}_{1}}{}\\ \\ \vdots\\ \\ \nu_{0}^{\mathcal{S}_{b}}{}\\ \nu_{1}^{\mathcal{S}_{b}}{}\\ \vdots\\ \nu_{T-l}^{\mathcal{S}_{b}}{}\end{bmatrix}+\ldots+\theta_{h1}\begin{bmatrix}\nu_{l-1}^{\mathcal{S}_{1}}{}\\ \nu_{l}^{\mathcal{S}_{1}}{}\\ \vdots\\ \nu_{T-1}^{\mathcal{S}_{1}}{}\\ \\ \vdots\\ \\ \nu_{l-1}^{\mathcal{S}_{b}}{}\\ \nu_{l}^{\mathcal{S}_{b}}{}\\ \vdots\\ \nu_{T-1}^{\mathcal{S}_{b}}{}\end{bmatrix}+\tau_{h}\begin{bmatrix}p_{l}^{\mathcal{S}_{1}}\\ p_{l+1}^{\mathcal{S}_{1}}\\ \vdots\\ p_{T}^{\mathcal{S}_{1}}\\ \\ \vdots\\ \\ p_{l}^{\mathcal{S}_{b}}\\ p_{l+1}^{\mathcal{S}_{b}}\\ \vdots\\ p_{T}^{\mathcal{S}_{b}}\end{bmatrix}+\lambda_{h}\begin{bmatrix}p_{l}^{\mathcal{S}_{1}}\nu_{l-1}^{\mathcal{S}_{1}}{}\\ p_{l+1}^{\mathcal{S}_{1}}\nu_{l}^{\mathcal{S}_{1}}{}\\ \vdots\\ p_{T}^{\mathcal{S}_{1}}\nu_{T-1}^{\mathcal{S}_{1}}{}\\ \\ \vdots\\ \\ p_{l}^{\mathcal{S}_{b}}\nu_{l-1}^{\mathcal{S}_{b}}{}\\ p_{l+1}^{\mathcal{S}_{b}}\nu_{l}^{\mathcal{S}_{b}}{}\\ \vdots\\ p_{T}^{\mathcal{S}_{b}}\nu_{T-1}^{\mathcal{S}_{b}}{}\end{bmatrix},$
or equivalently in a matrix form
$\vec{{\cal Y}}={\cal X}(\delta,\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top},$
where $\vec{{\cal Y}}$ represents the vector on the left-hand side of (95), while ${\cal X}$ denotes the matrix formed by the columns corresponding to the vectors on the right-hand side of (95).
For the regression model outlined in (95), we now show that the least squares estimator provides a strongly consistent estimate for the unknown coefficients. To this end, upon observing $\bm{w}{}$ and $\bm{Y}{}(\bm{w}{})$ within a system of $N$ experimental units, we define the following:
$\vec{{\cal Y}}(N)\sim{\cal X}(N)\vec{{\cal B}}(N),$ (96)
where $\vec{{\cal Y}}(N)$ denotes the sample mean of observed outcomes over time and across various subpopulations, corresponding to the vector on the left-hand side of (95). The vector $\vec{{\cal B}}(N)$ represents the unknown coefficients on the right-hand side of (95), while ${\cal X}(N)$ denotes a $b(T-l+1)$ by $21+l+2+l+$ 2 matrix. This matrix is aligned with the vectors on the right-hand side of (95), but with the corresponding sample means of observed outcomes and treatments replacing the asymptotic terms, see (97).
Proposition 10.5
Suppose that ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible. Let ^$\vec{\widehat{{\cal B}}}(N):=\big{(}{\cal X}(N)^{\top}{\cal X}(N)\big{)}^{-1}{\cal X}(N)^{\top}\vec{{\cal Y}}(N)$ be the least squares estimator. Then, ^$\vec{\widehat{{\cal B}}}(N)$ provides a strongly consistent estimator for the coefficient vector $(\delta,\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top}$.
Proof. Because the matrix ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible, the estimator ^$\vec{\widehat{{\cal B}}}(N)$ is a continuous function of the input data. Therefore, we can pass the limit through the estimator function as:
$\lim_{N\rightarrow\infty}\vec{\widehat{{\cal B}}}(N)=\left(\Big{(}\lim_{N\rightarrow\infty}{\cal X}(N)\Big{)}^{\top}\Big{(}\lim_{N\rightarrow\infty}{\cal X}(N)\Big{)}\right)^{-1}\Big{(}\lim_{N\rightarrow\infty}{\cal X}(N)\Big{)}^{\top}\Big{(}\lim_{N\rightarrow\infty}\vec{{\cal Y}}(N)\Big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;({\cal X}^{\top}{\cal X})^{-1}{\cal X}^{\top}\vec{{\cal Y}},$
where we used the result of Theorem 9.8. Now, note that (95) defines a deterministic regression model, and $({\cal X}^{\top}{\cal X})^{-1}{\cal X}^{\top}\vec{{\cal Y}}=(\delta,\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top}$. This concludes the proof. $\Box$
To ensure the invertibility condition of the matrix ${\cal X}(N)^{\top}{\cal X}(N)$, we need to set some simple conditions on the experimental design. Basically, conducting the experiment in more than one stage (equivalent to having two distinct values for elements of the vector $\big{(}\hat{p}_{l},\ldots,\hat{p}_{T}\big{)}^{\top}$ should suffice. This is needed to ensure that the first column of ${\cal X}(N)$ (i.e., $\vec{1}_{b(T-l+1)}$) is linearly independent of the column $\big{(}\hat{p}_{l},\ldots,\hat{p}_{T},\ldots,\hat{p}_{l},\ldots,\hat{p}_{T}\big{)}^{\top}$.
Assuming a non-zero treatment effect (i.e., $\tau_{h}\neq 0$), we can choose the batches to ensure enough variation across different batches. Therefore, upon a careful batching, columns of ${\cal X}(N)$ are linearly independent as each has its own specific variation patterns over time and/or subpopulations.
Although the exact value of $\tau$ is unknown, we suppose that contextual information suggests the presence of a non-zero direct treatment effect. In the case where $\tau=0$, according to contextual information, both the subpopulation sample mean $\nu_{t}^{\mathcal{S}}{}$ and the population sample mean $\nu_{t}{}$ are equal in (92) and (93), allowing for further simplification of the underlying model.
[…]
10.2 First-order Estimators
Given the observed outcomes $\bm{Y}{}(\bm{w}_{o}{})$, we can leverage Theorem 10.1 and Proposition 10.5 to consistently estimate counterfactuals under a desired treatment allocation $\bm{w}_{u}{}$. To this end, we propose two closely related families of estimators, which we detail below. Both family of estimators require that the delivered treatment allocation $\bm{w}_{o}{}$ and the desired treatment allocation $\bm{w}_{u}{}$ match during the first $l$ periods. These initial $l$ periods serve as the common foundation from which counterfactual trajectories are constructed.
Given subpopulation $\mathcal{S}$, Algorithms 4 and 5 aim to estimate counterfactuals under the desired treatment allocation over $\mathcal{S}$ using $b$ distinct subpopulations $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$ as the estimation samples. Both algorithms share their first two steps. In the first step, they compute sample means of observed outcomes for both the entire population and each subpopulation, along with sample means of delivered and desired treatment allocations. The second step estimates unknown parameters in the state evolution equation (93) using least squares estimation as detailed in Proposition 10.5. The algorithms then diverge in their third step, applying these results through two distinct approaches detailed below. The consistency proofs for both algorithms follow directly from earlier results and are omitted for brevity.
10.2.1 Semi-recursive Estimation Method
This estimator, outlined in Algorithm 4, builds on the observed sample means in its third step. It uses the parameter estimates from the second step and the state evolution equation (93) to modify the observed sample means by adjusting the treatment level to the desired one. This approach transfers the original complexities of the observed outcomes to the estimated counterfactual, making it particularly suitable for scenarios with strong time trends or complex baselines. This estimator generalizes the algorithm proposed in Shirani and Bayati, (2024) by accommodating broader model classes and providing more general estimands.
Algorithm 4 First-order semi-recursive counterfactual estimator
$\bm{Y}{}(\bm{w}_{o}{}),\bm{w}_{o}{}$, $\bm{w}_{u}{}$, estimation batch $\mathcal{S}$, sample batches $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$, and $l$
Step 1: Data processing
for $t=0,\ldots,T$ do
     $\hat{\nu}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{}(\bm{w}_{o}{})$
     $\hat{\nu}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t}{}(\bm{w}_{o}{})$
     $\hat{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{o,t}{}$
     $\hat{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{o,t}{}$
     $\check{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{u,t}{}$
     $\check{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{u,t}{}$
     for $j=1,\ldots,b$ do
         $\hat{\nu}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}Y^{i}_{t}{}(\bm{w}_{o}{})$
         $\hat{p}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}w^{i}_{o,t}{}$
     end for
end for
Step 2: Parameters estimation
$(\hat{\Delta},\hat{\Theta}_{gl},\ldots,\hat{\Theta}_{g1},\hat{\mathcal{T}}_{g},\hat{\Lambda}_{g},\hat{\Theta}_{hl},\ldots,\hat{\Theta}_{h1},\hat{\mathcal{T}}_{h},\hat{\Lambda}_{h})^{\top}\leftarrow\big{(}\hat{{\cal X}}^{\top}\hat{{\cal X}}\big{)}^{-1}\hat{{\cal X}}^{\top}\vec{\hat{{\cal Y}}}$
Step 3: Counterfactual estimation
$\Big{(}\widehat{\text{CFE}}_{0}{(\bm{w}_{u}{})},\ldots,\widehat{\text{CFE}}_{l-1}{(\bm{w}_{u}{})}\Big{)}\leftarrow\left(\hat{\nu}_{0}{},\ldots,\hat{\nu}_{l-1}{}\right)$
$\Big{(}\widehat{\text{CFE}}^{\mathcal{S}}_{0}{(\bm{w}_{u}{})},\ldots,\widehat{\text{CFE}}^{\mathcal{S}}_{l-1}{(\bm{w}_{u}{})}\Big{)}\leftarrow\left(\hat{\nu}_{0}^{\mathcal{S}}{},\ldots,\hat{\nu}_{l-1}^{\mathcal{S}}{}\right)$
for $t=l,\ldots,T$ do
     ${\cal R}_{g}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{gj}(\widehat{\text{CFE}}_{t-j}{(\bm{w}_{u}{})}-\hat{\nu}_{t-j}{})+\hat{\mathcal{T}}_{g}(\check{p}_{t}{}-\hat{p}_{t}{})+\hat{\Lambda}_{g}(\check{p}_{t}{}\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})}-\hat{p}_{t}{}\hat{\nu}_{t-1}{})$
     ${\cal R}_{h}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{hj}(\widehat{\text{CFE}}_{t-j}{(\bm{w}_{u}{})}-\hat{\nu}_{t-j}{})+\hat{\mathcal{T}}_{h}(\check{p}_{t}{}-\hat{p}_{t}{})+\hat{\Lambda}_{h}(\check{p}_{t}{}\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})}-\hat{p}_{t}{}\hat{\nu}_{t-1}{})$
     ${\cal R}_{h}^{\mathcal{S}}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{hj}(\widehat{\text{CFE}}^{\mathcal{S}}_{t-j}{(\bm{w}_{u}{})}-\hat{\nu}_{t-j}^{\mathcal{S}}{})+\hat{\mathcal{T}}_{h}(\check{p}_{t}^{\mathcal{S}}{}-\hat{p}_{t}^{\mathcal{S}}{})+\hat{\Lambda}_{h}(\check{p}_{t}^{\mathcal{S}}{}\widehat{\text{CFE}}^{\mathcal{S}}_{t-1}{(\bm{w}_{u}{})}-\hat{p}_{t}^{\mathcal{S}}{}\hat{\nu}_{t-1}^{\mathcal{S}}{})$
     $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}\leftarrow\hat{\nu}_{t}{}+{\cal R}_{g}+{\cal R}_{h}$
     $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}\leftarrow\hat{\nu}_{t}^{\mathcal{S}}{}+{\cal R}_{g}+{\cal R}_{h}^{\mathcal{S}}$
end for
$\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ and $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}$, for $t=0,\ldots,T$.
10.2.2 Recursive Estimation Method
This estimator, outlined in Algorithm 5, directly leverages the state evolution equation and parameter estimates from the second step. Specifically, it estimates counterfactuals recursively, using only the past $l$ terms and desired treatment levels. Unlike Algorithm 4, this algorithm can estimate counterfactuals even for time blocks where no outcome data were collected. For example, having observed data until December, it can predict counterfactual outcomes for January without requiring any observations during this month.
Algorithm 5 First-order recursive counterfactual estimator
$\bm{Y}{}(\bm{w}_{o}{}),\bm{w}_{o}{}$, $\bm{w}_{u}{}$, estimation batch $\mathcal{S}$, sample batches $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$, and $l$
Step 1: Data processing
for $t=0,\ldots,T$ do
     $\hat{\nu}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{}(\bm{w}_{o}{})$
     $\hat{\nu}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t}{}(\bm{w}_{o}{})$
     $\hat{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{o,t}{}$
     $\hat{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{o,t}{}$
     $\check{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{u,t}{}$
     $\check{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{u,t}{}$
     for $j=1,\ldots,b$ do
         $\hat{\nu}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}Y^{i}_{t}{}(\bm{w}_{o}{})$
         $\hat{p}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}w^{i}_{o,t}{}$
     end for
end for
Step 2: Parameters estimation
$(\hat{\Delta},\hat{\Theta}_{gl},\ldots,\hat{\Theta}_{g1},\hat{\mathcal{T}}_{g},\hat{\Lambda}_{g},\hat{\Theta}_{hl},\ldots,\hat{\Theta}_{h1},\hat{\mathcal{T}}_{h},\hat{\Lambda}_{h})^{\top}\leftarrow\big{(}\hat{{\cal X}}^{\top}\hat{{\cal X}}\big{)}^{-1}\hat{{\cal X}}^{\top}\vec{\hat{{\cal Y}}}$
Step 3: Counterfactual estimation
$\Big{(}\widehat{\text{CFE}}_{0}{(\bm{w}_{u}{})},\ldots,\widehat{\text{CFE}}_{l-1}{(\bm{w}_{u}{})}\Big{)}\leftarrow\left(\hat{\nu}_{0}{},\ldots,\hat{\nu}_{l-1}{}\right)$
$\Big{(}\widehat{\text{CFE}}^{\mathcal{S}}_{0}{(\bm{w}_{u}{})},\ldots,\widehat{\text{CFE}}^{\mathcal{S}}_{l-1}{(\bm{w}_{u}{})}\Big{)}\leftarrow\left(\hat{\nu}_{0}^{\mathcal{S}}{},\ldots,\hat{\nu}_{l-1}^{\mathcal{S}}{}\right)$
for $t=l,\ldots,T$ do
     ${\cal R}_{g}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{gj}\widehat{\text{CFE}}_{t-j}{(\bm{w}_{u}{})}+\hat{\mathcal{T}}_{g}\check{p}_{t}{}+\hat{\Lambda}_{g}\check{p}_{t}{}\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})}$
     ${\cal R}_{h}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{hj}\widehat{\text{CFE}}_{t-j}{(\bm{w}_{u}{})}+\hat{\mathcal{T}}_{h}\check{p}_{t}{}+\hat{\Lambda}_{h}\check{p}_{t}{}\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})}$
     ${\cal R}_{h}^{\mathcal{S}}\leftarrow\sum_{j=1}^{l}\hat{\Theta}_{hj}\widehat{\text{CFE}}^{\mathcal{S}}_{t-j}{(\bm{w}_{u}{})}+\hat{\mathcal{T}}_{h}\check{p}_{t}^{\mathcal{S}}{}+\hat{\Lambda}_{h}\check{p}_{t}^{\mathcal{S}}{}\widehat{\text{CFE}}^{\mathcal{S}}_{t-1}{(\bm{w}_{u}{})}$
     $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}\leftarrow{\cal R}_{g}+{\cal R}_{h}$
     $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}\leftarrow{\cal R}_{g}+{\cal R}_{h}^{\mathcal{S}}$
end for
$\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ and $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}$, for $t=0,\ldots,T$.
10.3 Higher-order Recursive Estimators
In light of Theorem 10.1, we can extend our approach to utilize higher-order approximations of $g_{t}$ and $h_{t}{}$. The estimator, outlined in Algorithm 6 for $l=1$ lag terms, incorporates up to order $m\geq 2$ moments of the unit outcomes. Precisely, we introduce two families of feature functions: $\vec{\phi}=(\phi_{1},\ldots,\phi_{n_{1}})^{\top}$ for population-level moments and $\vec{\psi}=(\psi_{1},\ldots,\psi_{n_{2}})^{\top}$ for subpopulation-level moments. The approach employs linear regression to estimate weights for a linear combination of these features. This can be realized as a generalization of Bayati et al., (2024)’s method and enables capturing more complex patterns in counterfactuals by leveraging richer information about the outcomes’ distributions over time. When the feature functions $\vec{\phi}$ and $\vec{\psi}$ are continuous, consistency follows from Theorem 10.1, though we defer rigorous treatment to future work.
Algorithm 6 Higher-order recursive counterfactual estimator
$\bm{Y}{}(\bm{w}_{o}{}),\bm{w}_{o}{}$, $\bm{w}_{u}{}$, $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$, $m\geq 2$, $\vec{\phi}=(\phi_{1},\ldots,\phi_{n_{1}})^{\top}$, and $\vec{\psi}=(\psi_{1},\ldots,\psi_{n_{2}})^{\top}$
Step 1: Data processing
for $t=0,\ldots,T$ do
    $\hat{\nu}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{}(\bm{w}_{o}{})$
    $\hat{\nu}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}Y^{i}_{t}{}(\bm{w}_{o}{})$
    for $k=2,\ldots,m$ do
         $\hat{\rho}_{t}^{(k)}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(Y^{i}_{t}{}(\bm{w}_{o}{})-\hat{\nu}_{t}{}\right)^{k}$
         $\hat{\rho}_{t}^{\mathcal{S},(k)}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\left(Y^{i}_{t}{}(\bm{w}_{o}{})-\hat{\nu}_{t}^{\mathcal{S}}{}\right)^{k}$
    end for
    $\hat{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{o,t}{}$
    $\hat{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{o,t}{}$
    $\check{p}_{t}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}w^{i}_{u,t}{}$
    $\check{p}_{t}^{\mathcal{S}}{}\leftarrow\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}w^{i}_{u,t}{}$
    for $j=1,\ldots,b$ do
         $\hat{\nu}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}Y^{i}_{t}{}(\bm{w}_{o}{})$
         for $k=2,\ldots,m$ do
             $\hat{\rho}_{t}^{\mathcal{S}_{j},(k)}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}\left(Y^{i}_{t}{}(\bm{w}_{o}{})-\hat{\nu}_{t}^{\mathcal{S}_{j}}{}\right)^{k}$
         end for
         $\hat{p}_{t}^{\mathcal{S}_{j}}{}\leftarrow\frac{1}{\left|{\mathcal{S}_{j}}\right|}\sum_{i\in\mathcal{S}_{j}}w^{i}_{o,t}{}$
    end for
end for
Step 2: Parameters estimation
Estimate $\bm{\Theta}_{g}\in\mathbb{R}^{m\times n_{1}}$ and $\bm{\Theta}_{h}\in\mathbb{R}^{m\times n_{2}}$ as $\widehat{\bm{\Theta}}_{g}$ and $\widehat{\bm{\Theta}}_{h}$:
$(\hat{\nu}_{t+1}^{\mathcal{S}_{j}}{},\hat{\rho}_{t+1}^{\mathcal{S}_{j},(2)}{},\ldots,\hat{\rho}_{t+1}^{\mathcal{S}_{j},(m)}{})^{\top}=\bm{\Theta}_{g}\vec{\phi}(\hat{\nu}_{t}{},\hat{\rho}_{t}^{(2)}{},\ldots,\hat{\rho}_{t}^{(m)}{},\hat{p}_{t+1}{})+\bm{\Theta}_{h}\vec{\psi}(\hat{\nu}_{t}^{\mathcal{S}_{j}}{},\hat{\rho}_{t}^{\mathcal{S}_{j},(2)}{},\ldots,\hat{\rho}_{t}^{\mathcal{S}_{j},(m)}{},\hat{p}_{t+1}^{\mathcal{S}_{j}}{}),$
where $j=1,\ldots,b$ and $t=0,\ldots,T-1$.
Step 3: Counterfactual estimation
$\widehat{\text{CFE}}_{0}{(\bm{w}_{u}{})}\leftarrow\hat{\nu}_{0}{}$ and $(\check{\rho}_{0}^{(2)}{},\ldots,\check{\rho}_{0}^{(m)}{})\leftarrow(\hat{\rho}_{0}^{(2)}{},\ldots,\hat{\rho}_{0}^{(m)}{})$
$\widehat{\text{CFE}}^{\mathcal{S}}_{0}{(\bm{w}_{u}{})}\leftarrow\hat{\nu}_{0}^{\mathcal{S}}{}$ and $(\check{\rho}_{0}^{\mathcal{S},(2)}{},\ldots,\check{\rho}_{0}^{\mathcal{S},(m)}{})\leftarrow(\hat{\rho}_{0}^{\mathcal{S},(2)}{},\ldots,\hat{\rho}_{0}^{\mathcal{S},(m)}{})$
for $t=1,\ldots,T$ do
    $\vec{\cal R}_{g}\leftarrow\widehat{\bm{\Theta}}_{g}\vec{\phi}(\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})},\check{\rho}_{t-1}^{(2)}{},\ldots,\check{\rho}_{t-1}^{(m)}{},,\check{p}_{t}{})$
    $\vec{\cal R}_{h}\leftarrow\widehat{\bm{\Theta}}_{h}\vec{\psi}(\widehat{\text{CFE}}_{t-1}{(\bm{w}_{u}{})},\check{\rho}_{t-1}^{(2)}{},\ldots,\check{\rho}_{t-1}^{(m)}{},\check{p}_{t}{})$
    $\vec{\cal R}_{h}^{\mathcal{S}}\leftarrow\widehat{\bm{\Theta}}_{h}\vec{\psi}(\widehat{\text{CFE}}^{\mathcal{S}}_{t-1}{(\bm{w}_{u}{})},\check{\rho}_{t-1}^{\mathcal{S},(2)}{},\ldots,\check{\rho}_{t-1}^{\mathcal{S},(m)}{},\check{p}_{t}^{\mathcal{S}}{})$
    $(\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})},\check{\rho}_{t}^{(2)}{},\ldots,\check{\rho}_{t}^{(m)}{})^{\top}\leftarrow\vec{\cal R}_{g}+\vec{\cal R}_{h}$
    $(\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})},\check{\rho}_{t}^{\mathcal{S},(2)}{},\ldots,\check{\rho}_{t}^{\mathcal{S},(m)}{})^{\top}\leftarrow\vec{\cal R}_{g}+\vec{\cal R}_{h}^{\mathcal{S}}$
end for
$\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ and $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}$, for $t=0,\ldots,T$.
11 Detrending for Temporal Patterns
While semi-recursive estimators (see §10.2.1) can capture complex temporal patterns in unit outcomes, they face two major limitations. First, they cannot estimate out-of-sample counterfactuals because their architecture relies directly on observed sample means. Second, unlike recursive estimators (see Algorithm 6), they cannot accommodate higher-order approximations of the outcome functions ($g_{t}$ and $h_{t}{}$). This limitation stems from their dependence on the closed-form of state evolution equation (as outlined in (93)), which may not exist for more complex characterizations of the outcome functions.
This section develops a two-stage estimation method that combines the advantages of both semi-recursive and recursive estimators. Although it requires an additional structural assumption on the outcome specification, this approach can handle complex temporal patterns while enabling both out-of-sample counterfactual estimation and higher-order approximations of the outcome functions. The method proceeds as follows: first, we employ a semi-recursive estimator with sufficient lag terms to accurately estimate temporal patterns. We use this to estimate the baseline outcome means (the counterfactual for all control units: $\text{CFE}_{t}{(\mathbf{0})},\;t=0,1,\ldots,T$). Next, we detrend the observed outcomes by subtracting the estimated baseline from them. We then apply a recursive estimator focused specifically on estimating treatment effects in the absence of temporal patterns. Finally, we add back the subtracted baseline to obtain the desired estimand (see Algorithm 7). The following sections provide more detailed expositions of the algorithm.
11.1 Baseline Outcome Estimation
Letting $\vec{y}^{\;}_{t}{}:=\vec{Y}_{t}{}(\bm{W}{}=\mathbf{0})$ denote the vector of baseline outcomes at time $t=0,1,\ldots,T$ under no treatment, we can write from (3):
$\vec{y}^{\;}_{t+1}{}=\vec{Y}_{t+1}{}(\bm{W}{}=\mathbf{0})$ $=\big{(}{\bf A}+{\bf B}_{t}\big{)}g_{t}\left(\vec{y}^{\;}_{t}{},\mathbf{0},\bm{X}\right)+h_{t}{}\left(\vec{y}^{\;}_{t}{},\mathbf{0},\bm{X}\right).$ (98)
Thus, the matrix $\bm{y}{}=[\vec{y}^{\;}_{0}{}|\ldots|\vec{y}^{\;}_{T}{}]$ represents the panel data of baseline outcomes that would be observed in the absence of any intervention.
In the first step of Algorithm 7, we employ a semi-recursive algorithm to estimate the sample means of the columns of $\bm{y}{}$, denoted by $\widehat{\text{CFE}}_{t}{(\mathbf{0})}$. The consistency of this estimation follows directly from the consistency of Algorithm 4.
11.2 Augmented Causal Message-Passing Model
The next two steps of Algorithm 7 require the following assumption. {assumption} For $t=0,1,\ldots,T-1$, we assume there exist families of functions $\tilde{g}_{t}$ and $\tilde{h}_{t}$ such that the potential outcomes $\vec{Y}_{t}{}(\bm{W}{})$ satisfy:
$\vec{Y}_{t+1}{}(\bm{W}{})=\vec{y}^{\;}_{t+1}{}+\big{(}{\bf A}+{\bf B}_{t}\big{)}\tilde{g}_{t}\left(\vec{Y}_{t}{}(\bm{W}{})-\vec{y}^{\;}_{t}{},\bm{W}{},\bm{X}\right)+\tilde{h}_{t}\left(\vec{Y}_{t}{}(\bm{W}{})-\vec{y}^{\;}_{t}{},\bm{W}{},\bm{X}\right).$ (99)
Additionally, $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$, and functions $\tilde{g}_{t}$ and $\tilde{h}_{t}$ satisfy the conditions of Assumption 9.4. Note that enforcing conditions $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$ ensures that $\vec{Y}_{t+1}{}(\mathbf{0})=\vec{y}^{\;}_{t+1}{}$, aligning with (98). Letting $\vec{{\cal Y}}_{t}{}(\bm{W}{}):=\vec{Y}_{t}{}(\bm{W}{})-\vec{y}^{\;}_{t}{}$, we can then rewrite the augmented model (99) to match the dynamics of the original outcome specification:
$\vec{{\cal Y}}_{t+1}{}(\bm{W}{})$ $=\big{(}{\bf A}+{\bf B}_{t}\big{)}\tilde{g}_{t}\left(\vec{{\cal Y}}_{t}{}(\bm{W}{}),\bm{W}{},\bm{X}\right)+\tilde{h}_{t}\left(\vec{{\cal Y}}_{t}{}(\bm{W}{}),\bm{W}{},\bm{X}\right).$ (100)
We emphasize that Equations (98) and (100) provide distinct characterizations of the outcomes, and Algorithm 7 requires both to hold simultaneously. Specifically, assuming the conditions of §9.4 hold in both settings, the baseline outcomes $\vec{y}^{\;}_{t}{}$ satisfy the state evolution equation corresponding to (98). However, in the context of (100), state evolution becomes relevant only when treatment is delivered (i.e., $\bm{W}{}\neq\mathbf{0}$). Indeed, under Assumption 11.2, the third condition in Assumption 9.4 indicates that state evolution can be derived only when there exists a non-zero treatment effect.
Then, the consistency of the second step estimation in Algorithm 7 holds in the context of the outcome specification (99). Finally, the consistency of the ultimate estimator follows from both the consistency of individual steps and the fact that Algorithm 7 can be viewed as a combination of continuous functions.
Algorithm 7 First-order counterfactual estimator with preprocessing
$\bm{Y}{}(\bm{w}_{o}{}),\bm{w}_{o}{}$, $\bm{w}_{u}{}$, estimation batch $\mathcal{S}$, sample batches $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$, and $l$
Step 1: Detrending
Use Algorithm 4 with $\bm{w}_{u}{}=\mathbf{0}$ to obtain $\widehat{\text{CFE}}_{t}{(\mathbf{0})}$, $t=0,\ldots,T$.
for $t=0,\ldots,T$ do
     $\vec{Y}^{{}^{\prime}}_{t}{}\leftarrow\vec{Y}_{t}{}(\bm{W}{}=\bm{w}_{o}{})-\widehat{\text{CFE}}_{t}{(\mathbf{0})}$
end for
Step 2: Counterfactual estimation with proprocessed data
Use Algorithm 5 with $\bm{Y}^{{}^{\prime}}{}$ to obtain $\widehat{\text{CFE}}^{{}^{\prime}\mathcal{S}}_{t}{(\bm{w}_{u}{})}$, $t=0,\ldots,T$.
Step 3: Post-processing
for $t=0,\ldots,T$ do
     $\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}\leftarrow\widehat{\text{CFE}}^{{}^{\prime}\mathcal{S}}_{t}{(\bm{w}_{u}{})}+\widehat{\text{CFE}}_{t}{(\mathbf{0})}$
end for
$\widehat{\text{CFE}}^{\mathcal{S}}_{t}{(\bm{w}_{u}{})}$, $t=0,\ldots,T$.
12 Auxiliary Results
We need the following strong law of large numbers (SLLN) for triangular arrays of independent but not identically distributed random variables. The form stated below is Theorem 3 in Bayati and Montanari, (2011) that is adapted from Theorem 2.1 in Hu and Taylor, (1997).
Theorem 12.1 (SLLN)
Let $\left\{X_{n,i}:1\leq i\leq n,\;n\geq 1\right\}$ be a triangular array of random variables such that $(X_{n,1},\ldots,X_{n,n})$ are mutually independent with a mean equal to zero for each $n$ and $\frac{1}{n}\sum_{i=1}^{n}E\left[|X_{n,i}|^{2+\kappa}\right]\leq cn^{\kappa/2}$ for some $10<\kappa<$ 1 and $c<\infty$. Then, we have
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=1}^{n}X_{n,i}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;0.$ (101)
We also need the following form of the law of large numbers which is an extension of Lemma 4 in Bayati and Montanari, (2011).
Theorem 12.2
Fix $k\geq 2$ and an integer $l$ and let $\left\{\bm{v}(N)\right\}_{N\geq 1}$ be a sequence of vectors that $\bm{v}(N)\in\mathbb{R}^{N\times l}$. That means, $\bm{v}(N)$ is a matrix with $N$ rows and $l$ columns. Assume that the empirical distribution of $\bm{v}(N)$, denoted by $\hat{p}_{N}$, converges weakly to a probability measure $p_{v}$ on $\mathbb{R}^{l}$ such that $\mathbb{E}_{p_{v}}\left[{\left\|\vec{V}\right\|}^{k}\right]<\infty$ and $\mathbb{E}_{\hat{p}_{N}}\left[{\left\|\vec{V}\right\|}^{k}\right]\rightarrow\mathbb{E}_{p_{v}}\left[{\left\|\vec{V}\right\|}^{k}\right]$ as $N\rightarrow\infty$. Then, for any continuous function $f:\mathbb{R}^{l}\mapsto\mathbb{R}$ with at most polynomial growth of order $k$, we have
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}f\big{(}\bm{v}_{n}(N)\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}_{p_{v}}[f(\vec{V})].$ (102)
Next, we present Lemma 9 from Li and Wei, (2022), and then employ it to prove Lemma 12.4.
Lemma 12.3
Fixing $N$ and $t<N$, consider a set of i.i.d. random vectors $\vec{\phi_{i}}\sim{\cal N}(0,\frac{1}{N}{\bf I}_{N}),\;i=1,\ldots,t$, and any unit vector $\vec{\beta}=(\beta^{1},\ldots,\beta^{t})^{\top}$ that might be statistically dependent on $\{\vec{\phi_{i}}\}_{i=1}^{t}$. Then, the 1-Wasserstein distance between the distribution of $\sum_{i=1}^{t}\beta^{i}\vec{\phi_{i}}$, denoted by $\mathcal{L}(\sum_{i=1}^{t}\beta^{i}\vec{\phi_{i}})$, and ${\cal N}(0,\frac{1}{N}{\bf I}_{N})$ obeys
$W_{1}\left(\mathcal{L}\left(\sum_{i=1}^{t}\beta^{i}\vec{\phi_{i}}\right),{\cal N}\left(0,\frac{1}{N}{\bf I}_{N}\right)\right)\leq c\sqrt{\frac{t\log N}{N}},$
for some constant $c$ that does not depend on $N$.
Lemma 12.4
Fixing $N$ and $t<N$, consider a set of i.i.d. random vectors $\vec{\phi_{i}}\sim{\cal N}(0,\frac{1}{N}{\bf I}_{N}),\;i=1,\ldots,t$, and any unit vector $\vec{\beta}=(\beta^{1},\ldots,\beta^{t})^{\top}$ that might be statistically dependent on $\{\vec{\phi_{i}}\}_{i=1}^{t}$. Let $\vec{\Phi}:=\sum_{i=1}^{t}\beta^{i}\vec{\phi_{i}}$ such that $\vec{\Phi}=(\Phi^{1},\ldots,\Phi^{N})^{\top}$ and $\mathcal{S}\subset[N]$ be subset of the indices. Then, the 1-Wasserstein distance between the distribution of $\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}\Phi^{n}$, denoted by $\mathcal{L}(\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}\Phi^{n})$, and ${\cal N}(0,{1}/{N})$ satisfies
$W_{1}\left(\mathcal{L}\left(\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}\Phi^{n}\right),{\cal N}\left(0,\frac{1}{N}\right)\right)\leq c\sqrt{\frac{t\log N}{N}},$
for some constant $c$ that does not depend on $N$ and $\mathcal{S}$.
Proof. Considering Kantorovich-Rubinstein duality, we use the dual representation of the 1-Wasserstein distance:
[…]
where $Z\sim{\cal N}(0,1)$. To proceed, fix the 1-Lipschitz function $f$ arbitrarily. Further, define the function $\psi:\mathbb{R}^{N}\mapsto\mathbb{R}$ such that for any vector $\vec{q}=(q^{1},\ldots,q^{N})^{\top}$, we have $\psi(\vec{q})=\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}q^{n}$. Then, we define $\tilde{f}=f\circ\psi:\mathbb{R}^{N}\mapsto\mathbb{R}$. Note that both $f$ and $\psi$ are continuous functions; as a result, the function $\tilde{f}$ is measurable. We show it is also $1$-Lipschitz. To this end, for vectors $\vec{q}_{1},\vec{q}_{2}\in\mathbb{R}^{N}$, we write:
$\left|\tilde{f}\left(\vec{q}_{2}\right)-\tilde{f}\left(\vec{q}_{1}\right)\right|=\left|f\left(\psi(\vec{q}_{2})\right)-f\left(\psi(\vec{q}_{1})\right)\right|\leq\left|\psi(\vec{q}_{2})-\psi(\vec{q}_{1})\right|\leq\frac{\sum_{n\in\mathcal{S}}\left|q^{n}_{2}-q^{n}_{1}\right|}{\sqrt{\left|{\mathcal{S}}\right|}}\leq\sqrt{\sum_{n\in\mathcal{S}}\left|q^{n}_{2}-q^{n}_{1}\right|^{2}}\leq{\left\|\vec{q}_{2}-\vec{q}_{1}\right\|},$
where we used the fact that $f$ is 1-Lipschitz and the Cauchy–Schwarz inequality. Therefore, the function $\tilde{f}$ is 1-Lipschitz and by the result of Lemma 12.3, we get
$\mathbb{E}\left[f\left(\frac{1}{\sqrt{N}}\sum_{n=1}^{N}\Phi^{n}\right)\right]-\mathbb{E}\left[f(\frac{Z}{\sqrt{N}})\right]=\mathbb{E}\left[\tilde{f}\left(\vec{\Phi}\right)\right]-\mathbb{E}\left[\tilde{f}\left(\frac{1}{\sqrt{N}}\vec{Z}\right)\right]\leq c\sqrt{\frac{t\log N}{N}},$
where $\vec{Z}\sim{\cal N}(0,{\bf I})$. Because $f$ is chosen arbitrarily, by Eq. (103), we obtain the desired result. $\Box$